### v1

# Tensor K-matrices for quantum symmetric pairs

Abstract.
Let ${\mathfrak{g}}$ be a symmetrizable Kac-Moody algebra, $U_{q}({\mathfrak{g}})$ its quantum group, and $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ a quantum symmetric pair subalgebra determined by a distinguished Lie algebra automorphism $\theta$. We introduce a category $\mathcal{W}_{\theta}$ of weight $U_{q}(\mathfrak{k})$-modules, which is acted upon by the category of weight $U_{q}({\mathfrak{g}})$-modules via tensor products. We construct a universal tensor K-matrix ${{\mathbb{K}}}$ (that is, a solution of a reflection equation) in a completion of $U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{g}})$. This yields a natural operator on any tensor product $M\otimes V$, where $M\in\mathcal{W}_{\theta}$ and $V\in{\mathcal{O}}_{\theta}$, i.e., $V$ is a $U_{q}({\mathfrak{g}})$-module in category ${\mathcal{O}}$ satisfying an integrability property determined by $\theta$. $\mathcal{W}_{\theta}$ is equipped with a canonical structure of a bimodule category over ${\mathcal{O}}_{\theta}$ and the action of ${{\mathbb{K}}}$ is encoded by a new categorical structure, which we refer to as a boundary structure on the bimodule category $\mathcal{W}_{\theta}$. This provides the most comprehensive algebraic framework to date for quantum integrability in the presence of boundaries. In particular, it generalizes a theorem of Kolb which describes a braided module structure on finite-dimensional $U_{q}(\mathfrak{k})$-modules when ${\mathfrak{g}}$ is finite-dimensional.
We apply our construction to the case of an affine Lie algebra, where it yields a formal tensor K-matrix valued in the endomorphisms of the tensor product of any module in $\mathcal{W}_{\theta}$ and any finite-dimensional module over the corresponding quantum affine algebra. We prove that this formal series can be normalized to a trigonometric K-matrix if the factors in the tensor product are both finite-dimensional irreducible modules over the quantum affine algebra.
Key words and phrases:
Reflection equation; quantum Kac-Moody algebras; quantum symmetric pairs
2020 Mathematics Subject Classification:
Primary: 81R50. Secondary: 16T25, 17B37, 81R12.
The first author is partially supported by the Programme FIL 2022 of the University of Parma and co-sponsored by Fondazione Cariparma.
1. Introduction
1.1.
The reflection equation was introduced by Cherednik in [Che84] in the context of quantum integrable systems with compatible boundary conditions, while the foundations for the underlying algebraic theory were laid by Sklyanin and Kulish in [Skl88, KS92]. One of Sklyanin’s observations in [Skl88] was that a solution of the reflection equation (a K-matrix) can be suitably combined with a solution of the Yang-Baxter equation (an R-matrix), to produce a new solution of the reflection equation which acts on a bigger space. Thus, more complex solutions can be obtained from simpler solutions. Additionally, in [KS92] it was shown that that coideal subalgebras of quasitriangular bialgebras (or, more generally, comodule algebras) play an essential role in the theory of the reflection equation. In this paper, we introduce a comprehensive algebraic framework for quantum integrability in the presence of boundaries and prove that a large class of examples arises from the theory of quantum symmetric pairs of arbitrary Kac-Moody type.
1.2.
The algebraic framework is summarized as follows. Let $A$ be any quasitriangular bialgebra with coproduct $\Delta:A\to A\otimes A$ and universal R-matrix $R\in(A\otimes A)^{\times}$. Next, let $B\subset A$ be any right coideal subalgebra: $\Delta(B)\subset B\otimes A$. In the case where $(A,B)$ is a quantum symmetric pair, see [Let02, Kol14], universal solutions of the reflection equation were recently found for increasingly general types in [BW18b, BK19, AV22a].
The axiomatic framework for such universal solutions is a cylindrical structure on a quasitriangular bialgebra $A$, introduced by the authors in [AV22a, Sec. 2]. This entails that there exists $K\in A^{\times}$, called (basic) universal K-matrix, satisfying a coproduct formula [^1] See Section A.2 for the precise definition.. The statement $R\cdot\Delta(K)=\Delta^{\sf op}(K)\cdot R$ is then equivalent to the following generalized universal (left) reflection equation, where $\psi:A\to A$ is an algebra automorphism called twist automorphism:
$R^{\psi\psi}_{21}\cdot(1\otimes K)\cdot R^{\psi}\cdot(K\otimes 1)=(K\otimes 1)\cdot R^{\psi}_{21}\cdot(1\otimes K)\cdot R.$ (1.1)
Here we have used the shorthand notations
$R^{\psi}=(\psi\otimes{\operatorname{id}})(R),\qquad R^{\psi}_{21}=(\psi\otimes{\operatorname{id}})(R)_{21},\qquad R_{21}^{\psi\psi}=(\psi\otimes\psi)(R)_{21}.$ (1.2)
In the case $A$ is an affine quantum group, we obtain a solid universal framework explaining the existence of matrix solutions of parameter-dependent reflection equations, see [AV22b], with the twist $\psi$ inducing the required inversion of the parameter. More precisely, it generates solutions of the generalized parameter-dependent reflection equation considered by Cherednik in [Che92, Sec. 4], which contains the one from [Che84] as a special case. Note the physical relevance of the framework of [Che92], which entails a system of two-boundary quantum Knizhnik-Zamolodchikov equations, naturally specializing to a formalism of Yang-type scattering matrices for a system of particles on an interval with generalized scattering at the boundaries.
In the framework of [AV22a], the subalgebra $B$ is related to the datum $(\psi,K)$ through the following linear relation:
$K\cdot b=\psi(b)\cdot K\qquad\forall\,b\in B.$ (1.3)
This provides a strong computational handle on $K$, and indeed this principle has already been used to compute K-matrices in particular representations, see e.g. [MN98, DG02, DM03]. Note however that the algebra $B$ appearing in (1.3) can be replaced by any subalgebra and hence it is natural to require a tighter relation between $B$ and $(\psi,K)$, bounding $B$ from below.
1.3.
In the present paper, we describe a natural universal 2-tensor version of the K-matrix, which we denote by ${{\mathbb{K}}}$ and call a tensor K-matrix. This generalizes the formalism with $\psi={\operatorname{id}}$ studied by Kolb in [Kol20], which builds on the works by Enriquez [Enr07] and Brochier [Bro13]. However, the latter does not apply to quantum groups of arbitrary type, and our construction aims to remove such an obstruction.
Sklyanin’s observations in [Skl88] inspire the following formula in terms of $R$, $\psi$ and $K$:
${{\mathbb{K}}}=(R^{\psi})_{21}\cdot(1\otimes K)\cdot R,$ (1.4)
which a priori lies in $A\otimes A$. If the support property
${{\mathbb{K}}}\in B\otimes A$ (1.5)
is satisfied, then ${{\mathbb{K}}}$ acts on tensor products in ${\operatorname{Mod}}(B)$, which is a right module category over the monoidal category ${\operatorname{Mod}}(A)$ owing to the coideal property.
Our approach to the support condition (1.5) is based on the following observation. We consider the automorphism $\xi=\operatorname{Ad}(K)^{-1}\circ\psi$ and the subalgebra
$B_{\xi}=\{x\in A\,|\,(\xi\otimes{\operatorname{id}})(\Delta(x))=\Delta(x)\}.$ (1.6)
It is easily checked that $B_{\xi}$ is the maximal coideal subalgebra contained in the fixed-point subalgebra $A^{\xi}$ and $B\subseteq B_{\xi}$ by (1.3). Moreover, it readily follows from (1.1) that ${{\mathbb{K}}}\in B_{\xi}\otimes A$. Hence, the support condition (1.5) is automatically satisfied whenever $B$ has the same maximality property of $B_{\xi}$, i.e.,
$B=B_{\xi}.$ (1.7)
1.4.
It is well-known that the universal R-matrix induces a braiding on ${\operatorname{Mod}}(A)$, and thereforee an action of the ordinary braid group on tensor products in ${\operatorname{Mod}}(A)$ In [Bro13, Kol20], the action of the tensor K-matrix is similarly encoded by a braiding on the module category ${\operatorname{Mod}}(B)$ over ${\operatorname{Mod}}(A)$, which induces an action of the cylindrical braid group [^2] Topologically, this is described by braids in the thick cylinder, see [Bri71, tD98].
Thus, a braided module category is to a braided monoidal category what the cylindrical braid group is to the ordinary braid group. However, even in this case, the setting is too restrictive to include the solutions of the generalized reflection equation (1.1), which arise from quantum symmetric pairs of arbitrary type. To remedy this, we introduce a new categorical structure, which we refer to as a boundary bimodule category, mirroring the algebraic framework described above. Interestingly, due to the presence of the twisting automorphism $\psi$, the symmetry group is replaced by a cylindrical braid groupoid, generated by the operators involved in (1.1), see Appendix B.
1.5.
Our framework naturally applies to quantum symmetric pairs of arbitrary Kac-Moody type. More precisely, let ${\mathfrak{g}}$ be an arbitray symmetrizable Kac-Moody algebra. Then, $A=U_{q}({\mathfrak{g}})$ is the corresponding quantized universal enveloping algebra (Drinfeld-Jimbo quantum group), while $B=U_{q}(\mathfrak{k})$ is a q-deformation of the universal enveloping algebra of a (pseudo-)fixed-point subalgebra of ${\mathfrak{g}}$, defined as a right coideal subalgebra of $U_{q}({\mathfrak{g}})$ in [Let02, Kol14, RV20, AV22a]. We will refer to it as a QSP subalgebra. In this work we will show (1.5) via (1.7), suitably interpreted in terms of completions with respect to certain subcategories of ${\operatorname{Mod}}(A)$ and ${\operatorname{Mod}}(B)$, see Theorem 4.1.2. The strategy is to study the classical limit of the basic universal K-matrix and of the inclusion $B\subseteq B_{\xi}$, relying on another maximality property of $U_{q}(\mathfrak{k})$ due to Kolb, see [Kol14], and the Iwasawa decomposition of ${\mathfrak{g}}$ with respect to ${\mathfrak{k}}$, see [RV22].
The relevant subcategories are defined in Section 2.3 for $U_{q}({\mathfrak{g}})$. In Section 5.2 we define a full subcategory $\mathcal{W}_{\theta}\subset{\operatorname{Mod}}(U_{q}(\mathfrak{k}))$ whose objects are called QSP weight modules. This category is significantly larger than the corresponding category (for quantum symmetric pairs of finite type) studied in [Kol20]. Indeed, the inclusion of the twist automorphism $\psi$, which is always close to a bialgebra isomorphism from $A$ to the co-opposite bialgebra $A^{\sf cop}$, is the reason that we can work with a larger subcategory of ${\operatorname{Mod}}(U_{q}(\mathfrak{k}))$, considering the formula for ${{\mathbb{K}}}$ and the polarization of the universal R-matrix. Moreover, ${{\mathbb{K}}}$ naturally acts on any tensor product $M\otimes V$, where $M\in\mathcal{W}_{\theta}$ and $V\in{\mathcal{O}}_{\theta}$, i.e., $V$ is a $U_{q}({\mathfrak{g}})$-module in category ${\mathcal{O}}$ satisfying an integrability property determined by $\theta$  [^3] More precisely, $V\in{\mathcal{O}}_{\theta}$ if $V$ is a $U_{q}({\mathfrak{g}})$-module in category ${\mathcal{O}}$ whose restriction to $U_{q}({\mathfrak{g}}_{X})$ is integrable, where $(X,\tau)$ is the generalized Satake diagram associated to $U_{q}(\mathfrak{k})$. In the rest of the paper, ${\mathcal{O}}_{\theta}$ is denoted by ${\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}$, see Section 2.3.. More concisely, we prove that $\mathcal{W}_{\theta}$ is equipped with a canonical structure of a bimodule category over ${\mathcal{O}}_{\theta}$ and the action of ${{\mathbb{K}}}$ yields a boundary structure on the bimodule category $\mathcal{W}_{\theta}$.
In particular, even in finity type, we obtain a generalization of the theorem of Kolb from [Kol20] which describes a braided module structure on finite-dimensional $U_{q}(\mathfrak{k})$-modules when ${\mathfrak{g}}$ is finite-dimensional.
1.6.
If ${\mathfrak{g}}$ is a Kac-Moody algebra of affine type then $U_{q}({\mathfrak{g}}^{\prime})$ has a rich category of finite-dimensional modules [CP94] which is of relevance in quantum integrability. In this case $U_{q}(\mathfrak{k})$ embeds in to $U_{q}({\mathfrak{g}}^{\prime})$ and in Section 6 we extend our previous work [AV22b] and show that ${{\mathbb{K}}}$ has a well-defined action on tensor products $M\otimes V$ for any $M\in\mathcal{W}_{\theta}$ and finite-dimensional $U_{q}({\mathfrak{g}})$-module $V$. Consider the subcategory ${\mathcal{C}}_{\theta}\subset\mathcal{W}_{\theta}$ of finite-dimensional modules. Relying on (1.4), we get trigonometric tensor K-matrices in the case that $M\in{\mathcal{C}}_{\theta}$ is the restriction of a finite-dimensional $U_{q}({\mathfrak{g}})$-module and both $M$ and $V$ are irreducible.
More generally, trigonometric tensor K-matrices act in generically irreducible tensor product objects in ${\mathcal{C}}_{\theta}$. However, it is unclear whether there are in fact generically irreducible finite-dimensional $U_{q}(\mathfrak{k})$-modules other than those coming from restricting $U_{q}({\mathfrak{g}})$-modules, see [AV22b]. Indeed, when $U_{q}({\mathfrak{g}}^{\prime})=U_{q}(\widehat{\mathfrak{sl}}_{2})$ and $U_{q}(\mathfrak{k})$ is the q-Onsager algebra (the coideal deformation of the universal enveloping algebra of the fixed-point subalgebra of the Chevalley involution $x\mapsto-x^{\sf T}$, also see [BK05] and references therein), Ito and Terwilliger proved in [IT09] that the objects in ${\mathcal{C}}_{\theta}$ are all restrictions of $U_{q}({\mathfrak{g}}^{\prime})$-modules, so that at least in this case we do not need to rely on generic irreducibility of finite-dimensional tensor products.
Recently Baseilhac, Gainutdinov and Lemarthe [LBG23] independently proposed axiomatics for a (tentative) element ${{\mathbb{K}}}$ of $B\otimes A$, which we present in Section A.3. In the case that $U_{q}(\mathfrak{k})$ is the q-Onsager algebra, given the affine lift $\pi_{z}$ with formal parameter $z$ of any irreducible finite-dimensional type 1 representation of $U_{q}(\mathfrak{sl}_{2})$, in ibid. fusion relations of a certain explicit element of $B\otimes\operatorname{End}(V)[[z]]$ were studied, which is conjectured to be proportional to $({\operatorname{id}}\otimes\pi_{z})({{\mathbb{K}}})$. It is a special case of the results of Section 6 that the element ${{\mathbb{K}}}$ indeed exists and that $({\operatorname{id}}\otimes\pi_{z})({{\mathbb{K}}})$ is a well-defined $U_{q}(\mathfrak{k})$-valued power series, as in this case $\mathcal{W}_{\theta}$ simplifies to ${\operatorname{Mod}}(U_{q}(\mathfrak{k}))$.
1.7.
We rely on the construction of the tensor K-matrix to address two open problems in the theory of affine quantum symmetric pairs. Below, we provide a brief overview of the strategy, which will be the focus of two forthcoming works.
1.7.1. Boundary q-characters
The theory of q-characters was introduced in [FR99, FM01] as a tool to study the finite-dimensional representation theory of quantum affine algebras. Roughly, the theory is a universal version of the well-known formalism of commuting transfer matrices in quantum integrability. The idea is to consider the grading-shifted universal R-matrix and evaluate only one factor on a finite-dimensional representation $V$, obtaining a semi-universal R-matrix with one factor in $\operatorname{End}(V)$. By taking the trace, one obtains a formal series $t_{V}(z)$ valued in the quantum analogue of the standard Borel subalgebra. From the Yang-Baxter equation, it follows that $[t_{V}(z),t_{W}(z^{\prime})]=0$.
Relying on Sklyanin’s formalism of two-boundary transfer matrices [Skl88], it is natural to expect that a two-boundary theory of q-characters should appear from the theory of quantum affine symmetric pairs. Namely, we consider two distinct QSP subalgebras of the quantum affine algebra assigned to each boundary. Provided that the underlying diagram automorphism is the same, the interplay between their tensor K-matrices is expected to yield, through evaluation and trace, a family of commuting elements $b_{V}(z)$. As in the standard case, the assignment is expected to provide a morphism from the Grothendieck ring of finite-dimensional modules over the quantum affine algebra, containing information about the structure of the modules under restriction to a QSP subalgebra.
1.7.2. The K-matrix formalism for QSP subalgebras
In [FRT90], Faddeev, Resheti-khin, and Takhtajan provided an alternative approach to quantum groups, which is commonly referred to as the FRT (or R-matrix) formalism. This yields a new presentation given in terms of a fixed matrix solution of the Yang-Baxter equation. The construction applies widely to quantum groups of both finite and affine type. More precisely, recently, Gautam, Rupert, and Wendlandt provided in [GRW22, RW23] a unified approach to the R-matrix formalism for $U_{q}({\mathfrak{g}})$, when ${\mathfrak{g}}$ is finite-dimensional. They prove that (an extension of) $U_{q}({\mathfrak{g}})$ can be realized as an Hopf quotient of the FRT algebra $U_{R}$ arising from the evaluation of the universal R-matrix on a finite-dimensional irreducible representation. The R-matrix formalism was extended to the case of classical quantum affine algebras in [RSTS90]. Moreover, building on the work of Olshanskii on twisted Yangians [Ols92], Molev, Ragoucy, and Sorba provided in [MRS] an analogue K-matrix presentation for certain quantum symmetric pairs.
We expect that in both finite and affine type, the evaluation of the universal tensor K-matrix associated to the quantum symmetric pair subalgebra $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ identifies $U_{q}(\mathfrak{k})$ with a subalgebra in the corresponding FRT algebra, thus yielding a uniform K-matrix presentation of $U_{q}(\mathfrak{k})$.
1.8. Outline
In Section 2 we recall the definition and fundamental properties of a quantized universal enveloping algebra $U_{q}({\mathfrak{g}})$ of a Kac-Moody algebra ${\mathfrak{g}}$ and its universal R-matrix. Next, in Section 3 we recall the definition of the subalgebra $U_{q}(\mathfrak{k})$ and the corresponding universal K-matrix; note that we work in the general setting determined by pseudo-involutions of ${\mathfrak{g}}$. The quantum group analogue of the key identity (1.7) is established in Section 4, see Theorem 4.2, via a maximality property of $U_{q}(\mathfrak{k})$ and by taking a classical limit. To promote this to the existence of the tensor K-matrix ${{\mathbb{K}}}$, in Section 5 we describe the appropriate categories of $U_{q}(\mathfrak{k})$-modules and $U_{q}({\mathfrak{g}})$-modules and operators acting on them; the main result is Theorem 5.16. In Section 6 we give the tensor K-matrix version of our previous work [AV22b] on K-matrices for finite-dimensional modules of quantum loop algebras.
We describe the general tensor K-matrix formalism for quasitriangular bialgebras in Appendix A. In Appendix B the corresponding categorical-topological framework is developed, resulting in the notion of a boundary bimodule category, which gives rise to actions of cylindrical braid groupoids on tensor products.
2. Quantum Kac-Moody algebras
Throughout this paper we will make use of the following notation, for any $k$-linear space $V$:
$X^{\xi}=(\xi\otimes{\operatorname{id}})(X)\,,\qquad X^{\xi}_{21}=(X^{\xi})_{21}\,,\qquad\mbox{and}\qquad X^{\xi\xi}=(\xi\otimes\xi)(X)$
for any $X\in V\otimes V$ and $k$-linear map $\xi:V\to V$. Here the field $k$ should be clear from the context.
Also, given a lattice $\mathsf{\Lambda}$, we shall denote its non-negative part by $\mathsf{\Lambda}^{+}$.
2.1. Kac-Moody algebras
Let ${\mathfrak{g}}$ be the Kac-Moody algebra defined over [^4] Or over any algebraically closed field of characteristic zero. ${\mathbb{C}}$ corresponding to a symmetrizable generalized Cartan matrix $A=(a_{ij})_{i,j\in I}$ [Kac90]. Let ${\mathfrak{h}}\subset{\mathfrak{g}}$ be the Cartan subalgebra. Denoting the simple roots by $\alpha_{i}$ and the simple coroots by $h_{i}$ ($i\in I$), we consider the root and coroot lattices:
$\mathsf{Q}=\operatorname{Sp}_{\mathbb{Z}}\{\alpha_{i}\,|\,i\in I\}\subset{\mathfrak{h}}^{*},\qquad\mathsf{Q}^{\vee}=\operatorname{Sp}_{\mathbb{Z}}\{h_{i}\,|\,i\in I\}\subset{\mathfrak{h}}.$ (2.1)
Let ${\bm{\Phi}}\subset\mathbb{Q}$ be the root system and ${\bm{\Phi}}^{+}\subset\mathsf{Q}^{+}$ the positive subsystem. Finally, set ${\mathfrak{g}}^{\prime}=[{\mathfrak{g}},{\mathfrak{g}}]$ and ${\mathfrak{h}}^{\prime}={\mathfrak{h}}\cap{\mathfrak{g}}^{\prime}=\mathsf{Q}^{\vee}\otimes_{\mathbb{Z}}{\mathbb{C}}$.
[…]
$\mathsf{Q}^{\vee}_{\sf ext}=\mathsf{Q}^{\vee}\oplus\bigoplus_{k=1}^{\operatorname{cork}(A)}{\mathbb{Z}}d_{k}\subset{{\mathfrak{h}}}\qquad\mbox{and}\qquad\mathsf{P}=\{\lambda\in{{\mathfrak{h}}}^{*}\;|\;\lambda(\mathsf{Q}^{\vee}_{\sf ext})\subset{\mathbb{Z}}\}.$
[…]
2.2. Drinfeld-Jimbo quantum groups
We set ${\mathbb{F}}=\overline{{\mathbb{C}}(q)}$ where $q$ is an indeterminate555 The results of this paper are also valid for ${\mathbb{F}}={\mathbb{C}}$ and $q\in{\mathbb{C}}^{\times}$not a root of unity.. Fix non-negative integers $\{\epsilon_{i}\;|\;i\in I\}$ such that the matrix $(\epsilon_{i}a_{ij})_{i,j\in I}$ is symmetric and set $q_{i}=q^{\epsilon_{i}}$. The quantum Kac-Moody algebra [Dri86, Jim85] associated to ${{\mathfrak{g}}}$ is the unital associative algebra $U_{q}({\mathfrak{g}})$ defined over ${\mathbb{F}}$ with generators $E_{i}$ and $F_{i}$ ($i\in I$), and $K_{h}$ ($h\in\mathsf{Q}^{\vee}_{\sf ext}$) subject to the following defining relations:
$K_{h}K_{h^{\prime}}=K_{h+h^{\prime}},\qquad K_{0}=1\,,$ (2.2)
$K_{h}E_{i}=q^{\alpha_{i}(h)}E_{i}K_{h}\,,\qquad K_{h}F_{i}=q^{-\alpha_{i}(h)}F_{i}K_{h}\,,$ (2.3)
$[E_{i},F_{j}]=\delta_{ij}\frac{K_{i}-K_{i}^{-1}}{q_{i}-q_{i}^{-1}}\,,$ (2.4)
for any $i,j\in I$ and $h,h^{\prime}\in\mathsf{Q}^{\vee}_{\sf ext}$, where $K_{i}^{\pm 1}=K_{\pm\epsilon_{i}h_{i}}$, together with the quantum Serre relations, see e.g., [Lus94, 3.1.1 (e)]. We consider the Hopf algebra structure on $U_{q}({\mathfrak{g}})$ uniquely determined by the following coproduct formulae
$\Delta(E_{i})$ $=E_{i}\otimes 1+K_{i}\otimes E_{i},$ $\Delta(F_{i})$ $=F_{i}\otimes K_{i}^{-1}+1\otimes F_{i},$ $\Delta(K_{h})$ $=K_{h}\otimes K_{h},$ (2.5)
for any $i\in I$ and $h\in\mathsf{Q}^{\vee}_{\sf ext}$. Let $\omega:U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})^{\operatorname{cop}}$ be the Chevalley involution
$\omega(K_{h})=K_{-h},\quad\omega(E_{i})=-F_{i},\quad\omega(F_{i})=-E_{i}\,.$ (2.6)
for any $i\in I$ and $h\in\mathsf{Q}^{\vee}_{\sf ext}$.
We denote by $U_{q}({\mathfrak{h}}^{\prime})$, $U_{q}{({{\mathfrak{n}}}^{+})}$, and $U_{q}{({{\mathfrak{n}}}^{-})}$ the subalgebras generated by the elements $\{K_{i}^{\pm 1}\}_{i\in I}$, $\{E_{i}\}_{i\in I}$ and $\{F_{i}\}_{i\in I}$, respectively. We set $U_{q}({\mathfrak{g}}^{\prime})=U_{q}{({{\mathfrak{n}}}^{+})}\cdot U_{q}({\mathfrak{h}}^{\prime})\cdot U_{q}{({{\mathfrak{n}}}^{-})}$ and $U_{q}{({{\mathfrak{b}}}^{\pm})}=U_{q}{({{\mathfrak{n}}}^{\pm})}\cdot U_{q}({\mathfrak{h}})$, where $U_{q}({\mathfrak{h}})$ is the commutative subalgebra generated by $K_{h}$, $h\in\mathsf{Q}^{\vee}_{\sf ext}$. The adjoint action of the $K_{h}$ induces a grading by the monoids ±$\pm\mathsf{Q}^{+}$ on the subalgebras $U_{q}{({{\mathfrak{n}}}^{\pm})}$, respectively, and we denote by $U_{q}{({{\mathfrak{n}}}^{\pm})}_{\pm\beta}$ the component of degree ±$\pm\beta$, for $\beta\in\mathsf{Q}^{+}$.
2.3. Categories of $U_{q}({\mathfrak{g}})$-modules
If $V$ is an $U_{q}({\mathfrak{h}})$–module and $\mu\in\mathsf{P}$, we denote the corresponding weight space of $V$ by
$V_{\mu}=\{v\in V\,|\,\forall h\in\mathsf{Q}^{\vee}_{\sf ext},\,\,K_{h}\,v=q^{\mu(h)}v\}\,.$
A $U_{q}({\mathfrak{g}})$–module $V$ is
  1. (C1)
    a (type $\bf 1$) weight module if $V=\bigoplus_{\mu\in\mathsf{P}}V_{\mu}$;
  2. (C2)
    integrable if it is a weight module, and the elements $\{E_{i},F_{i}\}_{i\in I}$ act locally nilpotently666 This implies that $V$ is completely reducible as a (possibly infinite) direct sum of simple finite–dimensional modules over the subalgebra generated by $E_{i},K_{i}^{\pm 1},F_{i}$, for each $i\in I$.;
  3. (C3)
    in category ${\mathcal{O}}$ if it is a weight module, and the action of $U_{q}{({{\mathfrak{n}}}^{+})}{}$ is locally finite777 This implies in particular that, for any $v\in V$, $U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}v=0$ for all but finitely many $\beta\in{\mathsf{Q}}^{+}$, cf. [Lus94, Sec. 3.4.7]..
We denote by $\mathcal{W}$, ${\mathcal{O}}$, ${\mathcal{O}}_{{\operatorname{int}}}$ the full subcategories of weight, category ${\mathcal{O}}$, integrable category ${\mathcal{O}}$ modules in ${\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$, respectively. The category $\mathcal{W}$ is only tensor, while the category ${\mathcal{O}}$ is braided, see [Dri86, Lus94]. The category ${\mathcal{O}}_{{\operatorname{int}}}$ is a semisimple braided subcategory of ${\mathcal{O}}$, whose simple modules are classified by dominant integral weights [Lus94, Thm 6.2.2, Cor. 6.2.3].
Finally, let $Z\subseteq I$ be any subset and $U_{q}({\mathfrak{g}}_{Z})\subseteq U_{q}({\mathfrak{g}})$ the subalgebra generated by $\{E_{i},F_{i},K_{i}^{\pm 1}\}_{i\in Z}$. We denote by ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}\subseteq{\mathcal{O}}$ the full subcategory of modules whose restriction to $U_{q}({\mathfrak{g}}_{Z})$ is integrable. Note that
${\mathcal{O}}_{{{\emptyset}\negthinspace\operatorname{-int}}}={\mathcal{O}}\qquad\mbox{and}\qquad{\mathcal{O}}_{{{I}\negthinspace\operatorname{-int}}}={\mathcal{O}}_{{\operatorname{int}}}\,.$
2.4. Completions
Let $A$ be an algebra, ${\mathcal{C}}\subset{\operatorname{Mod}}(A)$ a (full) subcategory, and $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ the algebra of endomorphisms of the forgetful functor ${\mathsf{f}}_{{\mathcal{C}}}\colon{\mathcal{C}}\to\operatorname{Vect}$. By definition, an element of $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ is a collection
$\varphi=\{\varphi_{V}\}_{V\in{\mathbb{C}}}\in\prod_{V\in{\mathcal{C}}}\operatorname{End}(V)$
which is natural, i.e., such that $f\circ\varphi_{V}=\varphi_{W}\circ f$ for any $f:V\to W$ in ${\mathbb{C}}$. The action of $A$ on any $V\in{\mathcal{C}}$ yields a morphism of algebras $A\to\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$, and factors through the action of $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ on $V$. We refer to $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ as the completion of $A$ with respect to the category ${\mathcal{C}}$. If $A$ is a bialgebra and ${\mathcal{C}}$ is closed under tensor products, the completion $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ is naturally endowed with a cosimplicial structure, see, e.g., [Dav07, ATL19].
For any subcategory ${\mathcal{C}}^{\prime}\subset{\mathcal{C}}$, there is a canonical morphism
$\eta\colon\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})\to\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}^{\prime}})$ (2.7)
given by the restriction to ${\mathcal{C}}^{\prime}$. A lift of $x\in\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}^{\prime}})$ is any element in $\eta^{-1}(x)\subset\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$. Since $\eta$ is in general neither surjective nor injective, a lift may not exist or be unique. However, in the following, we shall often consider operators whose action is given in terms of elements of $A$. Hence, if it exists, the lift is canonical (see, e.g.,  Section 2.6). In these cases, we shall make no distinction between an element and its canonical lift.
2.5. Completions of $U_{q}({\mathfrak{g}})$
We consider various completions of $U_{q}({\mathfrak{g}})^{\otimes n}$. Namely, for any $Z\subseteq I$ and $n>0$, let ${\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{n}\colon({\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}})^{\boxtimes n}\to\operatorname{Vect}$ be the $n$-fold forgetful functor given by
${\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{n}(V_{1},\dots,V_{n})=V_{1}\otimes\cdots\otimes V_{n}\,.$
We set ${\mathsf{f}}^{n}={\mathsf{f}}_{\emptyset}^{n}$, ${\mathsf{f}}_{{{I}\negthinspace\operatorname{-int}}}^{n}={\mathsf{f}}_{{\operatorname{int}}}^{n}$, and ${\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}={\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{1}$. Similarly, let ${\mathsf{f}}_{\mathcal{W}}^{n}:\mathcal{W}^{\boxtimes n}\to\operatorname{Vect}$ be the $n$-fold forgetful functor with respect to the category of weight $U_{q}({\mathfrak{g}})$-modules. Note that the chain of inclusions ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}\subset\mathcal{W}\subset{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$ induces a chain of morphisms of algebras $U_{q}({\mathfrak{g}})^{\otimes n}\to\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{n})\to\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{n})$. By [ATL24, Thm. 3.1], $U_{q}({\mathfrak{g}})^{\otimes n}$ embeds in $\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{n})$ and therefore in $\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{n})$.
2.6. The R-matrix
By [Dri86, Lus94], the universal R-matrix of $U_{q}({\mathfrak{g}})$ is a distinguished invertible element ${R}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$, whose action induces a braiding on the category ${\mathcal{O}}$. More precisely, ${R}$ satisfies the intertwining identity in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$
$R\cdot\Delta(x)=\Delta^{\operatorname{op}}(x)\cdot R\,,$ (2.8)
where $x\in U_{q}({\mathfrak{g}})$, and the coproduct identities in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}\boxtimes{\mathsf{f}})$
$(\Delta\otimes{\operatorname{id}})(R)=R_{13}\cdot R_{23}\qquad\mbox{and}\qquad({\operatorname{id}}\otimes\Delta)(R)=R_{13}\cdot R_{12}\,,$ (2.9)
where $\Delta$ denotes the coproduct and $\Delta^{\operatorname{op}}=(12)\circ\Delta$. More precisely, the morphisms
$\Delta,\Delta^{\operatorname{op}}\colon\operatorname{End}({\mathsf{f}})\to\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$
are defined as follows. For any $f\in\operatorname{End}({\mathsf{f}})$, we set
$\Delta(f)_{VW}=f_{V\otimes W}\quad\mbox{and}\quad\Delta^{\operatorname{op}}(f)_{VW}=(1\,2)\circ f_{W\otimes V}\circ(1\,2)$
where $V,W\in{\mathcal{O}}$. The morphisms $\Delta\otimes{\operatorname{id}}$ and ${\operatorname{id}}\otimes\Delta$ are similarly defined. From (2.8) and (2.9) it readily follows that $R$ is a solution of the Yang-Baxter equation
$R_{12}\cdot R_{13}\cdot R_{23}=R_{23}\cdot R_{13}\cdot R_{12}\,.$ (2.10)
in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}\boxtimes{\mathsf{f}})$.
The universal R-matrix arises from the Drinfeld double construction of $U_{q}({\mathfrak{g}})$ as the canonical tensor of a Hopf pairing between $U_{q}{({\mathfrak{b}}^{-})}$ and $U_{q}{({\mathfrak{b}}^{+})}$. More explicitly,
$R=\operatorname{C}^{{\operatorname{id}}}\cdot\Theta$ (2.11)
where
  • $\operatorname{C}^{{\operatorname{id}}}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{2})$ is a diagonal operator, acting on a tensor product of weight spaces $V_{\lambda}\otimes W_{\mu}$ as multiplication by $q^{(\lambda,\mu)}$, see also Section 5.3 ;
  • $\Theta=\sum_{\beta\in\mathsf{Q}^{+}}\Theta_{\beta}\in U_{q}{({{\mathfrak{n}}}^{-})}\widehat{\otimes}U_{q}{({{\mathfrak{n}}}^{+})}$, where
    $U_{q}{({{\mathfrak{n}}}^{-})}\widehat{\otimes}U_{q}{({{\mathfrak{n}}}^{+})}=\prod_{\beta}U_{q}{({{\mathfrak{n}}}^{-})}_{-\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}$
    and $\Theta_{\beta}\in U_{q}{({{\mathfrak{n}}}^{-})}_{-\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}$ is the canonical element with respect to the Hopf pairing.
The element $\Theta$ is commonly referred to as the quasi R-matrix [Lus94, Ch. 4]. Relying on the condition (C3), every element in $U_{q}{({{\mathfrak{n}}}^{-})}\widehat{\otimes}U_{q}{({{\mathfrak{n}}}^{+})}$ has a well-defined and natural action on any tensor product of the form $V\otimes W$ with $V\in{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$ and $W\in{\mathcal{O}}$. More precisely, let $\widetilde{{\mathsf{f}}}\colon{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))\to\operatorname{Vect}$ be the forgetful functor. Following [ATL24, Sec. 3], one proves that the canonical map
[…]
is an embedding. In particular, $\Theta\in\operatorname{End}(\widetilde{{\mathsf{f}}}\boxtimes{\mathsf{f}})$. Since the action of $\operatorname{C}^{{\operatorname{id}}}$ is defined only on weight modules, it follows that $R\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}})$. Then, it defines a braiding on ${\mathcal{O}}$ through the morphism
[…]
induced by the inclusion ${\mathcal{O}}\subset\mathcal{W}$.
3. Quantum symmetric pairs
We recall the definition of quantum symmetric pairs for quantum Kac-Moody algebras following [Kol14, RV20, AV22a].
3.1. Pseudo-involutions
Let $\operatorname{Aut}({A})$ be the group of diagram automorphisms of the Cartan matrix, i.e., the group of permutations $\tau$ of $I$ such that $a_{ij}=a_{\tau(i)\tau(j)}$. Any $\tau\in\operatorname{Aut}(A)$ canonically defines an automorphism of ${{\mathfrak{g}}^{\prime}}$, given on the generators by $\tau(h_{i})=h_{\tau(i)}$, $\tau(e_{i})=e_{\tau(i)}$, and $\tau(f_{i})=f_{\tau(i)}$. In [KW92, Sec. 4.9], Kac and Wang describe a procedure (depending on the choice of the subspace ${\mathfrak{h}}^{\prime\prime}\subset{\mathfrak{h}}$) to extend a diagram automorphism from ${{\mathfrak{h}}^{\prime}}$ to ${{\mathfrak{h}}}$. Then, we assume that the basis $\{d_{1},\dots,d_{\operatorname{cork}(A)}\}$ of ${\mathfrak{h}}^{\prime\prime}$ is chosen in such a way that the Kac-Wang extension of $\tau$ on ${\mathfrak{h}}$ preserves the extended root lattice $\mathsf{Q}^{\vee}_{\sf ext}$ and, by duality, the weight lattice $\mathsf{P}$.
Remark 3.1.1.
The problem of the existence of a basis of ${\mathfrak{h}}^{\prime\prime}$ compatible with a given $\tau\in\operatorname{Aut}(A)$ is discussed in [AV22a, Prop. 3.2] in the case $\operatorname{cork}(A)=1$. If $A$ is of affine type, such a basis exists for all $\tau$, see also [Kol14, Prop. 2.12]. normal-$\triangledown$
Let $X\subset I$ be a subset of indices such that the corresponding Cartan matrix $A_{X}$ is of finite type. We denote by ${\mathfrak{g}}_{X}\subset{\mathfrak{g}}$ the corresponding semisimple finite-dimensional Lie subalgebra. Let $\tau$ be an involutive diagram automorphism stabilizing $X$ such that $\tau|_{X}=\mathsf{oi}_{X}$, the involutive diagram automorphism of $X$ induced by the action of the longest element $w_{X}$ of the Weyl group $W_{X}$ on $\mathsf{Q}_{X}$. We consider the Lie algebra automorphism of ${\mathfrak{g}}$ given by
$\theta=\operatorname{Ad}(w_{X})\circ\omega\circ\tau$ (3.1)
where $\omega$ denotes the Chevalley involution on ${\mathfrak{g}}$ and $\operatorname{Ad}(w)$ for $w\in W$ is the corresponding braid group action realized in terms of triple exponentials. Note that $\operatorname{Ad}(w_{X})^{2}$ acts on the root space ${\mathfrak{g}}_{\lambda}$ ($\lambda\in\mathsf{Q}$) as multiplication by $(-1)^{\lambda(2\rho^{\vee}_{X})}$, with $\rho^{\vee}_{X}$ the half-sum of the positive coroots associated to ${\mathfrak{g}}_{X}$. Then $\theta$ fixes ${\mathfrak{g}}_{X}$ pointwise and the restriction of $\theta$ to ${\mathfrak{h}}$ is involutive; we use the notation $\theta$ also to denote the map on ${\mathfrak{h}}^{*}$ dual to $\theta|_{\mathfrak{h}}$, so that $\theta\big{(}{\mathfrak{g}}_{\alpha}\big{)}={\mathfrak{g}}_{\theta(\alpha)}$ for all $\alpha\in{\bm{\Phi}}$. We call $\theta$ a pseudo-involution (of the second kind).
Remark 3.1.2.
We will use the subscript $\theta$ even if objects are explicitly defined in terms of $(X,\tau)$. Note that the datum $(X,\tau)$ can be recovered from $\theta$ since $X=\{i\in I\,|\,\theta(h_{i})=h_{i}\}$. normal-$\triangledown$
Any ${\bm{y}}\in({\mathbb{C}}^{\times})^{I}$ may be viewed as a multiplicative character ${\bm{y}}:\mathsf{Q}\to{\mathbb{C}}^{\times}$ according to ${\bm{y}}(\alpha_{i})=y_{i}$ for $i\in I$. In turn, it gives rise to a Lie algebra automorphism $\operatorname{Ad}(\bm{y})$ of ${\mathfrak{g}}$ which acts on the root space ${\mathfrak{g}}_{\alpha}$ as multiplication by $\bm{y}(\alpha)$. For ${\bm{y}}\in({\mathbb{C}}^{\times})^{I}$ such that $y_{i}=1$ if $i\in X$, consider the following modification of $\theta$, which coincides with $\theta$ on $\langle{\mathfrak{g}}_{X},{\mathfrak{h}}\rangle$:
$\phi=\theta\circ\operatorname{Ad}({\bm{y}})^{-1}.$ (3.2)
3.2. Pseudo-fixed-point subalgebras
Let ${\mathfrak{k}}\subset{\mathfrak{g}}$ be the Lie subalgebra generated by ${\mathfrak{n}}^{+}_{X}$, ${\mathfrak{h}}^{\theta}$ and the elements
$f_{i}+\phi(f_{i})=f_{i}+y_{i}\theta(f_{i})$
for $i\in I$. We set ${\mathfrak{k}}^{\prime}={\mathfrak{k}}\cap{\mathfrak{g}}^{\prime}$.
Henceforth we assume that $(X,\tau)$ is a generalized Satake diagram, i.e., for all $i,j\in I$ such that $\theta(\alpha_{i})=-\alpha_{i}-\alpha_{j}$ we have $a_{ij}\neq-1$. Further, we assume that, for all $i\in I$,
$y_{i}=y_{\tau(i)}\qquad\text{if}\qquad(\alpha_{i},\theta(\alpha_{i}))=0.$ (3.3)
In this case the Lie subalgebra ${\mathfrak{k}}$ is close to the fixed-point subalgebra of an involution in the sense that ${\mathfrak{k}}\cap{\mathfrak{h}}={\mathfrak{h}}^{\theta}$, see [RV20, RV22, AV22a] and we call it a pseudo-fixed-point subalgebra.
3.3. Involutions and fixed-point subalgebras
The genuinely involutive case is described by the following refinement, cf. [Kol14, Sec. 2.4].
Lemma 3.3.1.
The following conditions are equivalent:
  1. (1)
    $\phi$ is an involution ;
  2. (2)
    $y_{i}y_{\tau(i)}^{-1}=(-1)^{\alpha_{i}(2\rho^{\vee}_{X})}$ for all $i\in I\setminus X$  ;
  3. (3)
    ${\mathfrak{k}}$ is the fixed-point subalgebra ${\mathfrak{g}}^{\phi}$
Note that condition (2) above implies that $\alpha_{i}(\rho^{\vee}_{X})\in{\mathbb{Z}}$ whenever $i=\tau(i)\in I\backslash X$, i.e., $(X,\tau)$ is a Satake diagram as defined in [Kol14, Def. 2.3].
Proof of Lemma 3.3.1..
Since $\theta^{2}=\operatorname{Ad}(w_{X})^{2}$, by direct inspection, $\phi$ is an involutive automorphism if and only if (2) holds. In this case, [Kol14, Lem. 2.8] implies that ${\mathfrak{k}}={\mathfrak{g}}^{\phi}$. Conversely, if ${\mathfrak{k}}$ is contained in ${\mathfrak{g}}^{\phi}$ then we must have $\phi^{2}(f_{i})=f_{i}$ for all $i\in I$ so that $\phi^{2}$ fixes ${\mathfrak{n}}^{-}$ pointwise. Noting that $\phi|_{\mathfrak{h}}=-w_{X}\circ\tau$ is an involution, we deduce that $\operatorname{ad}(f_{i})$ annihilates $e_{j}-\phi^{2}(e_{j})\in{\mathfrak{n}}^{+}$ for all $i,j\in I$. Using [Kac90, Lemma 1.5] we conclude that $\phi^{2}(e_{j})=e_{j}$ for all $j\in I$. Hence $\phi$ is an involution. ∎
3.4. Quantum pseudo-involutions
We now describe a convenient algebra automorphism $\theta_{q}$ of $U_{q}({\mathfrak{g}})$ which quantizes $\theta$. This is obtained by choosing a lift for each of the three factors in $\theta$. First, we consider the standard Chevalley involution on $U_{q}({\mathfrak{g}})$ given by (2.6). The diagram automorphism $\tau$ extends to an automorphism of $U_{q}({\mathfrak{g}})$ given on the generators by $\tau(E_{i})=E_{\tau(i)}$, $\tau(F_{i})=F_{\tau(i)}$, and $\tau(K_{h})=K_{\tau(h)}$.
The action of the Weyl group operator $w_{X}\in W_{X}$ is lifted to $U_{q}({\mathfrak{g}})$ as follows. Let $S_{X}$ be the braid group operator on integrable category ${\mathcal{O}}$ modules corresponding to $w_{X}$ [Lus94, Sec. 5]. More precisely, given a reduced expression $s_{i_{1}}\cdots s_{i_{\ell}}$ of $w_{X}$ in terms of fundamental reflections, one sets $S_{X}=S_{i_{1}}\cdots S_{i_{\ell}}$, where $S_{j}=T^{\prime\prime}_{j,1}$ in the notation from [Lus94, 5.2.1]. It follows from the braid relations that $S_{X}$ is independent of the chosen reduced expression.
In [AV22a, Sec. 4.9], we introduced a Cartan correction of $S_{X}$ given by
$\mathcal{S}_{\theta}=\kappa_{\theta}\cdot S_{X}$ (3.4)
Here, $\kappa_{\theta}$ is the operator defined on any weight vector of weight $\lambda$ as multiplication by $q^{\left(\theta(\lambda),\lambda\right)/2+\left(\lambda,\rho_{X}\right)}$, with $\rho_{X}$ the half-sum of the positive roots in ${\bm{\Phi}}_{X}$, the root system of the finite-dimensional semisimple Lie algebra ${\mathfrak{g}}_{X}$. Thus, $\mathcal{S}_{\theta}$ can be thought of as an element of $\operatorname{End}({\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$.
Remark 3.4.1.
In [AV22a, Sec. 6.6], we also introduced a Cartan correction of the universal R-matrix $R_{X}$ of $U_{q}({\mathfrak{g}}_{X})$ given by
$R_{\theta}=\operatorname{C}^{\theta}\cdot\Theta_{X}\in\operatorname{End}({\mathsf{f}}^{2})$ (3.5)
where $\Theta_{X}$ is the quasi R-matrix of $U_{q}({\mathfrak{g}}_{X})$. Then, by [AV22a, Lemma 6.8],
$R_{\theta}=(\mathcal{S}_{\theta}\otimes\mathcal{S}_{\theta})^{-1}\Delta(\mathcal{S}_{\theta})$ (3.6)
in $\operatorname{End}({\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}}^{2})$. normal-$\triangledown$
The operator $\mathcal{S}_{\theta}$ yields an algebra automorphism $\mathcal{T}_{\theta}=\operatorname{Ad}{(\mathcal{S}_{\theta})}$ of $\operatorname{End}({\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$, which preserves $U_{q}({\mathfrak{g}})$ [AV22a, Lem. 4.3 and 5.2]. Set
$\theta_{q}=\mathcal{T}_{\theta}\circ\omega\circ\tau$ (3.7)
As an algebra automorphism of $U_{q}({\mathfrak{g}})$, $\theta_{q}$ is the identity on $U_{q}({\mathfrak{g}}_{X})$, $\theta_{q}(K_{h})=K_{\theta(h)}$ for all $h\in\mathsf{Q}^{\vee}_{\sf ext}$ and $\theta_{q}\big{(}U_{q}({\mathfrak{g}})_{\lambda}\big{)}=U_{q}({\mathfrak{g}})_{\theta(\lambda)}$ for all $\lambda\in\mathsf{Q}$, see, e.g., [AV22a, Sec. 6.7].
We consider the tuple $\bm{\gamma}\in({\mathbb{F}}^{\times})^{I}$ as a multiplicative character $\bm{\gamma}:\mathsf{Q}\to{\mathbb{F}}^{\times}$, yielding a Hopf algebra automorphism $\operatorname{Ad}(\bm{\gamma})$ of $U_{q}({\mathfrak{g}})$ which acts on the root space ${\mathfrak{g}}_{\alpha}$ as multiplication by $\bm{\gamma}(\alpha)$. In analogy with (3.2), we set
$\phi_{q}=\theta_{q}\circ\operatorname{Ad}(\bm{\gamma}^{-1})$ (3.8)
so that $\phi_{q}(F_{i})=\bm{\gamma}_{i}\theta_{q}(F_{i})$.
3.5. QSP subalgebras
Let $\bm{\Gamma}\subset({\mathbb{F}}^{\times})^{I}$ and $\bm{\Sigma}\subset{\mathbb{F}}^{I}$ be the parameter sets defined in [AV22a, Def. 6.11]. Set $U_{q}({\mathfrak{h}}^{\theta})=\langle K_{h}\;|\;h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}\rangle$. The QSP subalgebra associated to $\theta$ with parameters $(\bm{\gamma},\bm{\sigma})\in\bm{\Gamma}\times\bm{\Sigma}$ is the subalgebra $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ generated by $U_{q}{({{\mathfrak{n}}}^{+})}_{X}$, $U_{q}({\mathfrak{h}}^{\theta})$, and the elements $B_{i}$ ($i\in I$), given by
$B_{i}=\left\{\begin{array}[]{cl}F_{i}&\mbox{if }\,i\in X\,,\\ F_{i}+\phi_{q}(F_{i})+\bm{\sigma}_{i}\,K_{i}^{-1}&\mbox{if }\,i\not\in X\,.\end{array}\right.$ (3.9)
We set $U_{q}({\mathfrak{k}^{\prime}})=U_{q}(\mathfrak{k})\cap U_{q}({\mathfrak{g}}^{\prime})$. If $\operatorname{cork}(A)\leqslant 1$ then $U_{q}({\mathfrak{k}^{\prime}})=U_{q}(\mathfrak{k})$, see [AV22a, Sec. 6.2].
Remarks 3.5.1.
  1. (1)
    Up to reparametrization, the definition of $U_{q}(\mathfrak{k})$ coincides with that given in [Let02, Kol14], see [AV22a, Rmk. 6.13].
  2. (2)
    It is known that $U_{q}(\mathfrak{k})$ is a right coideal, i.e., $\Delta(U_{q}(\mathfrak{k}))\subset U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{g}})$, with the additional property $U_{q}(\mathfrak{k})\cap U_{q}({\mathfrak{h}})=U_{q}({\mathfrak{h}}^{\theta})$; $U_{q}({\mathfrak{k}^{\prime}})$ satisfies analogous properties with respect to $U_{q}({\mathfrak{g}}^{\prime})$ and $U_{q}(({\mathfrak{h}}^{\prime})^{\theta})$, see [Kol14, AV22a]. $\triangledown$
3.6. Auxiliary R-matrices
Fix an extension of $\bm{\gamma}:\mathsf{Q}\to{\mathbb{F}}^{\times}$ to a character $\bm{\gamma}:\mathsf{P}\to{\mathbb{F}}^{\times}$. We consider the diagonal operator $\bm{\gamma}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}})$ acting on any weight space $V_{\lambda}$ as multiplication by $\bm{\gamma}(\lambda)$. Moreover, one has $\Delta(\bm{\gamma})=\bm{\gamma}\otimes\bm{\gamma}$ in $\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}}_{\mathcal{W}})$. Finally, $\operatorname{Ad}(\bm{\gamma})$ is an algebra automorphism of $\operatorname{End}({\mathsf{f}}_{\mathcal{W}})$, which restricts to a Hopf algebra automorphism of $U_{q}({\mathfrak{g}})$.
For any $V\in{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$, and any algebra automorphism $\psi:U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})$, let ${V^{\psi}}=\psi^{*}(V)\in{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$ be the pullback of $V$ under $\psi$.
Lemma 3.6.1.
In the case $\psi=\operatorname{Ad}(\bm{\gamma})\circ\theta_{q}^{-1}=\phi_{q}^{-1}$ the following statements hold.
  1. (1)
    For any $V\in\mathcal{W}$, ${V^{\psi}}\in\mathcal{W}$.
  2. (2)
    For any $V\in{\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}{}$, ${V^{\psi}}$ is in the opposite parabolic category ${\mathcal{O}}$, i.e., ${V^{\psi}}$ is a weight module which is locally finite over $U_{q}{({{\mathfrak{n}}}^{-})}$ and integrable over $U_{q}({\mathfrak{g}}_{X})$.
  3. (3)
    For any $V,W\in{\mathcal{O}}$, the operator
    $J=(\bm{\gamma}\otimes\bm{\gamma})\cdot R_{\theta}\cdot\Delta(\bm{\gamma})^{-1}=R_{\theta}\in\operatorname{End}({\mathsf{f}}^{2})$
    yields an intertwiner
    $J_{VW}\colon(V\otimes W)^{\psi}\to{V^{\psi}}\otimes^{\operatorname{op}}{W^{\psi}}\,,$
    where $\otimes^{\operatorname{op}}$ indicates that the action of $U_{q}({\mathfrak{g}})$ is given by the opposite coproduct $\Delta^{\operatorname{op}}$.
  4. (4)
    For any $V,W\in{\mathcal{O}}$, the action of the universal R-matrix on ${V^{\psi}}\otimes{W^{\psi}}$ is well-defined and it is given by
    $R_{{V^{\psi}}{W^{\psi}}}=J_{VW}\circ(R_{21})_{VW}\circ(J_{21})_{VW}^{-1}\,,$ (3.10)
[…]
Proof.
(1) It is enough to observe that, for any $h\in\mathsf{Q}^{\vee}_{\sf ext}$, $\psi(K_{h})=K_{\theta(h)}$. Therefore, ${V^{\psi}}\in\mathcal{W}$.
(2) Since $\theta_{q}=\mathcal{T}_{\theta}\circ\omega\circ\tau$, it is enough to observe that the restriction under $\mathcal{T}_{\theta}$ preserves $X$-integrable category ${\mathcal{O}}$ modules, see e.g., [AV22a, Lemma 5.2].
(3) and (4) are proved in [AV22a, Prop. 8.7]. In particular, note that
$(\bm{\gamma}\otimes\bm{\gamma})\cdot R_{\theta}=R_{\theta}\cdot(\bm{\gamma}\otimes\bm{\gamma})=R_{\theta}\cdot\Delta(\bm{\gamma})$ (3.11)
since $R_{\theta}$ is a weight preserving operator and $\bm{\gamma}$ is grouplike. ∎
Remark 3.6.2.
The identity (3.10) is somewhat surprising. Namely, by (2), the operator $R_{{V^{\psi}}{W^{\psi}}}$ is well-defined whenever $V\in{\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}{}$, since the locally finite action of $U_{q}{({{\mathfrak{n}}}^{-})}$ guarantees the convergence of $\Theta$, see Section 2.6. Instead, it is not clear a priori why the operator is well-defined even for category ${\mathcal{O}}$ modules which are not $X$-integrable. normal-$\triangledown$
Proposition 3.6.3.
  1. (1)
    There is a well-defined operator ${R}^{\psi}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}})$ given by the collection of maps
    ${R}_{VW}^{\psi}={R}_{{V^{\psi}}W}\in\operatorname{End}(V\otimes W)$
    where $V\in\mathcal{W}$ and $W\in{\mathcal{O}}$.
  2. (2)
    There is a well-defined operator ${R}_{21}^{\psi}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}_{\mathcal{W}})$ given by the collection of maps
[…]
    where $V,W\in{\mathcal{O}}$.
  3. (3)
    There is a well-defined operator ${R}^{\psi\psi}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ given by the collection of maps
    ${R}_{VW}^{\psi\psi}={R}_{{V^{\psi}}{W^{\psi}}}\in\operatorname{End}(V\otimes W)$
    where $V,W\in{\mathcal{O}}$.
Proof.
(1) By 2.6, $R\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}})$. Thus, by Lemma 3.6.1 (1), it follows that ${R}_{{V^{\psi}}W}$ is well-defined and natural in both $V$ and $W$.
(2) As in (1), it is enough to observe that
[…]
and $R_{21}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}_{\mathcal{W}})$.
(3) It follows readily from Lemma 3.6.1 (3). ∎
3.7. The quasi K-matrix
The QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})$ gives rise to a family of basic K-matrices in $U_{q}({\mathfrak{g}})$, see [AV22a, Thms. 8.8-8.9]. The result stems from the construction of the quasi K-matrix due to Bao and Wang [BW18b], subsequently generalized by Balagović and Kolb in [BK19] to the symmetrizable Kac-Moody setting. In [AV22a] it was shown that no additional assumptions on the parameters $\bm{\gamma}$, $\bm{\sigma}$ are necessary.
Theorem 3.7.1.
[AV22a, Thms. 8.8-8.9] There exists a unique element $\Upsilon\in\operatorname{End}({\mathsf{f}})$ of the form
$\Upsilon=1+\sum_{\begin{subarray}{c}\mu\in(\mathsf{Q}^{+})^{-\theta}\\ \mu\neq 0\end{subarray}}\Upsilon_{\mu}\in\prod_{\mu\in\mathsf{Q}^{+}}U_{q}{({{\mathfrak{n}}}^{+})}_{\mu}$ (3.14)
which satisfies
  • the intertwining identity
    $K\cdot b=\psi(b)\cdot K\,$ (3.15)
    for $b\in U_{q}({\mathfrak{k}^{\prime}})$;
  • the coproduct identity
    $\Delta(K)=J^{-1}\cdot 1\otimes K\cdot{R}^{\psi}\cdot\Upsilon\otimes 1\,,$ (3.16)
  • Cherednik’s generalized reflection equation
    $R^{\psi\psi}_{21}\cdot(1\otimes K)\cdot R^{\psi}\cdot(K\otimes 1)=(K\otimes 1)\cdot R^{\psi}_{21}\cdot(1\otimes K)\cdot R\,,$ (3.17)
where
[…]
Remarks 3.7.2.
  1. (1)
    By the weight decomposition of $\Upsilon$, one has $u\Upsilon=\Upsilon u$ for any $u\in U_{q}({\mathfrak{h}}^{\theta})$. Thus, the identity (3.15) for $U_{q}({\mathfrak{k}^{\prime}})$ implies the same identity for $U_{q}(\mathfrak{k})$ which appears in [AV22a, Thms. 8.8-8.9].
  2. (2)
    By Proposition 3.6.3, the identities (3.16) and (3.17) are in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$. $\triangledown$
In [AV22a], we introduced the notion of a cylindrical structure, encoding the formal properties of the datum $(\psi,J,K)$, see Section A.2. Since the action of the quasi K-matrix is defined at the level of category ${\mathcal{O}}$ modules, the result above can be rephrased as follows.
Corollary 3.7.3.
The datum $(\psi,J,K)=(\phi_{q}^{-1},R_{\theta},\Upsilon)$ is a cylindrical structure on $U_{q}({\mathfrak{g}})$ defined over ${\mathcal{O}}$ with respect to the QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})$.
3.8. Gauged K-matrices
In [AV22a, AV22b], we described a procedure to obtain new universal solutions of the generalized reflection equation (3.17) by gauging the quasi K-matrix, see [AV22b, Cor. 3.6]
Proposition 3.8.1.
Fix a subset $Z\subseteq I$. Let ${\mathcal{G}}_{Z}$ be the group of invertible elements ${\bf g}\in\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}})$ such that $\operatorname{Ad}({\bf g})$ preserves $U_{q}({\mathfrak{g}})\subset\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}})$. For any ${\bf g}\in{\mathcal{G}}_{Z}$, the datum $(\psi,J,K)$ given by
[…]
satisfies the identities (3.15), (3.16), and (3.17). In particular, $(\psi,J,K)$ is a cylindrical structure on $U_{q}({\mathfrak{g}})$ defined over ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}{}$ with respect to the QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})$.
Proof.
It is enough to observe that the results of Lemma 3.6.1 and Proposition 3.6.3 remain valid for this choice of $\psi$ and $J$, up to replacing ${\mathsf{f}}$ with ${\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}$. ∎
4. The classical limit of universal K-matrices
In [Kol14] a characterization of $U_{q}({\mathfrak{k}^{\prime}})$ is given as a maximal subspace with a prescribed specialization. In this section we rely on this result to characterize $U_{q}({\mathfrak{k}^{\prime}})$ as the maximal right coideal subalgebra of $U_{q}({\mathfrak{g}}^{\prime})$ contained in a fixed-point subalgebra determined by the quasi K-matrix.
We denote by $\operatorname{cl}(x)$ the specialization of $x$ at $q=1$ whenever well-defined. Henceforth, we assume that the QSP parameters $(\bm{\gamma},\bm{\sigma})\in\bm{\Gamma}\times\bm{\Sigma}$ are specializable at $q=1$.
4.1. Maximality for QSP subalgebras
Note that in Section 3.2 we defined subalgebras ${\mathfrak{k}}\subseteq{\mathfrak{g}}$ and ${\mathfrak{k}}^{\prime}={\mathfrak{k}}\cap{\mathfrak{g}}^{\prime}$ depending on a tuple $\bm{y}\in({\mathbb{C}}^{\times})^{I}$ satisfying (3.3). We recall the following characterization of $U_{q}(\mathfrak{k})^{\prime}$.
Theorem 4.1.1.
[Kol14, RV20] The QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})\subset U_{q}({\mathfrak{g}}^{\prime})$ is a maximal subspace such that $\operatorname{cl}(U_{q}({\mathfrak{k}^{\prime}}))=U({\mathfrak{k}}^{\prime})$ with $\operatorname{cl}(\bm{\gamma})=\bm{y}$.
Proof.
In the involutive case, we refer to [AV22a, Rmk. 6.13] for the equality between $U_{q}({\mathfrak{k}^{\prime}})$ and Kolb’s subalgebra $B_{{\bf c},{\bf s}}$. In this case, the statement is proved in [Kol14, Thms. 10.8, 10.11]. The proofs of loc. cit. naturally extend to generalized Satake diagrams, as was pointed out in [RV20]. In the same way, the proofs are still valid for general specializable $\bm{\gamma}\in\Gamma$, the difference being that the procedure yields the universal enveloping algebra of a Lie subalgebra ${\mathfrak{k}}^{\prime}\subset{\mathfrak{g}}^{\prime}$ depending on $\bm{y}=\operatorname{cl}(\bm{\gamma})$. ∎
Let $(\psi,J,K)$ be as in Proposition 3.8.1. We consider the automorphism of the completion $\operatorname{End}({\mathsf{f}}_{{\operatorname{int}}})$ given by888 There is a priori no reason why $U_{q}({\mathfrak{g}})$ should be $\operatorname{Ad}(\Upsilon)$-stable. In fact, it is easy to show that this is false even for ${\mathfrak{g}}=\mathfrak{sl}_{2}$ (see [DK17, Lemma 3.8]).
$\xi=\operatorname{Ad}(K)^{-1}\circ\psi=\operatorname{Ad}(\Upsilon)^{-1}\circ\phi_{q}^{-1}\,,$ (4.1)
where we recall $\phi_{q}^{-1}=\operatorname{Ad}(\bm{\gamma})\circ\theta_{q}^{-1}$. Then, we consider
$B^{\prime}_{\xi}=\{x\in U_{q}({\mathfrak{g}}^{\prime})\,|\,(\xi\otimes{\operatorname{id}})(\Delta(x))=\Delta(x)\}\subseteq U_{q}({\mathfrak{g}}^{\prime})^{\xi}\,.$ (4.2)
Clearly, $B^{\prime}_{\xi}$ is a coideal subalgebra in $U_{q}({\mathfrak{g}}^{\prime})$. Moreover, since $U_{q}({\mathfrak{k}^{\prime}})$ is a right coideal subalgebra and $K$ satisfies the local intertwining identity (3.15), one has $U_{q}({\mathfrak{k}^{\prime}})\subseteq B^{\prime}_{\xi}\subset U_{q}({\mathfrak{g}}^{\prime})^{\xi}$.
Theorem 4.1.2.
$U_{q}({\mathfrak{k}^{\prime}})=B^{\prime}_{\xi}$.
The proof is carried out in Section 4.3. It relies on the classical limit of the inclusion $U_{q}({\mathfrak{k}^{\prime}})\subseteq B^{\prime}_{\xi}$ and the properties of the classical limit of $\Upsilon$.
4.2. Classical limit of the quasi K-matrix
We shall now study the classical limit of $\Upsilon$. Note that $\operatorname{cl}(\theta_{q})=\theta$. Since $\operatorname{Ad}(\bm{\gamma})\in\operatorname{Aut}_{\sf alg}(U_{q}({\mathfrak{g}}))$ only depends on the values $\bm{\gamma}_{i}$, not on the choice of extension of $\bm{\gamma}:\mathsf{Q}\to{\mathbb{F}}^{\times}$ to a group homomorphism with domain $\mathsf{P}$, we obtain $\operatorname{cl}(\phi_{q})=\phi$, defined by (3.2) in terms of $\bm{y}=\operatorname{cl}(\bm{\gamma})$.
Proposition 4.2.1.
  1. (1)
    There is a unique element $\mathbf{Y}=\sum_{\mu\in{\bm{\Phi}}^{+}}\mathbf{Y}_{\mu}\in\prod_{\mu\in{\bm{\Phi}}^{+}}{\mathfrak{n}}^{+}_{\mu}$ such that $\operatorname{cl}(\Upsilon)=\exp(\mathbf{Y})$.
  2. (2)
    For any $b\in{\mathfrak{k}}^{\prime}$, $\exp(\operatorname{ad}(\mathbf{Y}))(b)=\phi^{-1}(b)$.
  3. (3)
    Let $\alpha\in{\bm{\Phi}}^{+}$ such that $\theta(\alpha)\neq-\alpha$. Then $\mathbf{Y}_{\alpha}=0$.
Proof.
(1) The coproduct identity (3.16) reads
$\Delta(\Upsilon)={R}_{\theta}^{-1}\cdot 1\otimes\Upsilon\cdot{R}^{\phi_{q}^{-1}}\cdot\Upsilon\otimes 1.$ (4.3)
Since the classical limit of ${R}$ and ${R}_{\theta}$ is trivial, it follows that $\operatorname{cl}(\Upsilon)\in\prod_{\mu\geqslant 0}U({\mathfrak{n}}^{+})_{\mu}$ is a grouplike element with $\operatorname{cl}(\Upsilon)_{0}=1$. Thus its logarithm is well-defined and primitive.
(2) The intertwining identity (3.15), viz. $\Upsilon\cdot b=\phi_{q}^{-1}(b)\cdot\Upsilon$ for all $b\in U_{q}({\mathfrak{k}^{\prime}})$, yields
$\exp(\mathbf{Y})\cdot b\cdot\exp(\mathbf{Y})^{-1}=\phi^{-1}(b)\qquad\text{for all }b\in{\mathfrak{k}}^{\prime}$ (4.4)
from which the desired equation readily follows.
(3) First consider an arbitrary element $\lambda\in\mathsf{Q}^{+}\backslash(\mathsf{Q}^{+})^{-\theta}$. By Theorem 3.7.1, $\Upsilon_{\lambda}=0$ and therefore $\operatorname{cl}(\Upsilon)_{\lambda}=0$. Since $\operatorname{cl}(\Upsilon)=\exp(\mathbf{Y})$, it follows that
$\sum_{k\geq 1}\frac{1}{k!}\sum_{\begin{subarray}{c}\beta_{1},\ldots,\beta_{k}\in{\bm{\Phi}}^{+}\\ \beta_{1}+\ldots+\beta_{k}=\lambda\end{subarray}}\mathbf{Y}_{\beta_{1}}\cdots\mathbf{Y}_{\beta_{k}}=0\,.$ (4.5)
In particular, if $\alpha\in{\bm{\Phi}}^{+}$ is a simple root such that $\theta(\alpha)\neq-\alpha$, then for $\lambda=\alpha$ the identity (4.5) reduces to $\mathbf{Y}_{\alpha}=0$, as required. We now proceed by induction on ${\mathsf{ht}}(\alpha)$. Let $\alpha\in{\bm{\Phi}}^{+}$ be a positive root such that $\theta(\alpha)\neq-\alpha$ and ${\mathsf{ht}}(\alpha)>1$. For any decomposition $\alpha=\beta_{1}+\cdots+\beta_{k}$ with $k>1$ and $\beta_{1},\dots,\beta_{k}\in{\bm{\Phi}}^{+}$ we have $\theta(\beta_{j})\neq-\beta_{j}$ for some $j\in\{1,\ldots,k\}$. Since ${\mathsf{ht}}(\beta_{j})<{\mathsf{ht}}(\alpha)$, by the induction hypothesis one has $\mathbf{Y}_{\beta_{j}}=0$. Thus, for $\lambda=\alpha$ the identity (4.5) reduces to
$\mathbf{Y}_{\alpha}=-\sum_{k\geq 2}\frac{1}{k!}\sum_{\begin{subarray}{c}\beta_{1},\ldots,\beta_{k}\in{\bm{\Phi}}^{+}\\ \beta_{1}+\ldots+\beta_{k}=\alpha\end{subarray}}\mathbf{Y}_{\beta_{1}}\cdots\mathbf{Y}_{\beta_{k}}=0\,$ (4.6)
and the result follows. ∎
The following statement is a generalization, to symmetric pairs of non-diagonal type, of the fact that the classical limit of the quasi R-matrix is $1\in U({\mathfrak{g}}\oplus{\mathfrak{g}})$.
Proposition 4.2.2.
$\mathbf{Y}=0$ if and only if $\phi$ is an involution.
Proof.
If $\mathbf{Y}=0$ then (4.4) implies that ${\mathfrak{k}}^{\prime}$ is contained in the fixed-point subalgebra of $\phi$. Re-running the proof of Lemma 3.3.1, we deduce that $\phi$ is an involution.
Conversely, suppose $\phi$ is an involution. Now Lemma 3.3.1 and (4.4) imply that $[\exp(\mathbf{Y}),b]=0$ for all $b\in{\mathfrak{k}}^{\prime}$. Since ${\mathfrak{k}}^{\prime}$ contains the elements $f_{i}$ for all $i\in X$ and $f_{i}+\phi(f_{i})$ for all $i\in I\backslash X$, this in turn implies
$\sum_{k\geq 1}\frac{1}{k!}[\mathbf{Y}^{k},f_{i}]=\begin{cases}0&\text{if }i\in X,\\ \displaystyle\sum_{k\geq 1}\frac{1}{k!}[\phi(f_{i}),\mathbf{Y}^{k}]&\text{if }i\in I\backslash X,\end{cases}$ (4.7)
which induces a recursive relation among the components of $\mathbf{Y}$ with respect to the principal grading:
$\mathbf{Y}=\sum_{h\geq 1}\mathbf{Y}(h),\qquad\mathbf{Y}(h)\in\bigoplus_{\begin{subarray}{c}\lambda\in{\bm{\Phi}}^{+}\\ {\mathsf{ht}}(\lambda)=h\end{subarray}}{\mathfrak{n}}^{+}_{\lambda}\subseteq{\mathfrak{n}}^{+}.$ (4.8)
To set this up, for $\bm{\ell}=(\ell_{1},\ldots,\ell_{k})\in{\mathbb{Z}}_{>0}^{k}$ we denote $\mathbf{Y}(\bm{\ell})=\mathbf{Y}(\ell_{1})\cdots\mathbf{Y}(\ell_{k})$, so that $\mathbf{Y}^{k}=\sum_{\bm{\ell}\in{\mathbb{Z}}_{>0}^{k}}\mathbf{Y}(\bm{\ell})$. For $m\in{\mathbb{Z}}$ consider the set of ordered partitions of $m$ in $k$ parts:
${\mathcal{P}}_{k}(m)=\{\bm{\ell}=(\ell_{1},\ldots,\ell_{k})\in{\mathbb{Z}}_{>0}^{k}\,|\,\ell_{1}+\ldots+\ell_{k}=m\}.$
Finally, for $i\in I\backslash X$, note that $\phi(f_{i})$ is homogeneous of degree $m_{i}={\mathsf{ht}}(w_{X}(\alpha_{\tau(i)}))\in{\mathbb{Z}}_{>0}$. With this notation, the component of (4.7) of degree $m-1$ ($m\in{\mathbb{Z}}_{>0}$) is given by
$\sum_{k\geq 1}\frac{1}{k!}\sum_{\bm{\ell}\in{\mathcal{P}}_{k}(m)}[\mathbf{Y}(\bm{\ell}),f_{i}]=\begin{cases}0&\text{if }i\in X,\\ \displaystyle\sum_{k\geq 1}\sum_{\bm{\ell}\in{\mathcal{P}}_{k}(m-m_{i}-1)}\frac{1}{k!}[\phi(f_{i}),\mathbf{Y}(\bm{\ell})]&\text{if }i\in I\backslash X.\end{cases}$ (4.9)
We now show by induction that $\mathbf{Y}(m)=0$ for all $m\in{\mathbb{Z}}_{>0}$, yielding $\mathbf{Y}=0$ as desired. The case $m=1$ follows directly from (4.9): the left-hand side equals $[\mathbf{Y}(1),f_{i}]$; the right-hand side equals 0 since ${\mathcal{P}}_{k}(-m_{i})=\emptyset$ for all $k$. Hence we obtain $[\mathbf{Y}(1),f_{i}]=0$ for all $i\in I$ so that $\mathbf{Y}(1)=0$ by [Kac90, Lemma 1.5].
Now let $m\in{\mathbb{Z}}_{\geq 2}$ and suppose that $\mathbf{Y}(m^{\prime})=0$ for all $m^{\prime}\in\{1,\ldots,m-1\}$. Since for all $k>1$ the elements of ${\mathcal{P}}_{k}(m)$ are tuples whose entries are less than $m$, the induction hypothesis implies that the left-hand side of (4.9) equals its $k=1$ term, namely $[\mathbf{Y}(m),f_{i}]$. Also, for all $k\geq 1$ the elements of ${\mathcal{P}}_{k}(m-m_{i}-1)$ are tuples whose entries are less than $m$. Hence the right-hand side of (4.9) vanishes by the induction hypothesis. We obtain that $[\mathbf{Y}(m),f_{i}]=0$ for all $i\in I$ and by [Kac90, Lemma 1.5] we deduce $\mathbf{Y}(m)=0$, which completes the proof. ∎
4.3. Proof of Theorem 4.1.2
Since $B^{\prime}_{\xi}$ is a right coideal subalgebra which contains $U_{q}({\mathfrak{k}^{\prime}})$, it is enough to prove that $\operatorname{cl}(B^{\prime}_{\xi})=U({\mathfrak{k}}^{\prime})$. Indeed, in this case Kolb’s maximality theorem, see Theorem 4.1.1, implies $B^{\prime}_{\xi}=U_{q}({\mathfrak{k}^{\prime}})$. We start with the following basic result.
Lemma 4.3.1.
Let $H$ be a Hopf algebra generated over a field $k$ by primitive elements. Then every coideal subalgebra of $H$ is a Hopf subalgebra.
Proof.
Since $H$ is cocommutative, every coideal subalgebra is a subbialgebra. It remains to prove that every subbialgebra of $H$ is preserved by the antipode $S$, which is a consequence from the following argument (see e.g. [EE05, Proof of Prop. 4.2] and [FGB05, Prop. 5.2]). Let $\eta$, $\varepsilon$, $m^{(n)}$ and $\Delta^{(n)}$ denote the unit, counit, $n$-fold product and $n$-fold coproduct, respectively. With respect to the convolution product on $\operatorname{End}_{k}(H)$, the antipode is the inverse of ${\operatorname{id}}_{H}$, the neutral element being $\eta\circ\varepsilon$. Hence, whenever the right-hand side is well-defined, we have
$S=\sum_{n\geq 0}m^{(n)}\circ(\eta\circ\varepsilon-{\operatorname{id}}_{H})^{\otimes n}\circ\Delta^{(n)}.$
Since any product of primitive elements is annihilated by $(\eta\circ\epsilon-{\operatorname{id}}_{H})^{\otimes n}\circ\Delta^{(n)}\in\operatorname{Hom}_{k}(H,H^{\otimes n})$ for sufficiently large $n$, we can express the antipode of any element of $H$ in terms of its bialgebra structure. ∎
Therefore, the coideal subalgebra $\operatorname{cl}(B^{\prime}_{\xi})$ of $U({\mathfrak{g}}^{\prime})$ is a Hopf subalgebra. Thus, by [MM65, Thm. 5.18, Prop. 6.13], $\operatorname{cl}(B^{\prime}_{\xi})=U({\mathfrak{f}})$ for some Lie subalgebra ${\mathfrak{f}}\subseteq{\mathfrak{g}}^{\prime}$. Moreover, since $\operatorname{cl}(U_{q}({\mathfrak{k}^{\prime}}))=U({\mathfrak{k}}^{\prime})$ and $\operatorname{cl}((U_{q}({\mathfrak{g}}^{\prime}))^{\xi})=U({\mathfrak{g}}^{\prime})^{\operatorname{cl}(\xi)}$, we have
${\mathfrak{k}}^{\prime}\subseteq{\mathfrak{f}}\subseteq({\mathfrak{g}}^{\prime})^{\operatorname{cl}(\xi)}\,.$ (4.10)
Note that $\operatorname{cl}(\xi)=\operatorname{Ad}(\operatorname{cl}(\Upsilon))\circ\phi$. By Proposition 4.2.1, $\operatorname{cl}(\Upsilon)=\exp(\mathbf{Y})$ for some $\mathbf{Y}=\sum_{\mu\in{\bm{\Phi}}^{+}}\mathbf{Y}_{\mu}\in\prod_{\mu\in{\bm{\Phi}}^{+}}{\mathfrak{n}}^{+}_{\mu}$. Thus,
$({\mathfrak{g}}^{\prime})^{\operatorname{cl}(\xi)}=\{x\in{\mathfrak{g}}^{\prime}\,|\,\operatorname{Ad}(\operatorname{cl}(\Upsilon))(x)=\phi^{-1}(x)\}=\{x\in{\mathfrak{g}}^{\prime}\,|\,\exp(\operatorname{ad}(\mathbf{Y}))(x)=\phi^{-1}(x)\}\,.$ (4.11)
Lemma 4.3.2.
${\mathfrak{k}}={\mathfrak{g}}^{\operatorname{cl}(\xi)}$.
Proof.
Clearly, ${\mathfrak{k}}\subseteq{\mathfrak{g}}^{\operatorname{cl}(\xi)}$. Denote ${\mathfrak{n}}^{+}_{\theta}={\mathfrak{n}}^{+}\cap\theta({\mathfrak{n}}^{-})={\mathfrak{n}}^{+}\cap\phi({\mathfrak{n}}^{-})$. By [RV22, Cor. 3.10], one has the Iwasawa decomposition:
${\mathfrak{g}}={\mathfrak{k}}\oplus{\mathfrak{h}}^{-\theta}\oplus{\mathfrak{n}}^{+}_{\theta}$ (4.12)
(as linear spaces). Now suppose $x\in{\mathfrak{h}}^{-\theta}\oplus{\mathfrak{n}}^{+}_{\theta}$ is fixed by $\operatorname{cl}(\xi)$, i.e.,
$\exp(\operatorname{ad}(\mathbf{Y}))(x)=\phi^{-1}(x).$ (4.13)
Denote by ${\mathfrak{b}}^{\pm}={\mathfrak{h}}\oplus{\mathfrak{n}}^{\pm}$ the standard Borel subalgebras. Note that $\exp(\operatorname{ad}(\mathbf{Y}))(x)\in\prod_{\mu\in\mathsf{Q}^{+}}{\mathfrak{b}}^{+}_{\mu}$ whereas $\phi^{-1}(x)\in{\mathfrak{b}}^{-}$. Hence (4.13) implies $x\in{\mathfrak{h}}$, so that $x\in{\mathfrak{h}}^{-\theta}={\mathfrak{h}}^{-\phi}$. Rewriting (4.13) as $\exp(\mathbf{Y})x=-x\exp(\mathbf{Y})$ and comparing weight-0 components, we obtain $x=0$. Now $({\mathfrak{h}}^{-\theta}\oplus{\mathfrak{n}}^{+}_{\theta})^{\operatorname{cl}(\xi)}=\{0\}$ implies the desired result. ∎
Intersecting with ${\mathfrak{g}}^{\prime}$, we obtain ${\mathfrak{k}}^{\prime}=({\mathfrak{g}}^{\prime})^{\operatorname{cl}(\xi)}$ so that ${\mathfrak{f}}={\mathfrak{k}}^{\prime}$. Theorem 4.1.1 now implies Theorem 4.1.2.
Remark 4.3.3.
If $\phi$ is an involution, Lemma 4.3.2 follows immediately from Proposition 4.2.2. normal-$\triangledown$
5. Tensor K-matrices
5.1. The QSP weight lattice $\mathsf{P}_{\theta}$
The involution $\theta:{\mathfrak{h}}\to{\mathfrak{h}}$ determines several lattices in ${\mathfrak{h}}$ and ${\mathfrak{h}}^{*}$ equipped with a natural pairing, see e.g.,  [BW18a, Wa21]. Adopting a slightly different approach, we consider the quotient lattice
$\mathsf{P}_{\theta}=\mathsf{P}/\mathsf{P}^{-\theta},$ (5.1)
with canonical projection $[\cdot]_{\theta}:\mathsf{P}\twoheadrightarrow\mathsf{P}_{\theta}$. The fixed-point sublattice of the ${\mathbb{Z}}$-linear automorphism $-\theta$ satisfies
$\mathsf{P}^{-\theta}=\mathsf{P}\cap\operatorname{span}_{\tfrac{1}{2}{\mathbb{Z}}}\{\lambda-\theta(\lambda)\,|\,\lambda\in\mathsf{P}\}.$ (5.2)
Given the bilinear pairing $\mathsf{P}\times\mathsf{Q}^{\vee}_{\sf ext}\to{\mathbb{Z}}$, define a pairing $\mathsf{P}_{\theta}\times\mathsf{Q}^{\vee}_{\sf ext}\to\tfrac{1}{2}{\mathbb{Z}}$ as follows: $\zeta(h)=\tfrac{1}{2}\lambda(h+\theta(h))$ for $\zeta=[\lambda]_{\theta}\in\mathsf{P}_{\theta}$ with $\lambda\in\mathsf{P}$ and $h\in\mathsf{Q}^{\vee}_{\sf ext}$ (note that it is independent of the choice of $\lambda$). This restricts to a nondegenerate pairing $\mathsf{P}_{\theta}\times(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}\to{\mathbb{Z}}$.
Remark 5.1.1.
Set $I_{\sf ns}=\{i\in I\,|\,\theta(\alpha_{i})=-\alpha_{i}\}$ so that $[\alpha_{i}]_{\theta}=0$ if and only if $i\in I_{\sf ns}$. Note that $\mathsf{P}_{\theta}$ identifies with a subgroup of the lattice $X^{\imath}$ considered in [Wa21, Sec. 3.2] of index $2^{|I_{\sf ns}|}$. normal-$\triangledown$
5.2. QSP weight modules
Recall that we have
$U_{q}({\mathfrak{h}}^{\theta})=\langle K_{h}\;|\;h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}\rangle=U_{q}(\mathfrak{k})\cap U_{q}({\mathfrak{h}}).$
Definition 5.2.1.
A $U_{q}(\mathfrak{k})$-module $M$ is a (type $\bf 1$) QSP weight module if, as a $U_{q}({\mathfrak{h}}^{\theta})$-module,
$M=\bigoplus_{\zeta\in\mathsf{P}_{\theta}}M_{\zeta}$ (5.3)
where
$M_{\zeta}=\{m\in M\,|\,\forall h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta},\,K_{h}\cdot m=q^{\zeta(h)}m\}\,.$ (5.4)
We denote the full subcategory of QSP weight modules by $\mathcal{W}_{\theta}\subseteq{\operatorname{Mod}}(U_{q}(\mathfrak{k}))$.
We summarize in the following some elementary properties of $\mathcal{W}_{\theta}$.
Proposition 5.2.2.
  1. (1)
    Let $M\in\mathcal{W}_{\theta}$. For any $\zeta\in\mathsf{P}_{\theta}$, $i\in I$ and $j\in X$, one has
[…]
  2. (2)
    $\mathcal{W}_{\theta}$ is a right module category over the monoidal category $\mathcal{W}$. Further, every weight $U_{q}({\mathfrak{g}})$-module is a QSP weight module under restriction to $U_{q}(\mathfrak{k})$.
  3. (3)
    The action of $U_{q}(\mathfrak{k})$ on $\mathcal{W}_{\theta}$ is faithful, i.e., $U_{q}(\mathfrak{k})$ embeds as a subalgebra into $\operatorname{End}({\mathsf{f}}_{\theta})$, where ${\mathsf{f}}_{\theta}:\mathcal{W}_{\theta}\to\operatorname{Vect}$ is the forgetful functor.
Proof.
(1) Note that $\bm{\sigma}_{i}=0$ if $\theta(\alpha_{i})\neq-\alpha_{i}$. Therefore, the result follows from the relations
$K_{h}B_{i}=q^{-\alpha_{i}(h)}B_{i}K_{h}\qquad\mbox{and}\qquad K_{h}E_{j}=q^{\alpha_{i}(h)}E_{j}K_{h}$
for any $i\in I$, $j\in X$, $h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}$.
(2) Since $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ is a (right) coideal subalgebra, ${\operatorname{Mod}}(U_{q}(\mathfrak{k}))$ is a right module category over ${\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$ with monoidal action given by the tensor product, cf. B.1. In particular, for any $M\in\mathcal{W}_{\theta}$ and $V\in\mathcal{W}$, one has $M\otimes V\in{\operatorname{Mod}}(U_{q}(\mathfrak{k}))$. Then,
$M\otimes V=\bigoplus_{\zeta\in\mathsf{P}_{\theta}}\underbrace{\left(\bigoplus_{\mu\in\mathsf{P}}M_{\zeta-[\mu]_{\theta}}\otimes V_{\mu}\right)}_{(M\otimes V)_{\zeta}}$
Let $h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}$ and $\zeta\in\mathsf{P}_{\theta}$. Then, $K_{h}$ acts on $(M\otimes V)_{\zeta}$ as multiplication by $q^{\zeta(h)}$. Namely, write $\zeta=[\lambda]_{\theta}$ for some $\lambda\in\mathsf{P}$. Then, it is enough to observe that $K_{h}$ is grouplike and it acts on $M_{\zeta-[\mu]_{\theta}}$ and $V_{\mu}$ via multiplication by $q^{(\lambda-\mu)(h)}$ and $q^{\mu(h)}$, respectively.
Finally, note that the restriction of $V$ to $U_{q}(\mathfrak{k})$ is given by the action of $V$ on the trivial representation $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})\stackrel{{\scriptstyle\varepsilon}}{{\longrightarrow}}{\mathbb{F}}$ , thus yielding a restriction functor $\mathcal{W}\to\mathcal{W}_{\theta}$.
(3) By (2), every module in ${\mathcal{O}}$ restricts to $\mathcal{W}_{\theta}$. This yields a morphism $\operatorname{End}({\mathsf{f}}_{\theta})\to\operatorname{End}({\mathsf{f}})$ and a commutative diagram
[…]
where the right vertical arrow is injective by [ATL24]. It follows that the map $U_{q}(\mathfrak{k})\to\operatorname{End}({\mathsf{f}}_{\theta})$ is also injective. ∎
Remark 5.2.3.
Consider the case of a split quantum symmetric pair, i.e., $\theta=\omega$. Then $\mathsf{P}_{\theta}=\{0\}$ and $\mathcal{W}_{\theta}={\operatorname{Mod}}(U_{q}(\mathfrak{k}))$. In particular, every $U_{q}(\mathfrak{k})$-module is a QSP weight module with a unique QSP weight space of weight 00. normal-$\triangledown$
5.3. Tensor weight operators
Let $f\colon\mathsf{P}\to{\mathbb{F}}$ be a ${\mathbb{Z}}$-linear functional which is self-adjoint with respect to the pairing on $\mathsf{P}$, i.e., $(f(\lambda),\mu)=(\lambda,f(\mu))$ for any $\lambda,\mu\in\mathsf{P}$. We consider the symmetric functional $\operatorname{C}^{f}:\mathsf{P}\times\mathsf{P}\to{\mathbb{F}}$ given by
$\operatorname{C}^{f}(\lambda,\mu)=q^{(f(\lambda),\mu)}\,.$ (5.6)
For any $V,W\in\mathcal{W}$, we denote by the same symbol the diagonal operator on $V\otimes W$, acting on the subspace $V_{\lambda}\otimes W_{\mu}$ as multiplication by $\operatorname{C}^{f}(\lambda,\mu)$. Let ${\mathsf{f}}_{\mathcal{W}}^{2}\colon\mathcal{W}\boxtimes\mathcal{W}\to\operatorname{Vect}$ be the 2-fold forgetful functor. Then, one readily checks that $\operatorname{C}^{f}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{2})$ and in $\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{3})$ the following identities hold:
$(\Delta\otimes{\operatorname{id}})(\operatorname{C}^{f})=\operatorname{C}^{f}_{13}\operatorname{C}^{f}_{23},\qquad({\operatorname{id}}\otimes\Delta))(\operatorname{C}^{f})=\operatorname{C}^{f}_{13}\operatorname{C}^{f}_{12}.$ (5.7)
5.4. Tensor QSP weight operators
Let $f\colon\mathsf{P}\to{\mathbb{F}}$ be a ${\mathbb{Z}}$-linear, self-adjoint functional such that $f\circ\theta=f$. Then, $f(\mathsf{P}^{-\theta})=0$ and $\operatorname{C}^{f}$ descends to a function $\mathsf{P}_{\theta}\times\mathsf{P}\to{\mathbb{F}}$. Proceeding as in Section 5.3, $\operatorname{C}^{f}$ determines a diagonal operator on any tensor product $M\otimes V$ with $M\in\mathcal{W}_{\theta}$ and $V\in\mathcal{W}$, which is natural in both $M$ and $V$. Let
${\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{\mathcal{W}}\colon\mathcal{W}_{\theta}\boxtimes\mathcal{W}\to\operatorname{Vect}$
be the 2-fold forgetful functor ${\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{\mathcal{W}}(M,V)=M\otimes V$. Then, $\operatorname{C}^{f}\in\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{\mathcal{W}})$.
5.5. Tensor Drinfeld algebras
For any $\beta\in\mathsf{Q}^{+}$, let ${\mathcal{B}}_{\beta}$ be a basis of $U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}$ and set ${\mathcal{B}}=\bigsqcup_{\beta\in\mathsf{Q}^{+}}{\mathcal{B}}_{\beta}$. Consider the Drinfeld algebra
${\mathcal{D}}=\left\{\left.\sum_{x\in{\mathcal{B}}}c_{x}x\,\right|\,c_{x}\in U_{q}{({\mathfrak{b}}^{-})}\right\}=\prod_{\beta\in\mathsf{Q}^{+}}U_{q}{({\mathfrak{b}}^{-})}U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\supseteq U_{q}({\mathfrak{g}})\,.$ (5.8)
One readily checks that ${\mathcal{D}}$ has a unique algebra structure which extends that of $U_{q}({\mathfrak{g}})$. Moreover, every element in ${\mathcal{D}}$ naturally acts on category ${\mathcal{O}}$ modules, since it reduces to a finite sum due to the condition (C3). By [ATL24, Sec. 3.1], ${\mathcal{D}}$ embeds as a subalgebra in both $\operatorname{End}({\mathsf{f}})$ and $\operatorname{End}({\mathsf{f}}_{{\operatorname{int}}})$, where ${\mathsf{f}}:{\mathcal{O}}\to\operatorname{Vect}$ and ${\mathsf{f}}_{{\operatorname{int}}}\colon{\mathcal{O}}_{{\operatorname{int}}}\to\operatorname{Vect}$ are the forgetful functors.
Remark 5.5.1.
By construction, the quasi K-matrix $\Upsilon$ is an element of ${\mathcal{D}}$, see Theorem 3.7.1. normal-$\triangledown$
In a similar way, we equip
${\mathcal{E}}_{\theta}$ =$=\left\{\left.\sum_{x\in{\mathcal{B}}}b_{x}\otimes c_{x}x\,\right|\,b_{x}\in U_{q}(\mathfrak{k})\,,\,c_{x}\in U_{q}{({\mathfrak{b}}^{-})}\right\}$ (5.9)
$=\prod_{\beta\in\mathsf{Q}^{+}}U_{q}(\mathfrak{k})\otimes U_{q}{({\mathfrak{b}}^{-})}U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\supseteq U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{g}})$ (5.10)
with the unique algebra structure extending that of $U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{g}})$. Every element in ${\mathcal{E}}_{\theta}$ acts in particular on tensor products of the form $M\otimes V$ with $M\in\mathcal{W}_{\theta}$ and $V\in{\mathcal{O}}$. The action is natural in $M$ and $V$, thus yielding an morphism of algebras ${\mathcal{E}}_{\theta}\to\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}})$. Proceeding as in [ATL24, Thm. 3.1], one further shows that the action of ${\mathcal{E}}_{\theta}$ is faithful on category ${\mathcal{O}}$ integrable $U_{q}({\mathfrak{g}})$-modules. Thus, we obtain a commutative diagram
[…]
where the first and second horizontal arrows are induced by the restriction functor ${\mathcal{O}}\to\mathcal{W}_{\theta}$ and the inclusion ${\mathcal{O}}_{{\operatorname{int}}}\hookrightarrow{\mathcal{O}}$, respectively.
Finally, set ${\mathcal{B}}_{X}=\bigsqcup_{\beta\in\mathsf{Q}_{X}^{+}}{\mathcal{B}}_{\beta}$. We consider the opposite algebra
${\mathcal{E}}^{\operatorname{op}}_{X}$ =$=\left\{\left.\sum_{x\in{\mathcal{B}}_{X}}a_{x}\otimes c_{x}\omega(x)\,\right|\,a_{x}\in U_{q}({\mathfrak{g}}_{X})\,,\,c_{x}\in U_{q}({\mathfrak{b}}_{X}^{+})\right\}$ (5.12)
$=\prod_{\beta\in\mathsf{Q}_{X}^{+}}U_{q}({\mathfrak{g}}_{X})\otimes U_{q}({\mathfrak{b}}_{X}^{+})U_{q}({\mathfrak{n}}_{X}^{-})_{-\beta}\supseteq U_{q}({\mathfrak{g}}_{X})\otimes U_{q}({\mathfrak{g}}_{X})$ (5.13)
We observe that ${\mathcal{E}}^{\operatorname{op}}_{X}$ acts on any tensor product $V\otimes W$ with $V\in{\operatorname{Mod}}(U_{q}({\mathfrak{g}}_{X}))$ and $W\in{\mathcal{O}}_{{\operatorname{int}}}(U_{q}({\mathfrak{g}}_{X}))$. Namely, by [Lus94, Thm. 6.2.2], a category ${\mathcal{O}}$ integrable module is completely reducible. Since $X\subset I$ is a subdiagram of finite type, $W$ decomposes as a (possibly infinite) sum of finite-dimensional irreducible $U_{q}({\mathfrak{g}}_{X})$-modules. Thus, the action of ${\mathcal{E}}^{\operatorname{op}}_{X}$ on $V\otimes W$ is well-defined. Through the restriction functor ${\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}\to{\mathcal{O}}_{{\operatorname{int}}}(U_{q}({\mathfrak{g}}_{X}))$ and the inclusion $U_{q}({\mathfrak{g}}_{X})\subset U_{q}(\mathfrak{k})$, we obtain an embedding
[…]
5.6. The tensor K-matrix
Let
$\psi=\operatorname{Ad}(\bm{\gamma})\circ\theta_{q}^{-1}=\phi_{q}^{-1}\,,\quad J=R_{\theta}\,,\quad\mbox{and}\quad K=\Upsilon\,,$ (5.15)
be the cylindrical structure on $U_{q}({\mathfrak{g}})$ with respect to the QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})$ constructed in Theorem 3.7.1. Let $R_{21}^{\psi}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ be the operator defined in Proposition 3.6.3 (2).
Theorem 5.6.1.
The operator
${{\mathbb{K}}}=R_{21}^{\psi}\cdot 1\otimes K\cdot R\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ (5.16)
has the following properties.
  1. (1)
    It satisfies the intertwining identity
    ${{\mathbb{K}}}\cdot\Delta(b)=({\operatorname{id}}\otimes\psi)(\Delta(b))\cdot{{\mathbb{K}}}$ (5.17)
    in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$, for all $b\in U_{q}(\mathfrak{k})$.
  2. (2)
    It satisfies the coproduct identities
    $(\Delta\otimes{\operatorname{id}})({{\mathbb{K}}})$ $=R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23}\,,$ (5.18)
    $({\operatorname{id}}\otimes\Delta)({{\mathbb{K}}})$ $=J^{-1}_{23}\cdot{{\mathbb{K}}}_{13}\cdot R^{\psi}_{23}\cdot{{\mathbb{K}}}_{12}\,$ (5.19)
    in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}\boxtimes{\mathsf{f}})$.
  3. (3)
    There are elements $\Xi_{X}\in{\mathcal{E}}^{\operatorname{op}}_{X}$ and $\Xi\in{\mathcal{E}}_{\theta}$ such that
    ${{\mathbb{K}}}=\operatorname{C}^{{\operatorname{id}}+\theta}\cdot\Xi_{X}\cdot\Xi\,.$
    In particular, ${{\mathbb{K}}}$ admits a canonical lift [^9] By construction, ${{\mathbb{K}}}$ acts on any tensor product $V\otimes W$ with $V,W\in{\mathcal{O}}$. This means that the action of ${{\mathbb{K}}}$ is extended to any tensor product $M\otimes W$ with $M\in\mathcal{W}_{\theta}$ but $W\in{\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}$. in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$, the identity (5.17) remains valid in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$, and the identities (5.18), (5.19) and (5.20) remain valid in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$101010 Since $\mathcal{W}_{\theta}$ is a module category over ${\mathcal{O}}$, the coproduct $\Delta$ naturally lifts to a morphism $\operatorname{End}({\mathsf{f}}_{\theta})\to\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}})$, cf. Section 2.6. .
  4. (4)
    It satisfies the generalized reflection equation
    $R^{\psi\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R^{\psi}_{23}\cdot{{\mathbb{K}}}_{12}={{\mathbb{K}}}_{12}\cdot R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23}\,.$ (5.20)
    in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$.
The proof is carried out in Sections 5.7 and 5.8.
Remark 5.6.2.
One has $K=(\varepsilon\otimes{\operatorname{id}})({{\mathbb{K}}})$. normal-$\triangledown$
5.7. Proof of Theorem 5.6.1 (1) and (2)
(1) For any $b\in U_{q}(\mathfrak{k})$ we have
$R^{\psi}_{21}\cdot 1\otimes\Upsilon\cdot R\cdot\Delta(b)$ $=R^{\psi}_{21}\cdot({\operatorname{id}}\otimes\psi)(\Delta^{\operatorname{op}}(b))\cdot 1\otimes\Upsilon\cdot R$ (5.21)
$=({\operatorname{id}}\otimes\psi)(R_{21}\cdot\Delta^{\operatorname{op}}(b))\cdot 1\otimes\Upsilon\cdot R$ (5.22)
$=({\operatorname{id}}\otimes\psi)(\Delta(b))\cdot R^{\psi}_{21}\cdot 1\otimes\Upsilon\cdot R,$ (5.23)
where the equality follows from the intertwining identity of $\Upsilon$ (3.15), the coideal property of $U_{q}(\mathfrak{k})$, and the intertwining property of $R$ (2.8).
(2) The first coproduct identity (5.18) follows from that of the R-matrix (2.9):
$(\Delta\otimes{\operatorname{id}})(R^{\psi}_{21}\cdot 1\otimes\Upsilon\cdot R)$ $=({\operatorname{id}}\otimes{\operatorname{id}}\otimes\psi)(R_{32}\cdot R_{31})\cdot(1\otimes 1\otimes\Upsilon)\cdot R_{13}\cdot R_{23}$ (5.24)
$=R^{\psi}_{32}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes\Upsilon)\cdot R_{13}\cdot R_{23}\,.$ (5.25)
On the other hand, the second coproduct identity (5.19) follows from that of $\Upsilon$ (4.3) and Lemma 3.6.1 (3). Namely, since $\Delta_{J}=\Delta^{\operatorname{op},\psi}$, one has
$({\operatorname{id}}\otimes\Delta)(R^{\psi}_{21})\cdot J^{-1}_{23}$ $=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\Delta_{J})(R^{\psi}_{21})$ (5.26)
$=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\Delta^{\operatorname{op},\psi})(R^{\psi}_{21})$ (5.27)
$=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\psi\otimes\psi)({\operatorname{id}}\otimes\Delta^{\operatorname{op}})(R_{21})$ (5.28)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot R^{\psi}_{21}\,.$ (5.29)
Therefore,
$({\operatorname{id}}\otimes\Delta)$ $(R^{\psi}_{21}\cdot 1\otimes\Upsilon\cdot R)=$ (5.30)
$=({\operatorname{id}}\otimes\Delta)(R^{\psi}_{21})\cdot J^{-1}_{23}\cdot(1\otimes 1\otimes\Upsilon)\cdot R^{\psi}_{23}\cdot(1\otimes\Upsilon\otimes 1)\cdot R_{13}\cdot R_{12}$ (5.31)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot R^{\psi}_{21}\cdot(1\otimes 1\otimes\Upsilon)\cdot R^{\psi}_{23}\cdot(1\otimes\Upsilon\otimes 1)\cdot R_{13}\cdot R_{12}$ (5.32)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes\Upsilon)\cdot R^{\psi}_{21}\cdot R^{\psi}_{23}\cdot R_{13}\cdot(1\otimes\Upsilon\otimes 1)\cdot R_{12}$ (5.33)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes\Upsilon)\cdot R_{13}\cdot R^{\psi}_{23}\cdot R^{\psi}_{21}\cdot(1\otimes\Upsilon\otimes 1)\cdot R_{12}$ (5.34)
where in the last equality follows from the Yang-Baxter equation.
5.8. Proof of Theorem 5.6.1 (3) and (4)
(3) We observe that, since $\theta|_{\mathfrak{h}}$ is involutive, the weight operator $\operatorname{C}^{{\operatorname{id}}+\theta}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}}_{\mathcal{W}})$ has a canonical lift in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{\mathcal{W}})$, see Section 5.4. We have to prove that
$\operatorname{C}^{-{\operatorname{id}}-\theta}\cdot\,{{\mathbb{K}}}=\Xi_{X}\cdot\Xi$
for some $\Xi_{X}\in{\mathcal{E}}^{\operatorname{op}}_{X}$ and $\Xi\in{\mathcal{E}}_{\theta}$.
The universal R-matrix is equipped with the factorization $R=\operatorname{C}^{{\operatorname{id}}}\cdot\,\Theta$, where
$\Theta=\sum_{\beta\in\mathsf{Q}^{+}}\Theta_{\beta}\in\prod_{\beta\in\mathsf{Q}^{+}}U_{q}{({{\mathfrak{n}}}^{-})}_{-\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\,$
see Section 2.6. By Lemma 3.6.1 (1), one readily checks that
$\operatorname{C}^{{\operatorname{id}}}_{V{W^{\psi}}}=\operatorname{C}^{\theta}_{VW}$
for any $V,W\in\mathcal{W}$. Hence, $R_{21}^{\psi}=\operatorname{C}^{\theta}\cdot\,\Theta_{21}^{\psi}$. Thus, we have
$\operatorname{C}^{-{\operatorname{id}}-\theta}\cdot\,{{\mathbb{K}}}=\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(\Theta^{\psi}_{21})\cdot\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(1\otimes\Upsilon)\cdot\Theta\,.$ (5.35)
Let ${R}_{X}$ be the universal R-matrix of $U_{q}({\mathfrak{g}}_{X})$ with the corresponding factorization ${R}_{X}=\operatorname{C}^{{\operatorname{id}}}\cdot\Theta_{X}$. By [AV22a, Prop. 4.3],
$\widetilde{\Theta}=\Theta_{X}^{-1}\cdot\Theta\in\prod_{\beta\in\mathsf{Q}^{+}\setminus\mathsf{Q}_{X}^{+}}U_{q}{({{\mathfrak{n}}}^{-})}_{-\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\,.$ (5.36)
Since $\psi$ is the identity on $U_{q}({\mathfrak{g}}_{X})$, we have $\Theta_{21}^{\psi}=\Theta_{X,21}\cdot\widetilde{\Theta}_{21}^{\psi}$. Set
$\Xi_{X}=\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(\Theta_{X,21})\quad\mbox{and}\quad\Xi=\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(\widetilde{\Theta}^{\psi}_{21})\cdot\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(1\otimes\Upsilon)\cdot\Theta\,.$ (5.37)
Clearly, $\Xi_{X}\in{\mathcal{E}}^{\operatorname{op}}_{X}$. It remains to prove that $\Xi\in{\mathcal{E}}_{\theta}$. Set
${\mathcal{D}}^{(2)}=\left\{\left.\sum_{x\in{\mathcal{B}}}a_{x}\otimes c_{x}x\,\right|\,a_{x}\in U_{q}({\mathfrak{g}})\,,\,c_{x}\in U_{q}{({\mathfrak{b}}^{-})}\right\}\supset{\mathcal{E}}_{\theta}\,.$
We claim that $\Xi\in{\mathcal{D}}^{(2)}$. Clearly, $\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(1\otimes\Upsilon)\cdot\Theta\in{\mathcal{D}}^{(2)}$. Then, it is enough to observe that $-\theta$ preserves $\mathsf{Q}^{+}\setminus\mathsf{Q}_{X}^{+}$. Therefore,
$\widetilde{\Theta}_{21}^{\psi}\in\prod_{\beta\in\mathsf{Q}^{+}\setminus\mathsf{Q}_{X}^{+}}U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\subset{\mathcal{D}}^{(2)}\,.$
Finally, set $\xi=\operatorname{Ad}(\Upsilon)^{-1}\circ\psi$ as in (4.1). By Theorem 4.1.2,
$U_{q}({\mathfrak{k}^{\prime}})=\{u\in U_{q}({\mathfrak{g}}^{\prime})\,|\,(\xi\otimes{\operatorname{id}})(\Delta(u))=\Delta(u)\}\,.$
Therefore, $\Xi\in{\mathcal{E}}_{\theta}$ if and only if
$(\xi\otimes{\operatorname{id}}\otimes{\operatorname{id}})\circ(\Delta\otimes{\operatorname{id}})(\Xi)=(\Delta\otimes{\operatorname{id}})(\Xi)\,.$ (5.38)
Note that (5.38) is automatically satisfied by $\Xi_{X}$, since $\xi$ is the identity on $U_{q}(\mathfrak{k})$. Moreover, it is also satisfied by $\operatorname{C}^{{\operatorname{id}}+\theta}$, since
  • the coproduct of $\operatorname{C}^{{\operatorname{id}}+\theta}$ is provided by (5.7);
  • for any $V,W\in\mathcal{W}$, one has $\operatorname{C}^{{\operatorname{id}}+\theta}_{{V^{\psi}}W}=\operatorname{C}^{{\operatorname{id}}+\theta}_{VW}$;
  • $\operatorname{C}^{{\operatorname{id}}+\theta}$ commutes with $\Upsilon$ because $\Upsilon_{\mu}=0$ whenever $\mu\not\in(\mathsf{Q}^{+})^{-\theta}$.
Thus, $\Xi\in{\mathcal{E}}_{\theta}$ if and only if ${{\mathbb{K}}}$ satisfies (5.38). Finally, it is enough to observe that, by (5.18), the identity (5.38) for ${{\mathbb{K}}}$ follows from the generalized reflection equation (3.17) for $\Upsilon$. Namely,
$(\xi\otimes{\operatorname{id}}\otimes{\operatorname{id}})$ $\circ(\Delta\otimes{\operatorname{id}})({{\mathbb{K}}})=$ (5.39)
=(5.18)${{=}}(\xi\otimes{\operatorname{id}}\otimes{\operatorname{id}})(R^{\psi}_{32}\cdot R^{\psi}_{31}\cdot 1\otimes 1\otimes K\cdot R_{13}\cdot R_{23})$ ) (5.40)
=(4.1)${{=}}R^{\psi}_{32}\cdot K^{-1}\otimes 1\otimes 1\cdot R^{\psi\psi}_{31}\cdot 1\otimes 1\otimes K\cdot R^{\psi}_{13}\cdot K\otimes 1\otimes 1\cdot R_{23}$ (5.41)
=(3.17)${{=}}R^{\psi}_{32}\cdot R^{\psi}_{31}\cdot 1\otimes 1\otimes K\cdot R_{13}\cdot R_{23}$ (5.42)
=(5.18)${{=}}(\Delta\otimes{\operatorname{id}})({{\mathbb{K}}}).$ ) . (5.43)
(4) The generalized reflection equation for ${{\mathbb{K}}}$ (5.20) follows from (3), (5.17), and (5.18).
5.9. Boundary bimodule structure
In the Appendix, we describe a categorical framework which encodes the defining properties of the tensor K-matrix ${{\mathbb{K}}}$. In Section B.4, we introduce the notion of a boundary bimodule category, which is a generalization of that of braided module category, see e.g., [Enr07, Bro13, Kol20].
Corollary 5.9.1.
Set
[…]
and let $R_{21}^{\psi}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ be as in Proposition 3.6.3 (2).
  1. (1)
    The category $\mathcal{W}_{\theta}$ has the following natural bimodule structure over ${\mathcal{O}}$, cf. B.3.
    • The right action is given by $M\triangleleft V=M\otimes V$ for any $M\in\mathcal{W}_{\theta}$ and $V\in{\mathcal{O}}$.
    • The left action is given by $M\blacktriangleleft V=M\otimes V^{\psi}$ with the associativity constraint
      $M\blacktriangleleft(V\otimes W)\to(M\blacktriangleleft W)\blacktriangleleft V$
      induced by $J_{VW}$, where $M\in\mathcal{W}_{\theta}$ and $V,W\in{\mathcal{O}}$.
    • The commutativity constraint
      $(M\triangleleft V)\blacktriangleleft W\to(M\blacktriangleleft W)\triangleleft V$
      is induced by $R^{\psi}_{VW}$, where $M\in\mathcal{W}_{\theta}$ and $V,W\in{\mathcal{O}}$.
  2. (2)
    The tensor K-matrix ${{\mathbb{K}}}$ defined in Theorem 5.6.1 is a boundary structure on the bimodule category $\mathcal{W}_{\theta}$ over ${\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}$.
Proof.
The notion of a boundary structure on a bimodule category is tailored to encode the datum of the auxiliary R-matrix $R^{\psi}$, the cylindrical structure $(\psi,J,K)$, and the corresponding tensor K-matrix ${{\mathbb{K}}}$. ∎
5.10. Gauged tensor K-matrices
The construction of the tensor K-matrix is compatible with gauging. We have the following analogue of Proposition 3.8.1.
Proposition 5.10.1.
Fix a subset $Z\subseteq I$. Let ${\mathcal{G}}_{Z}$ be the group of invertible elements ${\bf g}\in\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}})$ such that $\operatorname{Ad}({\bf g})$ preserves $U_{q}({\mathfrak{g}})\subset\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}})$. For any ${\bf g}\in{\mathcal{G}}_{Z}$, let $(\psi,J,K)$ be the datum given by
[…]
and set
${{\mathbb{K}}}=R_{21}^{\psi}\cdot 1\otimes K\cdot R\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}})\,.$ (5.46)
  1. (1)
    ${{\mathbb{K}}}$ admits a canonical lift in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X\cup Z}\negthinspace\operatorname{-int}}})$ which satisfies the identities (5.17) in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X\cup Z}\negthinspace\operatorname{-int}}})$ and (5.18), (5.19) and (5.20) in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X\cup Z}\negthinspace\operatorname{-int}}}\boxtimes{\mathsf{f}}_{{{X\cup Z}\negthinspace\operatorname{-int}}})$.
  2. (2)
    Consider the bimodule structure on $\mathcal{W}_{\theta}$ over ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}$ induced by $(\psi,J)$ as in Corollary 5.9.1. Then, ${{\mathbb{K}}}$ is a boundary structure on $\mathcal{W}_{\theta}$ over ${\mathcal{O}}_{{{X\cup Z}\negthinspace\operatorname{-int}}}$.
Proof.
It is enough to observe that
${{\mathbb{K}}}=R_{21}^{\psi}\cdot 1\otimes K\cdot R=1\otimes{\bf g}\cdot\left(R_{21}^{\phi_{q}^{-1}}\cdot 1\otimes\Upsilon\cdot R\right)=1\otimes{\bf g}\cdot{{\mathbb{K}}}_{{\bf g}=1}\,,$ (5.47)
where ${{\mathbb{K}}}_{{\bf g}=1}=R_{21}^{\phi_{q}^{-1}}\cdot 1\otimes\Upsilon\cdot R$ is the tensor K-matrix from Theorem 5.6.1. ∎
Remark 5.10.2.
The gauge group ${\mathcal{G}}_{Z}$ acts simultaneously on both the bimodule and the boundary structures of $\mathcal{W}_{\theta}$ over ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}$.
6. Trigonometric tensor K-matrices
In [AV22b] we showed that, for ${\mathfrak{g}}$ of untwisted affine type and for suitable twist automorphisms $\psi$, the universal K-matrix $K$ gives rise to a formal Laurent series with a well-defined action on finite-dimensional modules of $U_{q}({\mathfrak{g}})$. Here we consider similar questions for the tensor K-matrix ${{\mathbb{K}}}$.
6.1. Quantum loop algebras and finite-dimensional modules
Let ${\mathfrak{g}}$ be an untwisted affine Kac-Moody algebra. By [Kac90], ${\mathfrak{g}}^{\prime}$ is a central extension of the loop algebra ${\mathfrak{L}}=\mathring{{\mathfrak{g}}}[t,t^{-1}]$, where $\mathring{{\mathfrak{g}}}={\mathfrak{g}}_{I\backslash\{0\}}\subset{\mathfrak{g}}$ is a distinguished finite-dimensional Lie subalgebra appearing where the node $0\in I$ has been chosen as in [Kac90]. Let ${\mathcal{C}}$ be the category of finite-dimensional (type $\bf 1$) $U_{q}({\mathfrak{g}}^{\prime})$-modules and ${\mathsf{f}}_{{\mathcal{C}}}^{n}\colon{\mathcal{C}}^{\boxtimes n}\to\operatorname{Vect}$ the corresponding $n$-fold forgetful functor.
Let $\delta=\sum_{i\in I}a_{i}\alpha_{i}$ be the minimal imaginary root, i.e., the unique element of ${\operatorname{Ker}}(A)\cap\mathsf{Q}^{+}$ with setwise coprime coefficients, and set $K_{\delta}=\prod_{i\in I}K_{i}^{a_{i}}\in U_{q}({\mathfrak{g}}^{\prime})$. The element $K_{\delta}$ is central and acts trivially on any module in ${\mathcal{C}}$. Therefore, ${\mathcal{C}}$ is equivalent to the category of finite-dimensional (type $\bf 1$) modules over the quotient Hopf algebra $U_{q}(\mathfrak{L})=U_{q}({\mathfrak{g}}^{\prime})/(K_{\delta}-1)$, which is commonly referred to as the quantum loop algebra.
6.2. Trigonometric R-matrices [Dri86, FR92, KS95]
Let $z$ be a formal variable and consider the $\tau$-minimal grading shift $\Sigma_{z}:U_{q}(\mathfrak{L})\to U_{q}(\mathfrak{L})[z,z^{-1}]$ defined by
$\Sigma_{z}(E_{i})=z^{s_{i}}E_{i}\,,\quad\Sigma_{z}(F_{i})=z^{-s_{i}}F_{i}\,\quad\mbox{and}\quad\Sigma_{z}(K_{i})=K_{i}$
where $s_{0}=s_{\tau(0)}=1$ and $s_{i}=0$ if $i\notin\{0,\tau(0)\}$. For any $U_{q}(\mathfrak{L})$-module $V$, we consider the modules $V(z)=V\otimes{\mathbb{F}}(z)$ and ${V}(\negthinspace(z)\negthinspace)=V\otimes{{\mathbb{F}}}(\negthinspace(z)\negthinspace)$ with action given by the pullback through $\Sigma_{z}$ and extension of scalars.
Let $R\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ be the universal R-matrix of $U_{q}({\mathfrak{g}})$, see Section 2.6. Then,
$R(z)=({\operatorname{id}}\otimes\Sigma_{z})(R)=(\Sigma_{z^{-1}}\otimes{\operatorname{id}})(R)\in{\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}[\negthinspace[z]\negthinspace]$ (6.1)
that is $R(z)$ is a formal power series in $z$ whose coefficients have a well-defined and natural action on any tensor product in ${\mathcal{C}}$. Moreover, $R(z)$ satisfies the spectral Yang-Baxter equation
$R(w)_{12}\cdot R(wz)_{13}\cdot R(z)_{23}=R(z)_{23}\cdot R(wz)_{13}\cdot R(w)_{12}$ (6.2)
in ${\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}^{3})}[\negthinspace[w,z]\negthinspace]$. Thus, for any $U,V,W\in{\mathcal{C}}$, one obtains a matrix-valued formal series satisfying (6.2) in ${\operatorname{End}(U\otimes V\otimes W)}[\negthinspace[w,z]\negthinspace]$.
By [KS95, Sec. 4.2] and [Cha02, Thm. 3], when $V,W\in{\mathcal{C}}$ are irreducible, the tensor product $V\otimes W$ is generically irreducible, i.e., $V\otimes{W}(\negthinspace(z)\negthinspace)$ remains irreducible as a module over ${U_{q}(\mathfrak{L})}(\negthinspace(z)\negthinspace)$. This yields a factorization, uniquely defined up to a factor in ${\mathbb{F}}^{\times}$, of the form
$R_{VW}(z)=f_{VW}(z)\cdot{\mathbf{R}}_{VW}(z)\,,$ (6.3)
where $f_{VW}(z)\in{{\mathbb{F}}}(\negthinspace(z)\negthinspace)$ and ${\mathbf{R}}_{VW}(z)\in\operatorname{End}(V\otimes W)[z]$ is a non-vanishing operator, commonly referred to as a trigonometric R-matrix.
Since irreducible modules are highest weight and the R-matrix is weight zero, it is natural to normalize ${\mathbf{R}}_{VW}(z)$ such that the tensor product of the highest weight vectors has eigenvalue $1$. This yields a non-vanishing operator $\mathbf{R}^{\operatorname{norm}}_{VW}(z)$ satisfying the unitarity relation
[…]
Moreover, if $V\otimes W(\zeta)$ is irreducible for some $\zeta\in{\mathbb{F}}^{\times}$, then $\mathbf{R}^{\operatorname{norm}}_{VW}(\zeta)$ is well-defined and invertible.
6.3. Quantum affine symmetric pairs
Let $\theta$ be an automorphism of the second kind of ${\mathfrak{g}}$ of the form (3.1), $\theta_{q}$ the corresponding automorphism of $U_{q}({\mathfrak{g}})$ defined in (3.7), and $U_{q}(\mathfrak{k})\subseteq U_{q}({\mathfrak{g}})$ the corresponding QSP subalgebra.
Since $\delta$ is fixed by both the Weyl group and the diagram automorphisms, one has $\theta(\delta)=-\delta$. In particular, $[\delta]_{\theta}=[0]_{\theta}$, where $[\cdot]_{\theta}$ is the projection to the quotient lattice $\mathsf{P}_{\theta}$, see (5.1). Therefore, $\theta_{q}$ descends to an automorphism of $U_{q}(\mathfrak{L})$, and $U_{q}(\mathfrak{k})$ embeds in $U_{q}(\mathfrak{L})$ as a coideal subalgebra. We obtain the following analogue of Proposition 5.2.2 (2).
Lemma 6.3.1.
$\mathcal{W}_{\theta}$ is a right module category over ${\mathcal{C}}$. Further, every finite-dimensional $U_{q}(\mathfrak{L})$-module is a QSP weight module under restriction to $U_{q}(\mathfrak{k})$.
6.4. Spectral tensor K-matrices
Let ${\mathcal{G}}_{\theta,\bm{\gamma}}$ be the set of gauge transformations of the form ${\bf g}=S_{Y}^{-1}S_{X}{\bm{\beta}}^{-1}$ where
  • $Y\subseteq I\backslash\{0,\tau(0)\}$;
  • ${\bm{\beta}}\colon\mathsf{P}\to{\mathbb{F}}^{\times}$ satisfies ${\bm{\beta}}(\delta)=\bm{\gamma}(\delta)$.
Note that the identity belongs to ${\mathcal{G}}_{\theta,\bm{\gamma}}$ if and only if $0\not\in X$. Let ${\bf g}\in{\mathcal{G}}_{\theta,\bm{\gamma}}$ and set
$\psi=\operatorname{Ad}({\bf g}\bm{\gamma})\circ\thinspace\theta_{q}^{-1}\,,\quad J=({\bf g}\otimes{\bf g})\cdot R_{\theta}\cdot\Delta({\bf g})^{-1}\quad\mbox{and}\quad K={\bf g}\cdot\Upsilon\,,$ (6.5)
cf. Proposition 5.10.1.
Lemma 6.4.1.
  1. (1)
    $\Sigma_{z}\circ\psi=\psi\circ\Sigma_{z^{-1}}$.
  2. (2)
    $(\Sigma_{z}\otimes{\operatorname{id}})(R^{\psi})=R(z)^{\psi}$
  3. (3)
    $\Sigma_{z}(J)=J$.
  4. (4)
    $K(z)=\Sigma_{z}(K)\in{\operatorname{End}(f)}(\negthinspace(z)\negthinspace)$. If $0\not\in X$, $K(z)\in{\operatorname{End}(f)}[\negthinspace[z]\negthinspace]$.
Proof.
  1. (1)
    It is enough to observe that $\Sigma_{z}$ is the $\tau$-minimal grading shift.
  2. (2)
    This follows from (1) and (6.1).
  3. (3)
    Since ${\bf g}=S_{Y}^{-1}S_{X}{\bm{\beta}}^{-1}$ and $R_{\theta}=(\mathcal{S}_{\theta}\otimes\mathcal{S}_{\theta})^{-1}\cdot\Delta(\mathcal{S}_{\theta})$ by (3.6), the twist $J=({\bf g}\otimes{\bf g})\cdot R_{\theta}\cdot\Delta({\bf g})^{-1}$ is given by a Cartan correction of the quasi R-matrix $\Theta_{Y}$ of $U_{q}({\mathfrak{g}}_{Y})$. The result follows, since $0,\tau(0)\not\in Y$.
  4. (4)
    This is readily verified by direct inspection, see e.g., [AV22b, Sec. 4]. ∎
Let $\Delta_{z}=({\operatorname{id}}\otimes\Sigma_{z})\circ\Delta$ be the shifted coproduct. We have the following.
Theorem 6.4.2.
Let ${{\mathbb{K}}}=R^{\psi}_{21}\cdot 1\otimes K\cdot R\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}_{X\cup Y})$ be the tensor K-matrix corresponding to ${\bf g}\in{\mathcal{G}}_{\theta,\bm{\gamma}}$ constructed in Theorem 5.6.1. Set
${{\mathbb{K}}}(z)=({\operatorname{id}}\otimes\Sigma_{z})({{\mathbb{K}}})=R(z)_{21}^{\psi}\cdot 1\otimes K(z)\cdot R(z)$ (6.6)
in ${\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z)\negthinspace)$ Then, ${{\mathbb{K}}}(z)$ has a canonical lift in ${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z)\negthinspace)$, which satisfies the following identities:
  • the intertwining identity
    ${{\mathbb{K}}}(z)\cdot\Delta_{z}(b)$ $=({\operatorname{id}}\otimes\psi)(\Delta_{z^{-1}}(b))\cdot{{\mathbb{K}}}(z)\qquad(b\in U_{q}(\mathfrak{k}))$ (6.7)
    in ${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z)\negthinspace)$;
  • the coproduct identities
    $(\Delta_{z_{1}}\otimes{\operatorname{id}})({{\mathbb{K}}}(z_{2}))$ $=R(z_{1}z_{2})^{\psi}_{32}\cdot{{\mathbb{K}}}(z_{2})_{13}\cdot R(\tfrac{z_{2}}{z_{1}})_{23},$ (6.8)
    $({\operatorname{id}}\otimes\Delta_{z_{2}/z_{1}})({{\mathbb{K}}}(z_{1}))$ $=J_{23}^{-1}\cdot{{\mathbb{K}}}(z_{2})_{13}\cdot R(z_{1}z_{2})^{\psi}_{23}\cdot{{\mathbb{K}}}(z_{1})_{12},$ (6.9)
    in ${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z_{1},z_{2}/z_{1})\negthinspace)$;
  • the generalized reflection equation
    $R(\tfrac{z_{2}}{z_{1}})^{\psi\psi}_{32}\cdot{{\mathbb{K}}}(z_{2})_{13}\cdot R(z_{1}z_{2})^{\psi}_{23}\cdot{{\mathbb{K}}}(z_{1})_{12}$ $={{\mathbb{K}}}(z_{1})_{12}\cdot R(z_{1}z_{2})^{\psi}_{32}\cdot{{\mathbb{K}}}(z_{2})_{13}\cdot R(\tfrac{z_{2}}{z_{1}})_{23}$ (6.10)
    in ${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z_{1},z_{2}/z_{1})\negthinspace)$.
Moreover, if $0\not\in X$, then ${{\mathbb{K}}}(z)\in{\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}[\negthinspace[z]\negthinspace]$.
Proof.
It is enough to observe that $S_{X},S_{Y}$ are well-defined operators on finite-dimensional $U_{q}(\mathfrak{L})$-modules, ${\bf g}$ acts on $V(z)$ through Laurent polynomials, and therefore ${{\mathbb{K}}}(z)\in{\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z)\negthinspace)$. Then, the result readily follows from Theorem 5.6.1. ∎
6.5. Trigonometric tensor K-matrices
Let ${\mathcal{C}}_{\theta}\subseteq\mathcal{W}_{\theta}$ be the full subcategory whose objects are finite-dimensional. The construction of the spectral tensor K-matrix from Theorem 6.4.2 immediately implies the following.
Lemma 6.5.1.
Let $M\in{\mathcal{C}}_{\theta}$ and $V\in{\mathcal{C}}$. Then, there exists a QSP intertwiner $\mathbb{K}^{\operatorname{trig}}_{MV}(z)\colon M\otimes V(z)\to M\otimes V^{\psi}(z^{-1})$.
Proof.
The proof is standard, see, e.g., [AV22b, Lemma 5.1.1]. ∎
At present, in general it is unclear if $\mathbb{K}^{\operatorname{trig}}_{MV}(z)$ is obtained by normalizing the action of ${{\mathbb{K}}}(z)$ on $M\otimes{V}(\negthinspace(z)\negthinspace)$. When $M$ and $V$ are irreducible, it would follow automatically if the tensor product $M\otimes V$ is (generically) irreducible over $U_{q}(\mathfrak{k})$, which is currently out of reach.
Nonetheless, by relying on the construction of the trigonometric K-matrices carried out in [AV22b], we can prove it when $M\in{\mathcal{C}}_{\theta}$ is the restriction of an irreducible $U_{q}(\mathfrak{L})$-module. Namely, in [AV22b, Thm. 5.1.2 (2)], we proved that, for any irreducible $V\in{\mathcal{C}}$, there exist a formal Laurent series $g_{V}(z)\in{{\mathbb{F}}}(\negthinspace(z)\negthinspace)$ and a non-vanishing operator-valued polynomial $\mathbf{K}_{V}(z)\in\operatorname{End}(V)[z]$, uniquely defined up to a scalar in ${\mathbb{F}}^{\times}$, such that
$K_{V}(z)=g_{V}(z)\cdot\mathbf{K}_{V}(z)\,.$ (6.11)
Then, we have the following.
Theorem 6.5.2.
Let $M\in{\mathcal{C}}_{\theta}$ be the restriction of an irreducible $U_{q}(\mathfrak{L})$-module. For any irreducible $V\in{\mathcal{C}}$, set
$\mathbb{K}^{\operatorname{trig}}_{MV}(z)=(\mathbf{R}^{\operatorname{norm}}_{V^{\psi}M}(z))_{21}\cdot({\operatorname{id}}_{M}\otimes\mathbf{K}_{V}(z))\cdot\mathbf{R}^{\operatorname{norm}}_{MV}(z)$ (6.12)
in $\operatorname{End}(M\otimes V)(z)$, where $\mathbf{R}^{\operatorname{norm}}(z)$ denotes the unitary R-matrix from Section 6.2.
  1. (1)
    The operator-valued rational function $\mathbb{K}^{\operatorname{trig}}_{MV}(z)$ is non-vanishing and there exist a formal Laurent series $h_{MV}(z)\in{{\mathbb{F}}}(\negthinspace(z)\negthinspace)$ such that
    ${{\mathbb{K}}}_{MV}(z)=h_{MV}(z)\cdot\mathbb{K}^{\operatorname{trig}}_{MV}(z)\,.$ (6.13)
  2. (2)
    For any irreducible $V,W\in{\mathcal{C}}$, the generalized spectral reflection equation hold
    ${\mathbf{R}^{\operatorname{norm}}_{{W^{\psi}}{V^{\psi}}}(\tfrac{z_{2}}{z_{1}})}_{32}\cdot\mathbb{K}^{\operatorname{trig}}_{MW}(z_{2})_{13}\cdot\mathbf{R}^{\operatorname{norm}}_{{V^{\psi}}W}(z_{1}z_{2})_{23}\cdot\mathbb{K}^{\operatorname{trig}}_{MV}(z_{1})_{12}=$ (6.14)
    $\qquad=\mathbb{K}^{\operatorname{trig}}_{MV}(z_{1})_{12}\cdot\mathbf{R}^{\operatorname{norm}}_{{W^{\psi}}V}(z_{1}z_{2})_{32}\cdot\mathbb{K}^{\operatorname{trig}}_{MW}(z_{2})_{13}\cdot\mathbf{R}^{\operatorname{norm}}_{VW}(\tfrac{z_{2}}{z_{1}})_{23}$
Proof.
The result follows immediately from (6.6) and [AV22b, Thm. 5.1.2]. ∎
Remarks 6.5.3.
  1. (1)
    Consider the q-Onsager algebra, which is the QSP subalgebra $U_{q}(\mathfrak{k})\subset U_{q}(\widehat{\mathfrak{sl}}_{2})$ with $\theta=\omega$, see [BK05]. In [IT09], Ito and Terwilliger proved that, up to isomorphism, every finite-dimensional irreducible module over $U_{q}(\mathfrak{k})$ is the restriction of a $U_{q}(\widehat{\mathfrak{sl}}_{2})$-module. This implies that, in this case, Theorem 6.5.2 is valid for every irreducible module in ${\mathcal{C}}_{\theta}$.
  2. (2)
    For more general modules in $M\in{\mathcal{C}}_{\theta}$, Theorem 6.5.2 follows at once whenever the space of QSP intertwiners
    $M\otimes{V}(\negthinspace(z)\negthinspace)\to M\otimes{V}(\negthinspace(z)\negthinspace)^{\psi}$
    is proved to be one-dimensional. $\triangledown$
6.6. Unitary tensor K-matrices
The construction of unitary K-matrices requires to consider only $V\in{\mathcal{C}}$ such that $V^{\psi^{2}}=({V^{\psi}})^{\psi}$ is isomorphic to $V$ as a $U_{q}(\mathfrak{k})$-module. This constraint yields no loss of generality. Choosing the gauge transformation ${\bf g}=\mathcal{S}_{\theta}\cdot\bm{\gamma}^{-1}\in{\mathcal{G}}_{\theta,\bm{\gamma}}$, we have $\psi=\omega\circ\tau$. In particular, $\psi$ is an involution so that $V^{\psi^{2}}=V$ for every $V\in{\mathcal{C}}$.
In [AV22b, Prop. 5.4.1], we proved that, in this setting, the operators $\mathbf{K}_{V}(z)$ and $\mathbf{K}_{V^{\psi}}(z)$ can be normalized to a pair of unitary K-matrices $\mathbf{K}^{\operatorname{norm}}_{V}(z)$ and $\mathbf{K}^{\operatorname{norm}}_{V^{\psi}}(z)$. This yields the following.
Proposition 6.6.1.
Let $M\in{\mathcal{C}}_{\theta}$ be the restriction of an irreducible $U_{q}(\mathfrak{L})$-module. For any irreducible $V\in{\mathcal{C}}$ such that $V^{\psi^{2}}\cong V$, set
$\mathbb{K}^{\operatorname{norm}}_{MV}(z)=(\mathbf{R}^{\operatorname{norm}}_{V^{\psi}M}(z))_{21}\cdot({\operatorname{id}}_{M}\otimes\mathbf{K}^{\operatorname{norm}}_{V}(z))\cdot\mathbf{R}^{\operatorname{norm}}_{MV}(z)$ (6.15)
in $\operatorname{End}(M\otimes V)(z)$. Then,
$\mathbb{K}^{\operatorname{norm}}_{MV}(z)^{-1}=\mathbb{K}^{\operatorname{norm}}_{MV^{\psi}}(z^{-1})\,.$ (6.16)
Moreover, if $V(\zeta)$ is $U_{q}(\mathfrak{k})$-irreducible and $M\otimes V(\zeta)$, $V^{\psi}(\zeta)\otimes M$ are $U_{q}(\mathfrak{L})$-irreducible for some $\zeta\in{\mathbb{F}}^{\times}$, then $\mathbb{K}^{\operatorname{norm}}_{MV}(z)$ is well-defined and invertible.
Remark 6.6.2.
As before, for more general modules in $M\in{\mathcal{C}}_{\theta}$, Proposition 6.6.1 follows at once whenever the space of QSP intertwiners
$M\otimes{V}(\negthinspace(z)\negthinspace)\to M\otimes{V}(\negthinspace(z)\negthinspace)^{\psi}$
is proved to be one-dimensional. normal-$\triangledown$
Appendix A Cylindrical and reflection bialgebras
We develop a generalization of the approach of [Kol20, Sec. 2] to a braided structure on the category of modules of certain coideal subalgebras of quasitriangular bialgebras. In part, this was developed independently in [LBG23, Sec. 2.3].
A.1. Quasitriangular bialgebras [Dri86]
Fix a base field ${\mathbb{F}}$. A quasitriangular bialgebra $(A,R)$ is a pair $(A,R)$ where $A$ is a bialgebra and $R\in A\otimes A$ is a universal R-matrix, i.e., an invertible element satisfying the intertwining identity
$R\cdot\Delta(x)=\Delta^{\operatorname{op}}(x)\cdot R\qquad(x\in A)$ (A.1)
and the coproduct identities
$(\Delta\otimes{\operatorname{id}})(R)=R_{13}\cdot R_{23}\qquad\mbox{and}\qquad({\operatorname{id}}\otimes\Delta)(R)=R_{13}\cdot R_{12}\,,$ (A.2)
where $\Delta$ denotes the coproduct and $\Delta^{\operatorname{op}}=(12)\circ\Delta$. Then $R$ is a solution of the Yang-Baxter equation
$R_{12}\cdot R_{13}\cdot R_{23}=R_{23}\cdot R_{13}\cdot R_{12}\,.$ (A.3)
Moreover, $(\varepsilon\otimes{\operatorname{id}})(R)=1=({\operatorname{id}}\otimes\varepsilon)(R)$, where $\varepsilon$ is the counit of $A$. Note also that the co-opposite $(A^{\operatorname{cop}},R_{21})$ is a quasitriangular bialgebra.
A quasitriangular bialgebra $(A,R)$ can be twisted as follows, also see [Dav07].
  • Let $J\in A\otimes A$ be a Drinfeld twist, i.e., an invertible element satisfying $(J\otimes 1)\cdot(\Delta\otimes{\operatorname{id}})(J)=(1\otimes J)\cdot({\operatorname{id}}\otimes\Delta)(J)$ and $(\varepsilon\otimes{\operatorname{id}})(J)=1=({\operatorname{id}}\otimes\varepsilon)(J)$. There is a quasitriangular bialgebra $(A_{J},R_{J})$ such that $A_{J}=A$ as an algebra, its coproduct is $\Delta_{J}=\operatorname{Ad}(J)\circ\Delta$, the counit is $\varepsilon$, and $R_{J}=J_{21}\cdot R\cdot J^{-1}$.
  • Let $\psi:A\to A$ be an algebra automorphism. There is a quasitriangular bialgebra $(A^{\psi},R^{\psi\psi})$ such that $A^{\psi}=A$ as an algebra, its coproduct is $\Delta^{\psi}=(\psi\otimes\psi)\circ\Delta\circ\psi^{-1}$, the counit is $\varepsilon^{\psi}=\varepsilon\circ\psi^{-1}$, and $R^{\psi\psi}=(\psi\otimes\psi)(R)$. By construction, $\psi$ is an isomorphism of quasitriangular bialgebras $(A,R)\to(A^{\psi},R^{\psi\psi})$.
If $A^{\operatorname{cop},\psi}=A_{J}$ and $R_{21}^{\psi\psi}=R_{J}$, the pair $(\psi,J)$ is referred to as a twist pair, see [AV22a].
Remark A.1.1.
Note that, up to right-multiplication by elements of the centralizer of $\Delta(A)$ in $A\otimes A$, $J$ is uniquely determined by $\psi$ via the constraint $(\Delta^{\operatorname{op}})^{\psi}=\Delta_{J}$, in the same way that $R$ is almost uniquely determined by (A.1). normal-$\triangledown$
A.2. Cylindrical bialgebras [AV22a]
A cylindrical bialgebra $(A,R,B,\psi,J,K)$ is the datum of
  • a quasitriangular bialgebra $(A,R)$;
  • a (right) coideal subalgebra $B\subseteq A$, i.e., $\Delta(B)\subseteq B\otimes A$;
  • a twist pair $(\psi,J)$;
  • a (basic) universal K-matrix $K\in A$ with respect to $(B,\psi,J)$, i.e., an invertible element satisfying the local intertwining identity
    $K\cdot b=\psi(b)\cdot K\qquad(b\in B)$ (A.4)
    and the coproduct identity
    $\Delta(K)=J^{-1}\cdot(1\otimes K)\cdot R^{\psi}\cdot(K\otimes 1)\,.$ (A.5)
Equivalently, we shall also say that $(\psi,J,K)$ is a cylindrical structure on $(A,R)$ with respect to $B$. It follows, see [AV22a, Prop. 2.4], from combining (A.1), (A.5) and the twist pair property that $K$ satisfies the generalized reflection equation:
$R^{\psi\psi}_{21}\cdot(1\otimes K)\cdot R^{\psi}\cdot(K\otimes 1)=(K\otimes 1)\cdot R^{\psi}_{21}\cdot(1\otimes K)\cdot R\,.$ (A.6)
Moreover, $\varepsilon(K)=1$. The notion of cylindrical bialgebra generalizes that of cylinder twist, due to tom Dieck and Häring-Oldenburg in [tD98, tDHO98], and that of cylinder-braided coideal subalgebra, due to Balagović and Kolb in [BK19, Def. 4.10].
It is beneficial to require an additional condition of a cylindrical structure.
Definition A.2.1.
Let $(\psi,J,K)$ be a cylindrical structure on $(A,R)$ with respect to $B$. If
$R^{\psi}_{21}\cdot 1\otimes K\cdot R\quad\in\quad B\otimes A$ (A.7)
then we call $(\psi,J,K)$ supported on $(A,R,B)$ or, if is it is clear what $(A,R)$ is, supported on $B$. normal-$\triangledown$
Note that (A.7) ties the objects $B$ and $K$ more closely together: without it, any cylindrical structure $(\psi,J,K)$ on $(A,R)$ with respect to $B$ would also be with respect to any coideal subalgebra $B^{\prime}$ of $A$ contained in $B$, see also [Kol20, Rmk. 2.11]. Further, we obtain a second derivation of (A.6): it follows directly from (A.7) and (A.4).
Finally, since $B$ is a coideal subalgebra, (A.7) permits the module category ${\operatorname{Mod}}(B)$ over the monoidal category ${\operatorname{Mod}}(A)$ to be endowed with a braided structure compatible with the braided structure on ${\operatorname{Mod}}(A)$. We develop the categorical framework in Appendix B. The key algebraic ingredient is the universal tensor K-matrix.
A.3. Tensor K-matrices and reflection bialgebras
In a similar spirit, we introduce the notion of a reflection bialgebra. This is the natural generalization (to arbitrary twist pairs) of the notions of reflection algebra, given by Enriquez in [Enr07, Def. 4.1], and quasitriangular comodule algebra, given by Kolb in [Kol20, Def. 2.7].
Definition A.3.1.
A reflection bialgebra consists of a tuple $(A,R,B,\psi,J)$ as in Section A.2 and a universal tensor K-matrix ${{\mathbb{K}}}$ with respect to $(B,\psi,J)$, i.e., an invertible element of $B\otimes A$ satisfying the local intertwining identity
${{\mathbb{K}}}\cdot\Delta(b)=({\operatorname{id}}\otimes\psi)(\Delta(b))\cdot{{\mathbb{K}}}\qquad(b\in B)$ (A.8)
and the coproduct identities
$(\Delta\otimes{\operatorname{id}})({{\mathbb{K}}})$ $=R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23}\,,$ (A.9)
$({\operatorname{id}}\otimes\Delta)({{\mathbb{K}}})$ $=J_{23}^{-1}\cdot{{\mathbb{K}}}_{13}\cdot R^{\psi}_{23}\cdot{{\mathbb{K}}}_{12}\,.$ (A.10)
Equivalently, we shall also say that $(\psi,J,{{\mathbb{K}}})$ is a reflection structure on $(A,R)$ with respect to $B$. normal-$\triangledown$
Note that by applying ${\operatorname{id}}\otimes{\operatorname{id}}\otimes\epsilon$ to (A.10) we obtain the normalization
$({\operatorname{id}}\otimes\varepsilon)({{\mathbb{K}}})=1.$
Similar to the situation for a cylindrical structure $(\psi,J,K)$ supported on $B$, one readily deduces the following reflection equation in $A^{\otimes 3}$:
$R^{\psi\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R^{\psi}_{23}\cdot{{\mathbb{K}}}_{12}={{\mathbb{K}}}_{12}\cdot R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23}.$ (A.11)
Indeed, one can combine ${{\mathbb{K}}}\in B\otimes A$ with (A.8) and (A.9) or use (A.1), (A.10) and the twist pair property.
We will see that a reflection structure is equivalent to a cylindrical structure supported on $B$. First of all, in analogy with (part of) [Kol20, Lem. 2.9] (cf. [Enr07, Rmk. 4.2]), reflection structures straightforwardly induce cylindrical structures.
Lemma A.3.2.
Let $(A,R,B,\psi,J,{{\mathbb{K}}})$ be a reflection bialgebra and set
$K=(\varepsilon\otimes{\operatorname{id}})({{\mathbb{K}}})\in A\,.$
Then $(A,R,B,\psi,J,K)$ is a cylindrical bialgebra, i.e., $K$ is a universal K-matrix with respect to $(B,\psi,J)$. Furthermore,
${{\mathbb{K}}}=R^{\psi}_{21}\cdot 1\otimes K\cdot R.$ (A.12)
Proof.
This follows from a straightforward computation, almost identical to [Kol20, Lem. 2.9]:
${{\mathbb{K}}}=(((\varepsilon\otimes{\operatorname{id}})\circ\Delta)\otimes{\operatorname{id}})({{\mathbb{K}}})=(\varepsilon\otimes{\operatorname{id}}\otimes{\operatorname{id}})(R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23})=R^{\psi}_{21}\cdot 1\otimes K\cdot R\,.$ (A.13)
Then (A.4) follows from (A.8), while (A.5) follows from (A.10). ∎
Remark A.3.3.
If $\psi$ is a bialgebra automorphism and $J=R_{21}^{-1}$, the setup of Kolb in [Kol20, Sec. 2] is recovered. Note that in Kolb’s approach the twist $\psi$ is removed from the axiomatics by replacing ${\operatorname{Mod}}(A)$ by an equivariantized category. However, we are mainly interested in the case where $A$ is a q-deformed Kac-Moody algebra $U_{q}({\mathfrak{g}})$, which is a reflection bialgebra up to completion with respect to a full subcategory ${\mathcal{C}}$ of ${\operatorname{Mod}}(A)$, see Theorem 5.6.1. Since ${\mathcal{C}}$ is not preserved by the pullback of $\psi$ (except when ${\mathfrak{g}}$ is of finite type, which is assumed in [Kol20]), equivariantization is not applicable in our setting. normal-$\triangledown$
Remark A.3.4.
The concept of a universal tensor K-matrix meaningfully generalizes to the case where $B$ is a right $A$-comodule algebra with coaction map [^11] In other words, $\Delta_{B}:B\to B\otimes A$ is linear such that $(\Delta_{B}\otimes{\operatorname{id}}_{A})\circ\Delta_{B}=({\operatorname{id}}_{B}\otimes\Delta)\circ\Delta_{B}$. $\Delta_{B}$. In order to formulate the generalized notion of universal tensor K-matrix, simply replace the coproduct $\Delta$ by $\Delta_{B}$ in axioms (A.8-A.9). The situation described above is recovered if we assume that $B$ is a subalgebra of $A$ and $\Delta_{B}=\Delta|_{B}$. More generally, if there exists an algebra map $\epsilon_{B}:B\to{\mathbb{F}}$ (i.e.,  $B$ is an augmented algebra) then $\widetilde{B}=((\epsilon_{B}\otimes{\operatorname{id}}_{A})\circ\Delta_{B})(B)$ is a right coideal subalgebra of $A$ and a universal tensor K-matrix ${{\mathbb{K}}}\in B\otimes A$ with respect to $(A,R,B,\psi,J)$ induces a cylindrical structure $(\psi,J,K)$ on $(A,R)$ with respect to $\widetilde{B}$ via $K=(\epsilon_{B}\otimes{\operatorname{id}})({{\mathbb{K}}})\in A$. It has the property $R^{\psi}_{21}\cdot 1\otimes K\cdot R\in B\otimes A$. A natural sufficient condition for this cylindrical structure to be supported on $\widetilde{B}$ is isomorphicity of $B$ and $\widetilde{B}$ as right $A$-comodule algebras, i.e., injectivity of $(\epsilon_{B}\otimes{\operatorname{id}}_{A})\circ\Delta_{B}$. normal-$\triangledown$
A.4. From K-matrices to tensor K-matrices
We keep following the approach of [Kol20] closely. It is natural to ask when a cylindrical bialgebra $(A,R,B,\psi,J,K)$ can be promoted to a reflection bialgebra $(A,R,B,\psi,J,{{\mathbb{K}}})$ with ${{\mathbb{K}}}=R^{\psi}_{21}\cdot 1\otimes K\cdot R$. From Lemma A.3.2 it follows that the support condition (A.7) must be satisfied. It turns out this is sufficient.
Proposition A.4.1.
Let $(A,R,B,\psi,J,K)$ be a cylindrical bialgebra and define ${{\mathbb{K}}}\in A\otimes A$ by (A.12).
  1. (1)
    The identities (A.8-A.10) are satisfied and $K=(\varepsilon\otimes{\operatorname{id}})({{\mathbb{K}}})$.
  2. (2)
    ${{\mathbb{K}}}$ is a universal tensor K-matrix with respect to $(B,\psi,J)$ (in other words, $(A,R,B,\psi,J,{{\mathbb{K}}})$ is a reflection bialgebra) if and only if (A.7) is satisfied.
Proof.
(1) For any $b\in B$ we have
$R^{\psi}_{21}\cdot 1\otimes K\cdot R\cdot\Delta(b)$ $=R^{\psi}_{21}\cdot({\operatorname{id}}\otimes\psi)(\Delta^{\operatorname{op}}(b))\cdot 1\otimes K\cdot R$ (A.14)
$=({\operatorname{id}}\otimes\psi)(R_{21}\cdot\Delta^{\operatorname{op}}(b))\cdot 1\otimes K\cdot R$ (A.15)
$=({\operatorname{id}}\otimes\psi)(\Delta(b))\cdot R^{\psi}_{21}\cdot 1\otimes K\cdot R,$ (A.16)
where we have used (A.1), (A.4) and the coideal property $\Delta(B)\subseteq B\otimes A$, thus establishing the local intertwining identity (A.8). The first coproduct identity (A.9) follows from the identities (A.2) for the R-matrix:
$(\Delta\otimes{\operatorname{id}})(R^{\psi}_{21}\cdot 1\otimes K\cdot R)$ $=({\operatorname{id}}\otimes{\operatorname{id}}\otimes\psi)(R_{32}\cdot R_{31})\cdot(1\otimes 1\otimes K)\cdot R_{13}\cdot R_{23}$ (A.17)
$=R^{\psi}_{32}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes K)\cdot R_{13}\cdot R_{23}\,.$ (A.18)
On the other hand, the second coproduct identity (A.10) follows from (A.5) and the fact that $(\psi,J)$ is a twist pair. Indeed, since $\Delta_{J}=\Delta^{\operatorname{op},\psi}$, one has
$({\operatorname{id}}\otimes\Delta)(R^{\psi}_{21})\cdot J^{-1}_{23}$ $=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\Delta_{J})(R^{\psi}_{21})$ (A.19)
$=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\Delta^{\operatorname{op},\psi})(R^{\psi}_{21})$ (A.20)
$=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\psi\otimes\psi)({\operatorname{id}}\otimes\Delta^{\operatorname{op}})(R_{21})$ (A.21)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot R^{\psi}_{21}\,,$ (A.22)
therefore
$({\operatorname{id}}\otimes\Delta)$ $(R^{\psi}_{21}\cdot 1\otimes K\cdot R)=$ (A.23)
$=({\operatorname{id}}\otimes\Delta)(R^{\psi}_{21})\cdot J^{-1}_{23}\cdot(1\otimes 1\otimes K)\cdot R^{\psi}_{23}\cdot(1\otimes K\otimes 1)\cdot R_{13}\cdot R_{12}$ (A.24)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot R^{\psi}_{21}\cdot(1\otimes 1\otimes K)\cdot R^{\psi}_{23}\cdot(1\otimes K\otimes 1)\cdot R_{13}\cdot R_{12}$ (A.25)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes K)\cdot R^{\psi}_{21}\cdot R^{\psi}_{23}\cdot R_{13}\cdot(1\otimes K\otimes 1)\cdot R_{12}$ (A.26)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes K)\cdot R_{13}\cdot R^{\psi}_{23}\cdot R^{\psi}_{21}\cdot(1\otimes K\otimes 1)\cdot R_{12}$ (A.27)
where in the last equality we used (A.3). Finally, note that
$(\varepsilon\otimes{\operatorname{id}})({{\mathbb{K}}})=(\varepsilon\otimes{\operatorname{id}})(R^{\psi}_{21}\cdot 1\otimes K\cdot R)=K\,.$ (A.28)
(2) If ${{\mathbb{K}}}=R^{\psi}_{21}\cdot 1\otimes K\cdot R$ is a universal tensor K-matrix with respect to $(B,\psi,J)$ then (A.7) is just the defining condition ${{\mathbb{K}}}\in B\otimes A$. ∎
A.5. Gauge transformations
Given a quasitriangular bialgebra $(A,R)$ and a coideal subalgebra $B$. Then it is straightforward to check that the following assignments define an action of $A^{\times}$ on cylindrical structures called gauge transformation (also see [AV22a, Rmk. 8.11]):
$(\psi,J,K)\mapsto(\operatorname{Ad}(g)\circ\psi,(g\otimes g)\cdot J\cdot\Delta(g)^{-1},g\cdot K),\qquad g\in A^{\times}.$
This immediately lifts to an action on reflection structures, with the action on the tensor K-matrix given by
${{\mathbb{K}}}\mapsto(1\otimes g)\cdot{{\mathbb{K}}},$ (A.29)
which is compatible with Lemma A.3.2. Note that the support condition (A.7) is invariant under this action so that, in order to prove (A.7), it suffices to do so for a convenient choice of gauge transformation.
Appendix B Boundary bimodule categories
In this section, we introduce the notion of a boundary bimodule category tailored around the category of modules over a reflection bialgebra, see Section A.3. This framework encodes in particular the action of the tensor K-matrix of a quantum symmetric pair of arbitrary type.
As braided monoidal categories are equipped with a natural action of the braid group on the tensor powers of their objects, boundary bimodule categories give rise to representations of a cylindrical braid groupoid whose definition relies on Cherednik’s generalized reflection equation (see Section B.6).
Conventions. A boundary bimodule category is a category $\mathcal{M}$ acted upon by a braided monoidal category ${\mathcal{C}}$ and equipped with an extra structure. In the cases of our interest, the monoidal structure on ${\mathcal{C}}$ is always given by the ordinary tensor product of vector spaces. Hence the associativity and the unit constraints will be omitted for simplicity. In contrast, the monoidal action of ${\mathcal{C}}$ on $\mathcal{M}$, while being canonically unital, has a non–canonical associativity constraint, which will therefore not be omitted.
B.1. Module categories
We briefly review the notion of a module category $\mathcal{M}$ over a (braided) monoidal category ${\mathcal{C}}$ as presented in [Har01, Kol22]. Let $\mathcal{C}$ be a monoidal category with tensor product $\otimes$ and unit $\mathbf{1}$.
A (right) monoidal action of $\mathcal{C}$ on a category $\mathcal{M}$ is a functor :×$\vartriangleleft\colon\mathcal{M}\times\mathcal{C}\to\mathcal{M}$ together with an associativity constraint
$\Phi\colon\vartriangleleft\circ({\operatorname{id}}\times\otimes)\to\vartriangleleft\circ(\vartriangleleft\times{\operatorname{id}})$ (B.1)
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, $M\vartriangleleft{\bf 1}=M$ ;
  2. $\bullet$
    for any $M\in\mathcal{M}$, $V\in{\mathcal{C}}$, $\Phi_{M{\bf 1}V}={\operatorname{id}}_{M\vartriangleleft V}=\Phi_{MV{\bf 1}}$ ;
  3. $\bullet$
    for any $M\in\mathcal{M}$ and $U,V,W\in{\mathcal{C}}$,
[…]
${\bf J}_{MV}\colon F(M)\vartriangleleft V\to F(M\vartriangleleft V)$ (B.3)
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, ${\bf J}_{M{\bf 1}}={\operatorname{id}}_{F(M)}$ ;
  2. $\bullet$
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$,
[…]
Henceforth, we refer to the datum $(F,{\bf J})$ as a module functor $\mathcal{M}\to{\mathcal{N}}$.
Example B.1.1.
Set $\mathcal{C}={\operatorname{Mod}}{(U_{q}({\mathfrak{g}}))}$ and $\mathcal{M}={\operatorname{Mod}}{(U_{q}(\mathfrak{k}))}$. Since $U_{q}(\mathfrak{k})$ is a coideal subalgebra in $U_{q}({\mathfrak{g}})$, the usual tensor product induces a functor
$\vartriangleleft\colon{\operatorname{Mod}}{(U_{q}(\mathfrak{k}))}\times{\operatorname{Mod}}{(U_{q}({\mathfrak{g}}))}\to{\operatorname{Mod}}{(U_{q}(\mathfrak{k}))}$
which is readily verified to be a monoidal action of ${\operatorname{Mod}}{(U_{q}({\mathfrak{g}}))}$ on ${\operatorname{Mod}}{(U_{q}(\mathfrak{k}))}$. normal-$\triangledown$
B.2. Braided module categories
Let ${\mathcal{C}}$ be a braided monoidal category with braiding ${\bf c}\colon\otimes\to\otimes^{\scriptscriptstyle\operatorname{op}}$ and $\mathcal{M}$ a right module category over ${\mathcal{C}}$. For any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$, set
$({\operatorname{id}}_{M}\vartriangleleft{\bf c}_{VW})_{\Phi}=\Phi_{MWV}\circ({\operatorname{id}}_{M}\vartriangleleft{\bf c}_{VW})\circ\Phi_{MVW}^{-1}$
Following [Bro13, BZBJ18, Kol20], a module braiding on $\mathcal{M}$ is a natural automorphism
${\bf d}_{MV}\colon M\vartriangleleft V\to M\vartriangleleft V$
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, ${\bf d}_{M{\bf 1}}={\operatorname{id}}_{M}$ ;
  2. $\bullet$
    for any $M\in\mathcal{M}$, $V,W\in{\mathcal{C}}$,
[…]
  3. $\bullet$
    for any $M\in\mathcal{M}$, $V,W\in{\mathcal{C}}$,
[…]
A module functor $F:\mathcal{M}\to{\mathcal{N}}$ is braided if the module structure on $F$ intertwines the braidings, i.e., if the diagram
[…]
commutes for any $M\in\mathcal{M}$ and $V\in{\mathcal{C}}$.
Remarks B.2.1.
  1. (1)
    The identity ${\operatorname{id}}_{M\vartriangleleft V}$ does not define a module braiding on $\mathcal{M}$ unless the monoidal category ${\mathcal{C}}$ is symmetric, i.e., ${\bf c}_{WV}={\bf c}_{VW}^{-1}$.
  2. (2)
    It is pointed out in [BZBJ18, Rmk. 3.6] that there is an infinite family of possible axioms for a braided module category. The relations (B.5) and (B.6) constitute just one example of such axioms.
  3. (3)
    In [Kol20], Kolb proved that the action of the tensor K-matrix of a quantum symmetric pair $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ with $\dim{\mathfrak{g}}<\infty$ gives rise to a braided module structure on ${\operatorname{Mod}_{\sf fd}}(U_{q}(\mathfrak{k}))$ over ${\operatorname{Mod}_{\sf fd}}(U_{q}({\mathfrak{g}}))$. This is no longer true when $\dim{\mathfrak{g}}=\infty$, which motivates the notion of a boundary bimodule category, which we introduce in Sections B.3-B.4. $\triangledown$
B.3. Bimodule categories
Let ${\mathcal{C}}$ be a monoidal category. A left monoidal action of ${\mathcal{C}}$ on a category $\mathcal{M}$ is a functor :×$\blacktriangleleft\colon\mathcal{M}\times{\mathcal{C}}\to\mathcal{M}$ together with a natural isomorphism
$\Phi_{MVW}:M\blacktriangleleft(V\otimes W)\to(M\blacktriangleleft W)\blacktriangleleft V$
for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$.
A bimodule category over ${\mathcal{C}}$ is a category $\mathcal{M}$ equipped with
  1. $\bullet$
    a right monoidal action :×$\vartriangleleft\colon\mathcal{M}\times{\mathcal{C}}\to\mathcal{M}$ with associativity constraints $\Phi^{\vartriangleleft}$ ;
  2. $\bullet$
    a left monoidal action :×$\blacktriangleleft\colon\mathcal{M}\times{\mathcal{C}}\to\mathcal{M}$ with associativity constraints $\Phi^{\blacktriangleleft}$ ;
  3. $\bullet$
    a natural isomorphism
    ${\bf e}_{MVW}\colon(M\blacktriangleleft V)\vartriangleleft W\to(M\vartriangleleft W)\blacktriangleleft V$ (B.8)
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, $V\in{\mathcal{C}}$, ${\bf e}_{MV{\bf 1}}={\operatorname{id}}_{M\blacktriangleleft V}$ and ${\bf e}_{M{\bf 1}W}={\operatorname{id}}_{M\vartriangleleft W}$ ;
  2. $\bullet$
    for any $M\in\mathcal{M}$, $U,V,W\in{\mathcal{C}}$,
    {tikzcd} (B.9)
  3. $\bullet$
    for any $M\in\mathcal{M}$, $U,V,W\in{\mathcal{C}}$,
    {tikzcd} (B.10)
Remark B.3.1.
[…]
Let $\mathcal{M},{\mathcal{N}}$ be bimodule categories [^13] By abuse of notation, we use the same symbol to denote the monoidal actions on $\mathcal{M}$ and ${\mathcal{N}}$. and $F:\mathcal{M}\to{\mathcal{N}}$ a functor. A bimodule structure on $F$ is the datum of a pair of natural isomorphisms
${\bf J}^{\vartriangleleft}_{MV}\colon F(M)\vartriangleleft V\to F(M\vartriangleleft V)\quad\mbox{and}\quad{\bf J}^{\blacktriangleleft}_{MV}\colon F(M)\blacktriangleleft V\to F(M\blacktriangleleft V)$ (B.11)
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, $V,W\in{\mathcal{C}}$,
    {tikzcd} (B.12)
  2. $\bullet$
    for any $M\in\mathcal{M}$, $V,W\in{\mathcal{C}}$,
    {tikzcd} (B.13)
  3. $\bullet$
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$,
[…]
B.4. Boundary bimodule categories
Let ${\mathcal{C}}$ be a braided monoidal category with braiding ${\bf c}\colon\otimes\to\otimes^{\scriptscriptstyle\operatorname{op}}$ and $\mathcal{M}$ a bimodule category. A boundary structure on $\mathcal{M}$ is a natural isomorphism
${\bf f}_{MV}\colon M\vartriangleleft V\to M\blacktriangleleft V$ (B.15)
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$,
[…]
  2. $\bullet$
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$,
[…]
Remark B.4.1.
Consider ${\mathcal{C}}$ as a module category over itself. Then, every braided structure on ${\mathcal{C}}$ (see Section B.2) is also a boundary with respect the bimodule structure described in Remark B.3.1. normal-$\triangledown$
Let $\mathcal{M},{\mathcal{N}}$ be boundary bimodule categories and $F\colon\mathcal{M}\to{\mathcal{N}}$ a bimodule functor. Then, $F$ is boundary if the bimodule structures intertwines the boundary structures, i.e.,
[…]
for any $M\in\mathcal{M}$ and $V\in{\mathcal{C}}$.
B.5. From reflection bialgebras to boundary structures
The notion of a boundary structure on a bimodule category is tailored to encode the defining datum of a reflection bialgebra (see Section A.3). From
Proposition B.5.1.
Let $(A,R,\psi,J,B,{\mathbb{K}})$ be a reflection bialgebra. The following datum defines a boundary bimodule structure on ${\operatorname{Mod}}(B)$ over ${\operatorname{Mod}}(A)$.
  1. (1)
    Module structure. There is a right monoidal action $\vartriangleleft\colon{\operatorname{Mod}}(B)\times{\operatorname{Mod}}(A)\to{\operatorname{Mod}}(B)$ with trivial associativity constraint given by the standard tensor product, i.e., $M\vartriangleleft V=M\otimes V$
  2. (2)
[…]
    $\Phi^{\blacktriangleleft}_{MVW}\colon M\otimes(V\otimes W)^{\psi}\to M\otimes{W^{\psi}}\otimes{V^{\psi}}$
    is provided by the action of $J$. The commutativity constraint
    ${\bf e}_{MVW}\colon M\otimes{V^{\psi}}\otimes W\to M\otimes W\otimes{V^{\psi}}$
    is provided by the action of the R-matrix $R$ on the second and third factor.
  3. (3)
    Boundary structure. The boundary structure
    ${\bf f}_{MV}\colon M\otimes V\to M\otimes{V^{\psi}}$
    is provided by the action of the tensor K-matrix ${\mathbb{K}}$.
B.6. The model of a boundary bimodule category
In a braided monoidal category, every tensor power $V^{\otimes n}$ is naturally acted upon by the ordinary braid group (Artin-Tits group of type ${\mathsf{A}}_{n-1}$). Similarly, in a braided module category, every object of the form $W\vartriangleleft V^{\otimes n}$ is naturally acted upon by the cylindrical braid group (Artin-Tits group of type ${\mathsf{B}}_{n}$). In the case of boundary bimodule categories, the symmetry is encoded instead by a cylindrical braid groupoid ${\mathsf{M}}$, whose generators satisfy constant versions of Cherednik’s generalized reflection equation [Che92, Sec. 4]. Its definition can be thought of as a boundary analogue of the free PROP, see, e.g., [Law63, Mac65, KS20].
Let ${\mathsf{C}}$ be the strict monoidal category defined as follows.
  1. $\bullet$
    The objects of ${\mathsf{C}}$ are the non–negative integers. For any $n\in{\mathbb{Z}}_{\geqslant 0}$, the corresponding object is denoted $[n]$.
  2. $\bullet$
    The tensor product in ${\mathsf{C}}$ is the sum and the unit is zero.
  3. $\bullet$
    The morphisms in ${\mathsf{C}}$ are generated by an invertible element $\mathbf{t}\in\operatorname{End}_{{\mathsf{C}}}([2])$ satisfying the braid relation
    $\mathbf{t}_{12}\mathbf{t}_{23}\mathbf{t}_{12}=\mathbf{t}_{23}\mathbf{t}_{12}\mathbf{t}_{23}$ (B.19)
    in $\operatorname{End}_{{\mathsf{C}}}([3])$.
Remark B.6.1.
Let $V$ be an object in a braided monoidal category ${\mathcal{C}}$ with braiding $\mathbf{c}$. Then, there is an essentially unique141414 The functor ${\mathcal{G}}$ depends upon the choice of a bracketing on $n$ elements for every $n$, see, e.g., [ATL19, Sec. 7] for the precise definition. monoidal functor ${\mathcal{G}}:{\mathsf{C}}\to{\mathcal{C}}$ such that ${\mathcal{G}}([n])=V^{\otimes n}$ and ${\mathcal{G}}(\mathbf{t})=\mathbf{c}_{VV}$. normal-$\triangledown$
Let ${\mathsf{M}}$ be the strict bimodule category over ${\mathsf{C}}$ defined as follows.
  1. $\bullet$
    The objects of ${\mathsf{M}}$ are given by {}×$\{\star\}\times\coprod_{n\geqslant 0}{\mathsf{A}}^{n}$, where ${\mathsf{A}}=\{\pm\}$.
  2. $\bullet$
    The morphisms in ${\mathsf{M}}$ are generated by the following elements.
    • For any ${{\mathsf{s}}}\in{\mathsf{M}}$, three invertible elements
      $\mathbf{t}^{++}\in\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}++\,,\,{\mathsf{s}}++)$ (B.20)
      $\mathbf{t}^{--}\in\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}--\,,\,{\mathsf{s}}--)$ (B.21)
      $\mathbf{t}^{-+}\in\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}-+\,,\,{\mathsf{s}}+-)$ (B.22)
      subject to the mixed braid relations
      $\mathbf{t}^{\tau\zeta}_{12}\mathbf{t}^{\sigma\zeta}_{23}\mathbf{t}^{\sigma\tau}_{12}=\mathbf{t}^{\sigma\tau}_{23}\mathbf{t}^{\sigma\zeta}_{12}\mathbf{t}^{\tau\zeta}_{23}$ (B.23)
      in $\operatorname{Hom}_{\mathsf{C}}({\mathsf{s}}\,\sigma\,\tau\,\zeta,{\mathsf{s}}\,\zeta\,\tau\,\sigma)$, for all $\sigma,\tau,\zeta\in{\mathsf{A}}$ such that the corresponding operators are all defined.
    • For any ${{\mathsf{s}}}\in{\mathsf{M}}$, an invertible element
      $\mathbf{b}\in\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}\,+\,,\,{\mathsf{s}}\,-)$
      subject to Cherednik’s reflection equation
      $\mathbf{t}^{--}\mathbf{b}\mathbf{t}^{-+}\mathbf{b}=\mathbf{b}\mathbf{t}^{-+}\mathbf{b}\mathbf{t}^{++}$ (B.24)
      in $\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}\,++\,,\,{\mathsf{s}}\,--)$.
  3. $\bullet$
    The right and left actions of $[n]\in{\mathsf{C}}$ are given by concatenation with length-$n$ sequences of $+$ and --, respectively. The associativity constraints are trivial. The commutativity constraint (B.8) is given by $\mathbf{t}^{-+}$. Finally, the boundary structure (B.15) is given by $\mathbf{b}$.
Proposition B.6.2.
Let ${\mathcal{C}}$ be a braided monoidal category and $\mathcal{M}$ a boundary bimodule category.
  1. (1)
    For any $V\in{\mathcal{C}}$, the functor ${\mathcal{G}}:{\mathsf{C}}\to{\mathcal{C}}$ defined in Remark B.6.1 induces on $\mathcal{M}$ a boundary bimodule structure over ${\mathsf{C}}$.
  2. (2)
    For any $M\in\mathcal{M}$, there is an essentially unique boundary bimodule functor ${\mathcal{F}}:{\mathsf{M}}\to\mathcal{M}$ such that ${\mathcal{F}}(\star)=M$ and ${\mathcal{F}}(\mathbf{b})={\bf f}$.