### v2

# Tensor K-matrices for quantum symmetric pairs

Abstract.
Let ${\mathfrak{g}}$ be a symmetrizable Kac-Moody algebra, $U_{q}({\mathfrak{g}})$ its quantum group, and $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ a quantum symmetric pair subalgebra determined by a Lie algebra automorphism $\theta$. We introduce a category $\mathcal{W}_{\theta}$ of weight $U_{q}(\mathfrak{k})$-modules, which is acted on by the category of weight $U_{q}({\mathfrak{g}})$-modules via tensor products. We construct a universal tensor K-matrix ${{\mathbb{K}}}$ (that is, a solution of a reflection equation) in a completion of $U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{g}})$. This yields a natural operator on any tensor product $M\otimes V$, where $M\in\mathcal{W}_{\theta}$ and $V\in{\mathcal{O}}_{\theta}$, i.e., $V$ is a $U_{q}({\mathfrak{g}})$-module in category ${\mathcal{O}}$ satisfying an integrability property determined by $\theta$. Canonically, $\mathcal{W}_{\theta}$ is equipped with a structure of a bimodule category over ${\mathcal{O}}_{\theta}$ and the action of ${{\mathbb{K}}}$ is encoded by a new categorical structure, which we call a boundary structure on $\mathcal{W}_{\theta}$. This generalizes a result of Kolb which describes a braided module structure on finite-dimensional $U_{q}(\mathfrak{k})$-modules when ${\mathfrak{g}}$ is finite-dimensional.
We apply our construction to the case where ${\mathfrak{g}}$ is an affine Lie algebra, providing the most comprehensive universal framework to date for trigonometric K-matrices. In this case we obtain a formal tensor K-matrix valued in the endomorphisms of the tensor product of any module in $\mathcal{W}_{\theta}$ and any finite-dimensional module over the corresponding quantum affine algebra. This formal series can be normalized to a trigonometric K-matrix if the factors in the tensor product are both finite-dimensional irreducible modules over the quantum affine algebra.
Key words and phrases:
Reflection equation; quantum Kac-Moody algebras; quantum symmetric pairs
2020 Mathematics Subject Classification:
Primary: 81R50. Secondary: 16T25, 17B37, 81R12.
The first author is partially supported by the Programme FIL 2022 of the University of Parma and co-sponsored by Fondazione Cariparma, and by an INdAM Project Grant 2024. Both authors are partially supported by the International Scientists Program of the Beijing Natural Sciences Foundation (grant number IS24003).
1. Introduction
1.1.
The reflection equation was introduced by Cherednik [Che84] and Sklyanin [Skl88] in the context of quantum integrable systems with compatible boundary conditions. The underlying algebraic theory was outlined by Sklyanin and Kulish in [KS92]. In [AV22a] we proved that the Drinfeld–Jimbo quantum group $U_{q}({\mathfrak{g}})$, where ${\mathfrak{g}}$ is a symmetrizable Kac–Moody algebra, is a natural source of solutions, which are referred to as K-matrices, of a generalized version of the reflection equation, introduced by Cherednik [Che92].
Our result is a generalization of previous constructions by Bao and Wang [BW18b] and Balagović and Kolb [BK19]. In particular, it relies on a more refined initial datum, which is essential for extending the theory to arbitrary quantum Kac-Moody algebras, and subsequently to finite-dimensional modules over quantum affine algebras in [AV22b].
In [Skl88], Sklyanin observed that a K-matrix on a vector space $V$ could be promoted to a solution of the reflection equation acting on a larger tensor product. This is achieved by left- and right-multiplying the initial K-matrix with suitable R-matrices, i.e., solutions of the Yang-Baxter equation. In this paper, we apply this principle to the universal K-matrices obtained in [AV22a], leading to a natural construction of tensor K-matrices. This is a generalization of the result of Kolb [Kol20] for quantum groups of finite type. In the case of quantum affine algebras, this yields (a family of) trigonometric tensor K-matrices on any tensor product of irreducible finite-dimensional representations.
1.2.
The universal K-matrices constructed in [AV22a] for the quantum Kac-Moody algebra $U_{q}({\mathfrak{g}})$ depend on two data. The first one is a quantum symmetric pair subalgebra, i.e., a right coideal subalgebra $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$, which quantizes a (pseudo-)fixed-point Lie subalgebra with respect to a (pseudo-)involution $\theta$ of the second kind [Let02, Kol14, RV22, AV22a]. We refer to $U_{q}(\mathfrak{k})$ as a QSP subalgebra. The second one is a distinguished algebra automorphism $\psi$ of $U_{q}({\mathfrak{g}})$, which is essentially determined by $U_{q}(\mathfrak{k})$, see [AV22a]. We refer to it as a twist automorphism.
For any such pair $(U_{q}(\mathfrak{k}),\psi)$, we construct an operator $K$ on $U_{q}({\mathfrak{g}})$-modules in category ${\mathcal{O}}$, which satisfies the twisted reflection equation [^1] We use the shorthand notations $R^{\psi}=(\psi\otimes{\operatorname{id}})(R)$, and $R^{\psi\psi}=(\psi\otimes\psi)(R)$.
$R^{\psi\psi}_{21}\cdot(1\otimes K)\cdot R^{\psi}\cdot(K\otimes 1)=(K\otimes 1)\cdot R^{\psi}_{21}\cdot(1\otimes K)\cdot R,$ (1.1)
where $R$ is the universal R-matrix of $U_{q}({\mathfrak{g}})$. Moreover, $K$ satisfies the QSP intertwining equation
$K\cdot b=\psi(b)\cdot K$ (1.2)
where $b\in U_{q}(\mathfrak{k})$. In fact, $K$ is obtained as the unique solution of (1.2) among operators with a prescribed form. This is consistent with the examples of explicit K-matrices provided in [MN98, DG02, DM03].
More precisely, every universal K-matrix obtained in [AV22a] is eventually built out of the quasi-K-matrix $\Upsilon$, an operator originally introduced in [BW18b] as a canonical solution of a certain QSP intertwining equation for a QSP subalgebra of $U_{q}(\mathfrak{sl}_{N})$ and generalized in [BK19, AV22a] to almost arbitrary [^2] There is a natural constraint on $\theta$ extending to a map on the extended weight lattice, which is automatically satisfied for ${\mathfrak{g}}$ of finite and affine type, see [Kol14, AV22a]. QSP subalgebras of $U_{q}({\mathfrak{g}})$. Although it may appear different at first, this equation reduces to the form (1.2) after a simple algebraic manipulation. Relying on this observation, we prove in [AV22a] that $\Upsilon$ satisfies a generalized reflection equation of the form (1.1). In contrast with Balagović and Kolb’s universal K-matrix $K_{\sf BK}$ [BK19] defined in finite type, $\Upsilon$ acts on any $U_{q}({\mathfrak{g}})$-module in category ${\mathcal{O}}$.
New K-matrices are then obtained by gauging simultaneously $\Upsilon$ and its twist automorphism. Such gauge, however, may lead to convergence issues, which requires to consider a smaller category of $U_{q}({\mathfrak{g}})$-modules. For instance, when ${\mathfrak{g}}$ is of finite type, we obtain a family of universal K-matrices whose action interpolates between that of $\Upsilon$ on category ${\mathcal{O}}$ modules, and that of $K_{\sf BK}$ on integrable category ${\mathcal{O}}$ modules (i.e., finite-dimensional modules).
In the rest of the introduction we will focus on the quasi K-matrix $\Upsilon$, and use $\psi$ to denote the twist automorphism corresponding to $\Upsilon$.
1.3.
In the present paper, we describe a 2-tensor version of the universal K-matrix, which we denote by ${{\mathbb{K}}}$ and refer to as a tensor K-matrix. This generalizes the results from [Kol20] in the same way that the construction of the universal K-matrix given in [AV22a] generalizes the results from [BK19].
Inspired by Sklyanin’s construction in [Skl88], we define the standard tensor K-matrix as the operator
${{\mathbb{K}}}=R^{\psi}_{21}\cdot(1\otimes\Upsilon)\cdot R,$ (1.3)
which a priori acts only on a tensor product of $U_{q}({\mathfrak{g}})$-modules in category ${\mathcal{O}}$.
An analogous formula was considered by Kolb in [Kol20] for quantum groups of finite type, with $\Upsilon$ replaced by $K_{\sf BK}$ and $\psi$ by a diagram automorphism. By direct inspection, he proved that this datum is supported on $U_{q}(\mathfrak{k})$, i.e.,
${{\mathbb{K}}}\in U_{q}(\mathfrak{k})\widehat{\otimes}U_{q}({\mathfrak{g}})$ (1.4)
for some completion $U_{q}(\mathfrak{k})\widehat{\otimes}U_{q}({\mathfrak{g}})$ of $U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{g}})$. The upshot is that ${{\mathbb{K}}}$ acts on any tensor product of the form $M\otimes V$ with $M$ and $V$ finite-dimensional modules over $U_{q}(\mathfrak{k})$ and $U_{q}({\mathfrak{g}})$, respectively. [^3] In contrast with our construction, the tensor K-matrix constructed by Kolb does not act on a tensor product of non-integrable category ${\mathcal{O}}$ modules, e.g., Verma modules. Combined with the QSP intertwining and coproduct identities naturally satisfied by ${{\mathbb{K}}}$, the support property amounts to a concise categorical statement. Let $\mathcal{M}$ and ${\mathcal{C}}$ be the categories of finite-dimensional modules over $U_{q}(\mathfrak{k})$ and $U_{q}({\mathfrak{g}})$, respectively. Owing to the coideal property, $\mathcal{M}$ is a (right) module category over ${\mathcal{C}}$. Then, the action of ${{\mathbb{K}}}$ is easily seen to define a braided module structure444 More precisely, in order to have a genuine braided module structure, for certain QSP subalgebras, ${\mathcal{C}}$ needs to be a ${\mathbb{Z}}/2{\mathbb{Z}}$-equivariantization of the category of finite-dimensional $U_{q}({\mathfrak{g}})$-modules [Kol20, Sec. 3.2]. on $\mathcal{M}$ over ${\mathcal{C}}$.
For arbitrary ${\mathfrak{g}}$, however, this fails, since twisting by $\psi$ does not preserve ${\mathcal{O}}$.
1.4.
We develop instead an alternative approach to the support condition (1.4), which is based on the following. It is easily checked that $\operatorname{Ad}(\Upsilon)$ is not [^5] In the same way, $\operatorname{Ad}(R)$ is not an automorphism of $U_{q}({\mathfrak{g}})^{\otimes 2}$, although it preserves the subalgebra $\Delta(U_{q}({\mathfrak{g}}))$, acting as the simple transposition of tensor factors. an automorphism of $U_{q}({\mathfrak{g}})$, but it does act on $U_{q}(\mathfrak{k})$ as the restriction of $\psi\in\operatorname{Aut}_{\sf alg}(U_{q}({\mathfrak{g}}))$.
Instead, let $\mathscr{E}$ be the completion of $U_{q}({\mathfrak{g}})$ with respect to category ${\mathcal{O}}$ $\mathscr{E}$ is the algebra of endomorphisms of the forgetful functor from category ${\mathcal{O}}$ modules to vector spaces, see Section 2.4. By [ATL24], $U_{q}({\mathfrak{g}})$ embeds in $\mathscr{E}$. Since $\Upsilon\in\mathscr{E}$, $\operatorname{Ad}(\Upsilon)$ can be thought of as an endomorphism of $\mathscr{E}$, which does not preserve $U_{q}({\mathfrak{g}})$. Then we consider the subalgebra
$B_{\xi}=\{x\in U_{q}({\mathfrak{g}})\,|\,(\xi\otimes{\operatorname{id}})(\Delta(x))=\Delta(x)\},$ (1.5)
where
$\xi:=\operatorname{Ad}(\Upsilon)^{-1}\circ\psi$ (1.6)
and the identity $(\xi\otimes{\operatorname{id}})(\Delta(x))=\Delta(x)$ is to be understood in completion of $U_{q}({\mathfrak{g}})^{\otimes 2}$ with respect to category ${\mathcal{O}}$. Note that an automorphism similar to $\xi$ was considered in [KY20] and shown to exist for Nichols algebras of diagonal type.
By definition, $B_{\xi}$ is the maximal coideal subalgebra contained in the fixed-point subalgebra $U_{q}({\mathfrak{g}})^{\xi}$. Moreover, $U_{q}(\mathfrak{k})\subseteq B_{\xi}$ by (1.2). It readily follows from (1.1) that ${{\mathbb{K}}}$ belongs to a completion of $B_{\xi}\otimes U_{q}({\mathfrak{g}})$. Finally, we prove that $U_{q}(\mathfrak{k})$ has the same maximality property of $B_{\xi}$, i.e.,
$U_{q}(\mathfrak{k})=B_{\xi}.$ (1.7)
The proof relies on three ingredients: the study of the classical limit of $\Upsilon$, carried out in Section 4, the maximality property of $U_{q}(\mathfrak{k})$ proved by Kolb in [Kol14], and the Iwasawa decomposition of ${\mathfrak{g}}$ with respect to any pseudo-fixed-point subalgebra ${\mathfrak{k}}$, established by the second author and Regelskis in [RV22].
From (1.7), we obtain (1.4) as follows. We consider a full subcategory $\mathcal{W}_{\theta}\subset{\operatorname{Mod}}(U_{q}(\mathfrak{k}))$ whose objects are called QSP weight modules, see [BW18a, Wa24] for similar constructions.777 Up to twisting by an automorphism of $U_{q}(\mathfrak{k})$, any finite-dimensional $U_{q}(\mathfrak{k})$-module lies in $\mathcal{W}_{\theta}$. However, even in finite type, this category is significantly larger than the category of finite-dimensional $U_{q}(\mathfrak{k})$-modules considered in [Kol20]. Then, we prove that ${{\mathbb{K}}}$ naturally acts on any tensor product $M\otimes V$, where $M\in\mathcal{W}_{\theta}$ and $V$ belongs to a full subcategory888 We denote this category by ${\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}$. Its objects are modules in category ${\mathcal{O}}$ whose restriction to $U_{q}({\mathfrak{g}}_{X})$ is integrable, where $(X,\tau)$ is the generalized Satake diagram associated to $\theta$. of ${\mathcal{O}}$ satisfying an integrability property dictated by $U_{q}(\mathfrak{k})$.
1.5.
Let ${\mathfrak{g}}$ be an affine Kac-Moody algebra. In [AV22b], we proved that the action of the quasi-K-matrix $\Upsilon$ descends to finite–dimensional $U_{q}({\mathfrak{g}}^{\prime})$-modules, and induces trigonometric K-matrices on any irreducible module. In Section 6 we extend this results to tensor K-matrices. Specifically, we observe that the action of ${{\mathbb{K}}}$ descends to tensor products of the form $M\otimes V$ with $M\in\mathcal{W}_{\theta}$ and $V$ a finite-dimensional $U_{q}({\mathfrak{g}})$-module. Then we consider the full subcategory ${\mathcal{C}}_{\theta}\subset\mathcal{W}_{\theta}$ whose objects are finite-dimensional. By (1.3), we obtain trigonometric tensor K-matrices whenever $V$ is irreducible and $M\in{\mathcal{C}}_{\theta}$ is the restriction of an irreducible $U_{q}({\mathfrak{g}})$-module.
Following [AV22b], we actually obtain trigonometric tensor K-matrices whenever $M\otimes V\in{\mathcal{C}}_{\theta}$ is generically irreducible. However, it is unclear whether this would produce new examples of trigonometric tensor K-matrices other than those obtained by restrictions. For instance, when $U_{q}({\mathfrak{g}}^{\prime})=U_{q}(\widehat{\mathfrak{sl}}_{2})$ and $U_{q}(\mathfrak{k})$ is the q-Onsager algebra, i.e., the QSP subalgebra corresponding to the Chevalley involution, Ito and Terwilliger proved that every irreducible object in ${\mathcal{C}}_{\theta}$ is obtained by restriction of an irreducible $U_{q}({\mathfrak{g}}^{\prime})$-module [IT09]. Thus, a posteriori, in this case we obtain a trigonometric K-matrix on any tensor product $M\otimes V$ with $M$ irreducible in ${\mathcal{C}}_{\theta}$. For other affine quantum symmetric pairs, however, the problem is completely open.
1.6.
The tensor K-matrix allows us to address several open problems in the representation theory of affine quantum symmetric pairs with applications to quantum integrable systems with two boundaries. Below, we provide a brief overview, which will be the focus of forthcoming works.
1.6.1.
An important application of the universal formalism of transfer matrices built out of R-matrices is the study of q-characters for finite-dimensional modules of quantum affine algebras $U_{q}({\mathfrak{g}}^{\prime})$ [FR99, FM01]. In the original setting one considers the grading-shifted universal R-matrix of a quantum affine algebra $U_{q}({\mathfrak{g}}^{\prime})$, and evaluates only one factor on a finite-dimensional representation $V$ of $U_{q}({\mathfrak{g}}^{\prime})$. By taking the trace, one obtains a formal series ${\mathsf{t}}_{V}(z)$ valued in the quantum Borel subalgebra $U_{q}({\mathfrak{b}}^{+})\subset U_{q}({\mathfrak{g}})$. From the Yang-Baxter equation, it readily follows that $[{\mathsf{t}}_{V}(z),{\mathsf{t}}_{W}(z^{\prime})]=0$.
For the tensor K-matrix, we obtain an analogous result by applying Sklyanin’s formalism of two-boundary transfer matrices [Skl88]. This involves a correction of the universal tensor K-matrix of an affine quantum symmetric pair $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}}^{\prime})$ by fixing a dual QSP subalgebra, associated to the opposite boundary and hence by nature a left coideal subalgebra. Through evaluation of the second factor on $V$ and taking the trace, we obtain a family of commuting series ${\mathsf{b}}_{V}(z)$ valued in $U_{q}(\mathfrak{k})$, and a morphisms of rings
${\left[{\operatorname{Mod}_{\sf fd}}(U_{q}({\mathfrak{g}}^{\prime}))\right]}$${U_{q}(\mathfrak{k})}$$\scriptstyle{{\mathsf{b}}}$ (1.8)
where $\left[{\operatorname{Mod}_{\sf fd}}(U_{q}({\mathfrak{g}}^{\prime}))\right]$ is the Grothendieck ring. It is natural to expect that ${\mathsf{b}}$ extends to a boundary analogue of the q-character map, providing a refined tool for the finite-dimensional representation theory of $U_{q}({\mathfrak{g}}^{\prime})$. This requires the composition with an Harish-Chandra map associated to a Drinfeld (loop) presentation of $U_{q}(\mathfrak{k})$, which was recently found by Lu, Wang and Zhang for all restrictable999 This means that the affine node is fixed by the diagram involution $\tau$ associated to the QSP subalgebra. quasi-split types in [LW21, Zh22, LWZ23a, LWZ23b].
For split types, Li and Przeździecki consider instead an opposite q-character map
${\left[{\operatorname{Mod}_{\sf fd}}(U_{q}(\mathfrak{k}))\right]}$${U_{q}(L\mathring{{\mathfrak{h}}})}$ (1.9)
where $L\mathring{{\mathfrak{h}}}\subset{\mathfrak{g}}^{\prime}$ is the loop Cartan subalgebra [Prz24, LP24]. This is a module map, directly defined in terms of the spectra of the abelian loop generators of $U_{q}(\mathfrak{k})$. We expect that even this version can be obtained as above, by evaluating the first factor of the corrected universal tensor K-matrix on finite-dimensional $U_{q}(\mathfrak{k})$-modules.
1.6.2.
A representation-theoretic approach to Baxter Q-operators for quantum integrable systems and their functional relations was first set out in the 1990s [BLZ, AF97], by evaluating one leg of the universal R-matrix on a certain infinite-dimensional module of $U_{q}({\mathfrak{b}}^{+})$ called asymptotic representation, and taking a trace. In [HJ12, FH15] the corresponding representation theory of $U_{q}({\mathfrak{b}}^{+})$ was developed, an approach to Baxter’s functional relations was established (for all untwisted quantum affine algebras) and a conjecture from [FR99] about expressions of the eigenvalues of transfer matrices in terms of q-characters was proved.
In quasi-split type, the tensor K-matrices ${{\mathbb{K}}}$ considered here lie in a completion of $U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{b}}^{+})$. Hence it is natural to attempt such an approach to functional relations for QSP subalgebras. Note that trigonometric Baxter Q-operators defined in the style of Sklyanin’s two-boundary transfer matrices have been considered, initially in [YNZ06]. The algebraic theory for the special case of the affine quantum symmetric pair consisting of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ and the augmented q-Onsager algebra is particularly well-studied, see [BT18, VW20, Ts21]. In [CVW24] this was connected to the universal K-matrix $K$ residing in a completion of $U_{q}({\mathfrak{b}}^{+})$ constructed in [AV22a, AV22b].
1.6.3.
Transfer matrices built out of R-matrices are closely related to Yang’s scattering matrices and hence to the operators appearing in quantum Knizhnik-Zamolodchikov (KZ) equations. Two-boundary analogues of scattering matrices and quantum KZ operators built from solutions of Yang-Baxter and reflection equations were first considered in [Che91, Che92]. In the untwisted case, the general connection between two-boundary transfer and scattering matrices was given in [Vl15]. Solutions of various types for quantum KZ equations with K-matrices coming from QSP subalgebras of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ were given in [JKKMW, RSV, SV15, HL21].
As mentioned above, the quantum affine algebra $U_{q}({\mathfrak{g}}^{\prime})$ has two important categories of representations: ${\mathcal{O}}^{\sf int}$ and finite-dimensional modules. In [FR92] quantum KZ equations built out of R-matrices were connected to a certain intertwiner from an irreducible object in ${\mathcal{O}}^{\sf int}$ to shifted tensor products of another irreducible in ${\mathcal{O}}^{\sf int}$ and a finite-dimensional module. We expect that the universal tensor K-matrix of a QSP algebra $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}}^{\prime})$ will allow for 2-boundary quantum KZ equations to be related to analogous intertwiners for certain infinite-dimensional $U_{q}(\mathfrak{k})$-modules and finite-dimensional $U_{q}({\mathfrak{g}}^{\prime})$-modules, with the construction from [JKKMW] appearing as a special case. It will also be interesting to connect such constructions to recent works on differential operators (asymptotic Knizhnik-Zamolodchikov-Bernard operators) defined in terms of classical dynamical r- and k-matrices for symmetric spaces [RS].
1.6.4.
The last application is a universal K-matrix formalism for QSP subalgebras. In [FRT90], Faddeev, Reshetikhin, and Takhtajan provided an alternative approach to Drinfeld-Jimbo quantum groups, which is commonly referred to as the R-matrix (or FRT) formalism. This yields a new presentation given in terms of a fixed matrix solution of the Yang-Baxter equation. The R-matrix formalism was extended to quantum affine algebras of classical Lie type in [RSTS90].
More recently, Gautam, Rupert, and Wendlandt provided in [GRW22, RW23] a unified approach to the R-matrix formalism for $U_{q}({\mathfrak{g}})$ when ${\mathfrak{g}}$ is finite-dimensional,101010 Their results naturally extend to the case of a symmetrizable Kac-Moody algebra. similar in spirit to the main result of [RSTS90]. They study the R-matrix algebra $U_{R}$ arising from the evaluation of the universal R-matrix of $U_{q}({\mathfrak{g}})$ on an arbitrary finite-dimensional irreducible representation. Then, they prove that (an extension of) $U_{q}({\mathfrak{g}})$ can be realized as a Hopf quotient of $U_{R}$.
Relying on the construction of universal tensor K-matrices, we can extend this approach to quantum symmetric pairs. Namely, through the evaluation of the universal tensor K-matrix associated to a quantum symmetric pair subalgebra $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$, we identify $U_{q}(\mathfrak{k})$ with a coideal subalgebra $U_{K}\subset U_{R}$. In the affine case, it is expected to recover and extend the results by Molev, Ragoucy, and Sorba [MRS] and Chen, Guay, and Ma [CGM].
1.7.
Finally, we conclude with some remarks about the algebraic and the categorical frameworks underlying our construction, which we develop in Appendices B and C, respectively.
The algebraic framework revolves around the notion of a reflection bialgebra, and it is a natural extension of the notion of a cylindrical bialgebra introduced in [AV22a]. As the latter could be regarded as a generalization of the original definition given by tom Dieck and Häring-Oldenburg [tD98, tDHO98], and later by Balagović and Kolb [BK19], the former is a straightforward adaptation of the original definitions given by Enriquez [Enr07], and later by Kolb [Kol20]. It also appears as a special case of a weakly quasitriangular comodule algebra as defined in [KY20].
Recently, Baseilhac, Gainutdinov and Lemarthe independently proposed this axiomatic scheme for comodule algebras, and applied it to the case of the alternating central extension ${\mathcal{A}}_{q}$ of the q-Onsager algebra [LBG23, Le23]. Specifically, they study fusion relations of a certain explicit element of ${{\mathcal{A}}_{q}\otimes\operatorname{End}(V)}[\negthinspace[z]\negthinspace]$, where $V$ is an evaluation representation of $U_{q}(\widehat{\mathfrak{sl}}_{2})$, and conjecture that this element is given by the action of a universal tensor K-matrix for ${\mathcal{A}}_{q}$. It is natural to expect a connection111111 The comparison is subtle because in [LBG23, Le23] a twist automorphism is used which is a priori inaccessible for the q-Onsager algebra in our framework: $\psi=\omega\circ\tau$ with $\tau$ the nontrivial diagram automorphism. This obstacle can be resolved by noting that in ibid. only modules $V$ are considered where $\tau$ acts as conjugation by an antidiagonal matrix in any basis of $V$ consisting of $U_{q}(\mathfrak{sl}_{2})$-weight vectors. with our results specialized to the q-Onsager algebra.
Our main results are also concisely summarized in terms of the categorical framework developed in Appendix C. Specifically, the action of ${{\mathbb{K}}}$ yields a boundary structure on $\mathcal{W}_{\theta}$ with respect to its bimodule structure over ${\mathcal{O}}_{\theta}$ determined by the coideal property and $\psi$. Even in finite type, this yields a generalization of the braided module structure described by Kolb [Kol20].
1.8. Outline
In Section 2 we recall the definition and fundamental properties of a quantized universal enveloping algebra $U_{q}({\mathfrak{g}})$ of a Kac-Moody algebra ${\mathfrak{g}}$ and its universal R-matrix. Next, in Section 3 we recall the definition of the subalgebra $U_{q}(\mathfrak{k})$ and the corresponding universal K-matrix in the general setting determined by pseudo-involutions of ${\mathfrak{g}}$. The quantum group analogue of the key identity (1.7) is established in Section 4, see Theorem 4.2, via a maximality property of $U_{q}(\mathfrak{k})$ and by taking a classical limit. To promote this to the existence of the tensor K-matrix ${{\mathbb{K}}}$, in Section 5 we describe the appropriate categories of $U_{q}(\mathfrak{k})$-modules and $U_{q}({\mathfrak{g}})$-modules and operators acting on them; the main result is Theorem 5.17. In Section 6 we give the tensor K-matrix version of our previous work [AV22b] on K-matrices for finite-dimensional modules of quantum loop algebras.
We describe the general tensor K-matrix formalism for quasitriangular bialgebras in Appendix B. In Appendix C the corresponding categorical-topological framework is developed, resulting in the notion of a boundary bimodule category, which gives rise to actions of cylindrical braid groupoids on tensor products.
1.9. Acknowledgments
The authors would like to thank Pascal Baseilhac, Azat Gainutdinov, and Stefan Kolb for useful discussions and their interest in this work.
2. Quantum Kac-Moody algebras
Throughout this paper we will make use of the following notation, for any $k$-linear space $V$:
$X^{\xi}=(\xi\otimes{\operatorname{id}})(X)\,,\qquad X^{\xi}_{21}=(X^{\xi})_{21}\,,\qquad\mbox{and}\qquad X^{\xi\xi}=(\xi\otimes\xi)(X)$
for any $X\in V\otimes V$ and $k$-linear map $\xi:V\to V$. Here the field $k$ should be clear from the context.
Also, given a lattice $\mathsf{\Lambda}$ whose ${\mathbb{Z}}$-basis is clear from the context, we shall denote its non-negative part by $\mathsf{\Lambda}^{+}$.
2.1. Kac-Moody algebras
Let ${\mathfrak{g}}$ be the Kac-Moody algebra defined over121212 Or over any algebraically closed field of characteristic zero. ${\mathbb{C}}$ in terms of a symmetrizable generalized Cartan matrix $A=(a_{ij})_{i,j\in I}$ [Kac90]. Let ${\mathfrak{h}}\subset{\mathfrak{g}}$ be the Cartan subalgebra. Denoting the simple roots by $\alpha_{i}$ and the simple coroots by $h_{i}$ ($i\in I$), we consider the root and coroot lattices:
$\mathsf{Q}=\operatorname{Sp}_{\mathbb{Z}}\{\alpha_{i}\,|\,i\in I\}\subset{\mathfrak{h}}^{*},\qquad\mathsf{Q}^{\vee}=\operatorname{Sp}_{\mathbb{Z}}\{h_{i}\,|\,i\in I\}\subset{\mathfrak{h}}.$ (2.1)
Let ${\bm{\Phi}}\subset\mathbb{Q}$ be the root system and ${\bm{\Phi}}^{+}\subset\mathsf{Q}^{+}$ the positive subsystem. Finally, set ${\mathfrak{g}}^{\prime}=[{\mathfrak{g}},{\mathfrak{g}}]$ and ${\mathfrak{h}}^{\prime}={\mathfrak{h}}\cap{\mathfrak{g}}^{\prime}=\mathsf{Q}^{\vee}\otimes_{\mathbb{Z}}{\mathbb{C}}$.
[…]
$\mathsf{Q}^{\vee}_{\sf ext}=\mathsf{Q}^{\vee}\oplus\bigoplus_{k=1}^{\operatorname{cork}(A)}{\mathbb{Z}}d_{k}\subset{{\mathfrak{h}}}\qquad\mbox{and}\qquad\mathsf{P}=\{\lambda\in{{\mathfrak{h}}}^{*}\;|\;\lambda(\mathsf{Q}^{\vee}_{\sf ext})\subset{\mathbb{Z}}\}.$
[…]
2.2. Drinfeld-Jimbo quantum groups
Denote by ${\mathbb{F}}$ the algebraic closure of ${\mathbb{C}}(q)$ where $q$ is an indeterminate.131313 The results of this paper are also valid for ${\mathbb{F}}={\mathbb{C}}$ and $q\in{\mathbb{C}}^{\times}$not a root of unity. Fix non-negative integers $\{\epsilon_{i}\;|\;i\in I\}$ such that the matrix $(\epsilon_{i}a_{ij})_{i,j\in I}$ is symmetric and set $q_{i}=q^{\epsilon_{i}}$. The quantum Kac-Moody algebra [Dri86, Jim85] associated to ${{\mathfrak{g}}}$ is the unital associative algebra $U_{q}({\mathfrak{g}})$ defined over ${\mathbb{F}}$ with generators $E_{i}$ and $F_{i}$ ($i\in I$), and $K_{h}$ ($h\in\mathsf{Q}^{\vee}_{\sf ext}$) subject to the following defining relations:
$K_{h}K_{h^{\prime}}=K_{h+h^{\prime}},\qquad K_{0}=1,$ (2.2)
$K_{h}E_{i}=q^{\alpha_{i}(h)}E_{i}K_{h},\qquad K_{h}F_{i}=q^{-\alpha_{i}(h)}F_{i}K_{h},$ (2.3)
$[E_{i},F_{j}]=\delta_{ij}\frac{K_{i}-K_{i}^{-1}}{q_{i}-q_{i}^{-1}},$ (2.4)
for any $i,j\in I$ and $h,h^{\prime}\in\mathsf{Q}^{\vee}_{\sf ext}$, where $K_{i}^{\pm 1}=K_{\pm\epsilon_{i}h_{i}}$, together with the quantum Serre relations (see e.g., [Lus94, Cor. 33.1.5])
$\operatorname{Serre}_{ij}(E_{i},E_{j})=\operatorname{Serre}_{ij}(F_{i},F_{j})=0\qquad\text{for all }i\neq j$ (2.5)
where $\operatorname{Serre}_{ij}$ denotes the following polynomial in two noncommuting variables:
$\operatorname{Serre}_{ij}(x,y)=\sum_{r=0}^{1-a_{ij}}(-1)^{n}\binom{1-a_{ij}}{r}_{q_{i}}x^{1-a_{ij}-r}yx^{r}$ (2.6)
with the q-deformed binomial coefficient defined in e.g., [Lus94, 1.3.3].
We consider the Hopf algebra structure on $U_{q}({\mathfrak{g}})$ uniquely determined by the following coproduct formulae
$\Delta(E_{i})$ $=E_{i}\otimes 1+K_{i}\otimes E_{i},$ $\Delta(F_{i})$ $=F_{i}\otimes K_{i}^{-1}+1\otimes F_{i},$ $\Delta(K_{h})$ $=K_{h}\otimes K_{h},$ (2.7)
for any $i\in I$ and $h\in\mathsf{Q}^{\vee}_{\sf ext}$. Let $\omega:U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})^{\operatorname{cop}}$ be the Chevalley involution
$\omega(K_{h})=K_{-h},\quad\omega(E_{i})=-F_{i},\quad\omega(F_{i})=-E_{i}.$ (2.8)
for any $i\in I$ and $h\in\mathsf{Q}^{\vee}_{\sf ext}$.
We denote by $U_{q}({\mathfrak{h}}^{\prime})$, $U_{q}{({{\mathfrak{n}}}^{+})}$, and $U_{q}{({{\mathfrak{n}}}^{-})}$ the subalgebras generated by the elements $\{K_{i}^{\pm 1}\}_{i\in I}$, $\{E_{i}\}_{i\in I}$ and $\{F_{i}\}_{i\in I}$, respectively. We set $U_{q}({\mathfrak{g}}^{\prime})=U_{q}{({{\mathfrak{n}}}^{+})}\cdot U_{q}({\mathfrak{h}}^{\prime})\cdot U_{q}{({{\mathfrak{n}}}^{-})}$ and $U_{q}{({{\mathfrak{b}}}^{\pm})}=U_{q}{({{\mathfrak{n}}}^{\pm})}\cdot U_{q}({\mathfrak{h}})$, where $U_{q}({\mathfrak{h}})$ is the commutative subalgebra generated by $K_{h}$, $h\in\mathsf{Q}^{\vee}_{\sf ext}$. The adjoint action of the $K_{h}$ induces a grading by the monoids ±$\pm\mathsf{Q}^{+}$ on the subalgebras $U_{q}{({{\mathfrak{n}}}^{\pm})}$, respectively, and we denote by $U_{q}{({{\mathfrak{n}}}^{\pm})}_{\pm\beta}$ the component of degree ±$\pm\beta$, for $\beta\in\mathsf{Q}^{+}$.
2.3. Categories of $U_{q}({\mathfrak{g}})$-modules
If $V$ is an $U_{q}({\mathfrak{h}})$–module and $\mu\in\mathsf{P}$, we denote the corresponding weight space of $V$ by
$V_{\mu}=\{v\in V\,|\,\forall h\in\mathsf{Q}^{\vee}_{\sf ext},\,\,K_{h}\,v=q^{\mu(h)}v\}.$
A $U_{q}({\mathfrak{g}})$–module $V$ is
  1. (C1)
    a (type $\bf 1$) weight module if $V=\bigoplus_{\mu\in\mathsf{P}}V_{\mu}$;
  2. (C2)
    integrable if it is a weight module, and the elements $\{E_{i},F_{i}\}_{i\in I}$ act locally nilpotently;141414 This implies that $V$ is completely reducible as a (possibly infinite) direct sum of simple finite–dimensional modules over the subalgebra generated by $E_{i},K_{i}^{\pm 1},F_{i}$, for each $i\in I$.
  3. (C3)
    in category ${\mathcal{O}}$ if it is a weight module, and the action of $U_{q}{({{\mathfrak{n}}}^{+})}{}$ is locally finite.151515 This implies in particular that, for any $v\in V$, $U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}v=0$ for all but finitely many $\beta\in{\mathsf{Q}}^{+}$. Therefore, ${\mathcal{O}}$ coincides with the category ${\mathcal{C}}^{\operatorname{hi}}$ used in [Lus94, Sec. 3.4.7].
We denote by $\mathcal{W}$, ${\mathcal{O}}$, ${\mathcal{O}}_{{\operatorname{int}}}$ the full subcategories of weight, category ${\mathcal{O}}$, integrable category ${\mathcal{O}}$ modules in ${\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$, respectively. The category $\mathcal{W}$ is only tensor, while the category ${\mathcal{O}}$ is braided, see [Dri86, Lus94]. The category ${\mathcal{O}}_{{\operatorname{int}}}$ is a semisimple braided subcategory of ${\mathcal{O}}$, whose simple modules are classified by dominant integral weights [Lus94, Thm 6.2.2, Cor. 6.2.3].
Finally, let $Z\subseteq I$ be any subset and $U_{q}({\mathfrak{g}}_{Z})\subseteq U_{q}({\mathfrak{g}})$ the subalgebra generated by $\{E_{i},F_{i},K_{i}^{\pm 1}\}_{i\in Z}$. We denote by ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}\subseteq{\mathcal{O}}$ the full subcategory of modules whose restriction to $U_{q}({\mathfrak{g}}_{Z})$ is integrable. In particular, ${\mathcal{O}}_{{{\emptyset}\negthinspace\operatorname{-int}}}={\mathcal{O}}$ and ${\mathcal{O}}_{{{I}\negthinspace\operatorname{-int}}}={\mathcal{O}}_{{\operatorname{int}}}$.
2.4. Completions
Let $A$ be an algebra, ${\mathcal{C}}\subset{\operatorname{Mod}}(A)$ a (full) subcategory, and $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ the algebra of endomorphisms of the forgetful functor ${\mathsf{f}}_{{\mathcal{C}}}\colon{\mathcal{C}}\to\operatorname{Vect}$. By definition, an element of $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ is a collection
$\varphi=\{\varphi_{V}\}_{V\in{\mathcal{C}}}\in\prod_{V\in{\mathcal{C}}}\operatorname{End}(V)$
which is natural, i.e., such that $f\circ\varphi_{V}=\varphi_{W}\circ f$ for any $f:V\to W$ in ${\mathcal{C}}$. The action of $A$ on any $V\in{\mathcal{C}}$ yields a morphism of algebras $A\to\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$, and factors through the action of $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ on $V$. We refer to $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ as the completion of $A$ with respect to the category ${\mathcal{C}}$. If $A$ is a bialgebra and ${\mathcal{C}}$ is closed under tensor products, the completion $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ is naturally endowed with a cosimplicial structure, see, e.g., [Dav07, ATL19].
For any subcategory ${\mathcal{C}}^{\prime}\subset{\mathcal{C}}$, there is a canonical morphism
$\eta\colon\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})\to\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}^{\prime}})$ (2.9)
given by the restriction to ${\mathcal{C}}^{\prime}$. A lift of $x\in\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}^{\prime}})$ is any element in $\eta^{-1}(x)\subset\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$. Since $\eta$ is in general neither surjective nor injective, a lift may not exist or be unique. However, in the following, we shall often consider operators whose action is given in terms of elements of $A$. Hence, if it exists, the lift is canonical (see, e.g.,  Section 2.6). In these cases, we shall make no distinction between an element and its canonical lift.
2.5. Completions of $U_{q}({\mathfrak{g}})^{\otimes n}$
We consider various completions of $U_{q}({\mathfrak{g}})^{\otimes n}$. Namely, for any $Z\subseteq I$ and $n>0$, let ${\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{n}\colon({\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}})^{\boxtimes n}\to\operatorname{Vect}$ be the $n$-fold forgetful functor given by
${\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{n}(V_{1},\dots,V_{n})=V_{1}\otimes\cdots\otimes V_{n}.$
We set ${\mathsf{f}}^{n}={\mathsf{f}}_{\emptyset}^{n}$, ${\mathsf{f}}_{{{I}\negthinspace\operatorname{-int}}}^{n}={\mathsf{f}}_{{\operatorname{int}}}^{n}$, and ${\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}={\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{1}$. Similarly, let ${\mathsf{f}}_{\mathcal{W}}^{n}:\mathcal{W}^{\boxtimes n}\to\operatorname{Vect}$ be the $n$-fold forgetful functor with respect to the category of weight $U_{q}({\mathfrak{g}})$-modules. Note that the chain of inclusions ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}\subset\mathcal{W}\subset{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$ induces a chain of morphisms of algebras $U_{q}({\mathfrak{g}})^{\otimes n}\to\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{n})\to\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{n})$. By [ATL24, Thm. 3.1], $U_{q}({\mathfrak{g}})^{\otimes n}$ embeds in $\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}^{n})$ and therefore in $\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{n})$.
2.6. The R-matrix
By [Dri86, Lus94], the universal R-matrix of $U_{q}({\mathfrak{g}})$ is a distinguished invertible element ${R}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$, whose action induces a braiding on the category ${\mathcal{O}}$. More precisely, ${R}$ satisfies the intertwining identity in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$
$R\cdot\Delta(x)=\Delta^{\operatorname{op}}(x)\cdot R,$ (2.10)
where $x\in U_{q}({\mathfrak{g}})$, and the coproduct identities in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}\boxtimes{\mathsf{f}})$
$(\Delta\otimes{\operatorname{id}})(R)=R_{13}\cdot R_{23}\qquad\mbox{and}\qquad({\operatorname{id}}\otimes\Delta)(R)=R_{13}\cdot R_{12},$ (2.11)
where $\Delta$ denotes the coproduct and $\Delta^{\operatorname{op}}=(12)\circ\Delta$. More precisely, the morphisms
$\Delta,\Delta^{\operatorname{op}}\colon\operatorname{End}({\mathsf{f}})\to\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$
are defined as follows. For any $f\in\operatorname{End}({\mathsf{f}})$, we set
$\Delta(f)_{VW}=f_{V\otimes W}\quad\mbox{and}\quad\Delta^{\operatorname{op}}(f)_{VW}=(1\,2)\circ f_{W\otimes V}\circ(1\,2)$
where $V,W\in{\mathcal{O}}$. The morphisms $\Delta\otimes{\operatorname{id}}$ and ${\operatorname{id}}\otimes\Delta$ are similarly defined. From (2.10) and (2.11) it readily follows that $R$ is a solution of the Yang-Baxter equation
$R_{12}\cdot R_{13}\cdot R_{23}=R_{23}\cdot R_{13}\cdot R_{12}.$ (2.12)
in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}\boxtimes{\mathsf{f}})$.
The universal R-matrix arises from the Drinfeld double construction of $U_{q}({\mathfrak{g}})$ as the canonical tensor of a Hopf pairing between $U_{q}{({\mathfrak{b}}^{-})}$ and $U_{q}{({\mathfrak{b}}^{+})}$. More explicitly,
$R=\operatorname{C}^{{\operatorname{id}}}\cdot\Theta$ (2.13)
where
  • $\operatorname{C}^{{\operatorname{id}}}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{2})$ is a diagonal operator, acting on a tensor product of weight spaces $V_{\lambda}\otimes W_{\mu}$ as multiplication by $q^{(\lambda,\mu)}$, see also Section 5.3;
  • $\Theta=\sum_{\beta\in\mathsf{Q}^{+}}\Theta_{\beta}\in U_{q}{({{\mathfrak{n}}}^{-})}\widehat{\otimes}U_{q}{({{\mathfrak{n}}}^{+})}$, where
    $U_{q}{({{\mathfrak{n}}}^{-})}\widehat{\otimes}U_{q}{({{\mathfrak{n}}}^{+})}=\prod_{\beta}U_{q}{({{\mathfrak{n}}}^{-})}_{-\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}$
    and $\Theta_{\beta}\in U_{q}{({{\mathfrak{n}}}^{-})}_{-\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}$ is the canonical element with respect to the Hopf pairing.
The element $\Theta$ is commonly referred to as the quasi R-matrix [Lus94, Ch. 4]. Relying on the condition (C3), every element in $U_{q}{({{\mathfrak{n}}}^{-})}\widehat{\otimes}U_{q}{({{\mathfrak{n}}}^{+})}$ has a well-defined and natural action on any tensor product of the form $V\otimes W$ with $V\in{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$ and $W\in{\mathcal{O}}$. More precisely, let $\widetilde{{\mathsf{f}}}\colon{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))\to\operatorname{Vect}$ be the forgetful functor. Following [ATL24, Sec. 3], one proves that the canonical map
${U_{q}{({{\mathfrak{n}}}^{-})}\widehat{\otimes}U_{q}{({{\mathfrak{n}}}^{+})}}$${\operatorname{End}(\widetilde{{\mathsf{f}}}\boxtimes{\mathsf{f}})}$ (2.14)
is an embedding. In particular, $\Theta\in\operatorname{End}(\widetilde{{\mathsf{f}}}\boxtimes{\mathsf{f}})$. Since the action of $\operatorname{C}^{{\operatorname{id}}}$ is defined only on weight modules, it follows that $R\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}})$. Then, it defines a braiding on ${\mathcal{O}}$ through the morphism
${\operatorname{End}({{\mathsf{f}}_{\mathcal{W}}}\boxtimes{\mathsf{f}})}$${\operatorname{End}({\mathsf{f}}^{2})}$ (2.15)
induced by the inclusion ${\mathcal{O}}\subset\mathcal{W}$.
3. Quantum symmetric pairs
We recall the definition of quantum symmetric pairs for quantum Kac-Moody algebras following [Kol14, RV20, AV22a].
3.1. Pseudo-involutions
Let $\operatorname{Aut}({A})$ be the group of diagram automorphisms of the Cartan matrix, i.e., the group of permutations $\tau$ of $I$ such that $a_{ij}=a_{\tau(i)\tau(j)}$. Any $\tau\in\operatorname{Aut}(A)$ canonically defines an automorphism of ${{\mathfrak{g}}^{\prime}}$, given on the generators by $\tau(h_{i})=h_{\tau(i)}$, $\tau(e_{i})=e_{\tau(i)}$, and $\tau(f_{i})=f_{\tau(i)}$. In [KW92, Sec. 4.9], Kac and Wang describe a procedure (depending on the choice of the subspace ${\mathfrak{h}}^{\prime\prime}\subset{\mathfrak{h}}$) to extend a diagram automorphism from ${{\mathfrak{h}}^{\prime}}$ to ${{\mathfrak{h}}}$. Then, we assume that the basis $\{d_{1},\dots,d_{\operatorname{cork}(A)}\}$ of ${\mathfrak{h}}^{\prime\prime}$ is chosen in such a way that the Kac-Wang extension of $\tau$ on ${\mathfrak{h}}$ preserves the extended root lattice $\mathsf{Q}^{\vee}_{\sf ext}$ and, by duality, the weight lattice $\mathsf{P}$.
Remark 3.1.1.
The problem of the existence of a basis of ${\mathfrak{h}}^{\prime\prime}$ compatible with a given $\tau\in\operatorname{Aut}(A)$ is discussed in [AV22a, Prop. 3.2] in the case $\operatorname{cork}(A)=1$. If $A$ is of affine type, such a basis exists for all $\tau$, see also [Kol14, Prop. 2.12]. $\triangledown$
Let $X\subset I$ be a subset of indices such that the corresponding Cartan matrix $A_{X}$ is of finite type. We denote by ${\mathfrak{g}}_{X}\subset{\mathfrak{g}}$ the corresponding semisimple finite-dimensional Lie subalgebra. Let $\tau$ be an involutive diagram automorphism stabilizing $X$ such that $\tau|_{X}=\mathsf{oi}_{X}$, the involutive diagram automorphism of $X$ induced by the action of the longest element $w_{X}$ of the Weyl group $W_{X}$ on $\mathsf{Q}_{X}$. We consider the Lie algebra automorphism of ${\mathfrak{g}}$ given by
$\theta=\operatorname{Ad}(w_{X})\circ\omega\circ\tau$ (3.1)
where $\omega$ denotes the Chevalley involution on ${\mathfrak{g}}$ and $\operatorname{Ad}(w)$ for $w\in W$ is the corresponding braid group action realized in terms of triple exponentials. Note that $\operatorname{Ad}(w_{X})^{2}$ acts on the root space ${\mathfrak{g}}_{\lambda}$ ($\lambda\in\mathsf{Q}$) as multiplication by $(-1)^{\lambda(2\rho^{\vee}_{X})}$, with $\rho^{\vee}_{X}$ the half-sum of the positive coroots associated to ${\mathfrak{g}}_{X}$. Then $\theta$ fixes ${\mathfrak{g}}_{X}$ pointwise and exchanges positive and negative root spaces not contained in ${\mathfrak{g}}_{X}$. Further, $\theta$ restricts to ${\mathfrak{h}}$ as the involutive map $-w_{X}\circ\tau$; we use the notation $\theta$ also to denote the map on ${\mathfrak{h}}^{*}$ dual to $\theta|_{\mathfrak{h}}$, so that $\theta\big{(}{\mathfrak{g}}_{\alpha}\big{)}={\mathfrak{g}}_{\theta(\alpha)}$ for all $\alpha\in{\bm{\Phi}}$.
We call $\theta$ a pseudo-involution (of the second kind). The datum $(X,\tau)$ can be recovered from $\theta$ since $X=\{i\in I\,|\,\theta(h_{i})=h_{i}\}$ and hence we will use the subscript $\theta$ for objects explicitly defined in terms of $(X,\tau)$.
Any ${\bm{y}}\in({\mathbb{C}}^{\times})^{I}$ may be viewed as a multiplicative character ${\bm{y}}:\mathsf{Q}\to{\mathbb{C}}^{\times}$ according to ${\bm{y}}(\alpha_{i})=y_{i}$ for $i\in I$. In turn, it gives rise to a Lie algebra automorphism $\operatorname{Ad}(\bm{y})$ of ${\mathfrak{g}}$ which acts on the root space ${\mathfrak{g}}_{\alpha}$ as multiplication by $\bm{y}(\alpha)$. For ${\bm{y}}\in({\mathbb{C}}^{\times})^{I}$ such that $y_{i}=1$ if $i\in X$, consider the following modification of $\theta$, which coincides with $\theta$ on $\langle{\mathfrak{g}}_{X},{\mathfrak{h}}\rangle$:
$\phi=\theta\circ\operatorname{Ad}({\bm{y}})^{-1}.$ (3.2)
3.2. Pseudo-fixed-point subalgebras
Let ${\mathfrak{k}}\subset{\mathfrak{g}}$ be the Lie subalgebra generated by ${\mathfrak{n}}^{+}_{X}$, ${\mathfrak{h}}^{\theta}$ and the elements
$f_{i}+\phi(f_{i})=f_{i}+y_{i}\theta(f_{i})$
for $i\in I$. We set ${\mathfrak{k}}^{\prime}={\mathfrak{k}}\cap{\mathfrak{g}}^{\prime}$.
Henceforth we assume that $(X,\tau)$ is a generalized Satake diagram, i.e., for all $i,j\in I$ such that $\theta(\alpha_{i})=-\alpha_{i}-\alpha_{j}$ we have $a_{ij}\neq-1$. Further, we assume that, for all $i\in I$,
$y_{i}=y_{\tau(i)}\qquad\text{if}\qquad(\alpha_{i},\theta(\alpha_{i}))=0.$ (3.3)
With these assumptions, the Lie subalgebra ${\mathfrak{k}}$ resembles the fixed-point subalgebra of an involution in the sense that ${\mathfrak{k}}\cap{\mathfrak{h}}={\mathfrak{h}}^{\theta}$, see [RV20, RV22, AV22a], and we call it a pseudo-fixed-point subalgebra.
3.3. Involutions and fixed-point subalgebras
The genuinely involutive case is described by the following refinement, cf. [Kol14, Sec. 2.4].
Lemma 3.3.1.
The following conditions are equivalent:
  1. (1)
    $\phi$ is an involution;
  2. (2)
    $y_{i}y_{\tau(i)}^{-1}=(-1)^{\alpha_{i}(2\rho^{\vee}_{X})}$ for all $i\in I\setminus X$;
  3. (3)
    ${\mathfrak{k}}$ is the fixed-point subalgebra ${\mathfrak{g}}^{\phi}$.
Note that condition (2) above implies that $\alpha_{i}(\rho^{\vee}_{X})\in{\mathbb{Z}}$ whenever $i=\tau(i)\in I\backslash X$, i.e., $(X,\tau)$ is a Satake diagram as defined in [Kol14, Def. 2.3].
Proof of Lemma 3.3.1..
Since $\theta^{2}=\operatorname{Ad}(w_{X})^{2}$, by direct inspection, $\phi$ is an involutive automorphism if and only if (2) holds. In this case, [Kol14, Lem. 2.8] implies that ${\mathfrak{k}}={\mathfrak{g}}^{\phi}$. Conversely, if ${\mathfrak{k}}$ is contained in ${\mathfrak{g}}^{\phi}$ then we must have $\phi^{2}(f_{i})=f_{i}$ for all $i\in I$ so that $\phi^{2}$ fixes ${\mathfrak{n}}^{-}$ pointwise. Noting that $\phi|_{\mathfrak{h}}=-w_{X}\circ\tau$ is an involution, we deduce that $\operatorname{ad}(f_{i})$ annihilates $e_{j}-\phi^{2}(e_{j})\in{\mathfrak{n}}^{+}$ for all $i,j\in I$. Using [Kac90, Lemma 1.5] we conclude that $\phi^{2}(e_{j})=e_{j}$ for all $j\in I$. Hence $\phi$ is an involution. ∎
3.4. Quantum pseudo-involutions
We now describe a convenient algebra automorphism $\theta_{q}$ of $U_{q}({\mathfrak{g}})$ which quantizes $\theta$. This is obtained by choosing a lift for each of the three factors in $\theta$. First, we consider the standard Chevalley involution on $U_{q}({\mathfrak{g}})$ given by (2.8). The diagram automorphism $\tau$ extends to an automorphism of $U_{q}({\mathfrak{g}})$ given on the generators by $\tau(E_{i})=E_{\tau(i)}$, $\tau(F_{i})=F_{\tau(i)}$, and $\tau(K_{h})=K_{\tau(h)}$.
The action of the Weyl group operator $w_{X}\in W_{X}$ is lifted to $U_{q}({\mathfrak{g}})$ as follows. Let $S_{X}$ be the braid group operator on integrable category ${\mathcal{O}}$ modules corresponding to $w_{X}$ [Lus94, Sec. 5]. More precisely, given a reduced expression $s_{i_{1}}\cdots s_{i_{\ell}}$ of $w_{X}$ in terms of fundamental reflections, one sets $S_{X}=S_{i_{1}}\cdots S_{i_{\ell}}$, where $S_{j}=T^{\prime\prime}_{j,1}$ in the notation from [Lus94, 5.2.1]. It follows from the braid relations that $S_{X}$ is independent of the chosen reduced expression.
In [AV22a, Sec. 4.9], we introduced a Cartan correction of $S_{X}$ given by
$\mathcal{S}_{\theta}=\kappa_{\theta}\cdot S_{X}$ (3.4)
Here, $\kappa_{\theta}$ is the operator defined on any weight vector of weight $\lambda$ as multiplication by $q^{\left(\theta(\lambda),\lambda\right)/2+\left(\lambda,\rho_{X}\right)}$, with $\rho_{X}$ the half-sum of the positive roots in ${\bm{\Phi}}_{X}$, the root system of the finite-dimensional semisimple Lie algebra ${\mathfrak{g}}_{X}$. Thus, $\mathcal{S}_{\theta}$ can be thought of as an element of $\operatorname{End}({\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$.
Remark 3.4.1.
In [AV22a, Sec. 6.6], we also introduced a Cartan correction of the universal R-matrix $R_{X}$ of $U_{q}({\mathfrak{g}}_{X})$ given by
$R_{\theta}=\operatorname{C}^{\theta}\cdot\Theta_{X}\in\operatorname{End}({\mathsf{f}}^{2})$ (3.5)
where $\Theta_{X}$ is the quasi R-matrix of $U_{q}({\mathfrak{g}}_{X})$. Then, by [AV22a, Lemma 6.8],
$R_{\theta}=(\mathcal{S}_{\theta}\otimes\mathcal{S}_{\theta})^{-1}\Delta(\mathcal{S}_{\theta})$ (3.6)
in $\operatorname{End}({\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}}^{2})$. $\triangledown$
The operator $\mathcal{S}_{\theta}$ yields an algebra automorphism $\mathcal{T}_{\theta}=\operatorname{Ad}{(\mathcal{S}_{\theta})}$ of $\operatorname{End}({\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$, which preserves $U_{q}({\mathfrak{g}})$ [AV22a, Lem. 4.3 and 5.2]. Set
$\theta_{q}=\mathcal{T}_{\theta}\circ\omega\circ\tau$ (3.7)
As an algebra automorphism of $U_{q}({\mathfrak{g}})$, $\theta_{q}$ is the identity on $U_{q}({\mathfrak{g}}_{X})$, $\theta_{q}(K_{h})=K_{\theta(h)}$ for all $h\in\mathsf{Q}^{\vee}_{\sf ext}$ and $\theta_{q}\big{(}U_{q}({\mathfrak{g}})_{\lambda}\big{)}=U_{q}({\mathfrak{g}})_{\theta(\lambda)}$ for all $\lambda\in\mathsf{Q}$, see, e.g., [AV22a, Sec. 6.7].
We consider the tuple $\bm{\gamma}\in({\mathbb{F}}^{\times})^{I}$ as a multiplicative character $\bm{\gamma}:\mathsf{Q}\to{\mathbb{F}}^{\times}$, yielding a Hopf algebra automorphism $\operatorname{Ad}(\bm{\gamma})$ of $U_{q}({\mathfrak{g}})$ which acts on the root space ${\mathfrak{g}}_{\alpha}$ as multiplication by $\bm{\gamma}(\alpha)$. In analogy with (3.2), we set
$\phi_{q}=\theta_{q}\circ\operatorname{Ad}(\bm{\gamma}^{-1})$ (3.8)
so that $\phi_{q}(F_{i})=\bm{\gamma}_{i}\theta_{q}(F_{i})$.
3.5. QSP subalgebras
Let $\bm{\Gamma}\subset({\mathbb{F}}^{\times})^{I}$ and $\bm{\Sigma}\subset{\mathbb{F}}^{I}$ be the parameter sets defined in [AV22a, Def. 6.11].. The QSP subalgebra associated to $\theta$ with parameters $(\bm{\gamma},\bm{\sigma})\in\bm{\Gamma}\times\bm{\Sigma}$ is the subalgebra $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ generated by the subalgebra $U_{q}({\mathfrak{n}}^{+}_{X}):=\langle E_{j}\rangle_{j\in X}$, the commutative subalgebra $U_{q}({\mathfrak{h}}^{\theta}):=\langle K_{h}\;|\;h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}\rangle$, and the elements elements $B_{i}$ ($i\in I$), given by
$B_{i}=\left\{\begin{array}[]{cl}F_{i}&\mbox{if }\,i\in X,\\ F_{i}+\phi_{q}(F_{i})+\bm{\sigma}_{i}\,K_{i}^{-1}&\mbox{if }\,i\not\in X.\end{array}\right.$ (3.9)
We set $U_{q}({\mathfrak{k}^{\prime}})=U_{q}(\mathfrak{k})\cap U_{q}({\mathfrak{g}}^{\prime})$.
Remarks 3.5.1.
  1. (1)
    Up to reparametrization, the definition of $U_{q}(\mathfrak{k})$ coincides with that given in [Let02, Kol14], see [AV22a, Rmk. 6.13].
  2. (2)
    It is known that $U_{q}(\mathfrak{k})$ is a right coideal, i.e., $\Delta(U_{q}(\mathfrak{k}))\subset U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{g}})$, with the additional property $U_{q}(\mathfrak{k})\cap U_{q}({\mathfrak{h}})=U_{q}({\mathfrak{h}}^{\theta})$; $U_{q}({\mathfrak{k}^{\prime}})$ satisfies analogous properties with respect to $U_{q}({\mathfrak{g}}^{\prime})$ and $U_{q}(({\mathfrak{h}}^{\prime})^{\theta})$, see [Kol14, AV22a].
  3. (3)
    If $\operatorname{cork}(A)\leqslant 1$ then $U_{q}({\mathfrak{k}^{\prime}})=U_{q}(\mathfrak{k})$, see [AV22a, Sec. 6.2]. $\triangledown$
3.6. Auxiliary R-matrices
Fix an extension of $\bm{\gamma}:\mathsf{Q}\to{\mathbb{F}}^{\times}$ to a character $\bm{\gamma}:\mathsf{P}\to{\mathbb{F}}^{\times}$. We consider the diagonal operator $\bm{\gamma}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}})$ acting on any weight space $V_{\lambda}$ as multiplication by $\bm{\gamma}(\lambda)$. Moreover, one has $\Delta(\bm{\gamma})=\bm{\gamma}\otimes\bm{\gamma}$ in $\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}}_{\mathcal{W}})$. Finally, $\operatorname{Ad}(\bm{\gamma})$ is an algebra automorphism of $\operatorname{End}({\mathsf{f}}_{\mathcal{W}})$, which restricts to a Hopf algebra automorphism of $U_{q}({\mathfrak{g}})$.
For any $V\in{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$, and any algebra automorphism $\psi:U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})$, let ${V^{\psi}}=\psi^{*}(V)\in{\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$ be the pullback of $V$ under $\psi$.
Lemma 3.6.1.
In the case $\psi=\operatorname{Ad}(\bm{\gamma})\circ\theta_{q}^{-1}=\phi_{q}^{-1}$ the following statements hold.
  1. (1)
    For any $V\in\mathcal{W}$, ${V^{\psi}}\in\mathcal{W}$.
  2. (2)
    For any $V\in{\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}{}$, ${V^{\psi}}$ is in the opposite parabolic category ${\mathcal{O}}$, i.e., ${V^{\psi}}$ is a weight module which is locally finite over $U_{q}{({{\mathfrak{n}}}^{-})}$ and integrable over $U_{q}({\mathfrak{g}}_{X})$.
  3. (3)
    For any $V,W\in{\mathcal{O}}$, the operator
    $J=(\bm{\gamma}\otimes\bm{\gamma})\cdot R_{\theta}\cdot\Delta(\bm{\gamma})^{-1}=R_{\theta}\in\operatorname{End}({\mathsf{f}}^{2})$
    yields an intertwiner
    $J_{VW}\colon(V\otimes W)^{\psi}\to{V^{\psi}}\otimes^{\operatorname{op}}{W^{\psi}},$
    where $\otimes^{\operatorname{op}}$ indicates that the action of $U_{q}({\mathfrak{g}})$ is given by the opposite coproduct $\Delta^{\operatorname{op}}$.
  4. (4)
    For any $V,W\in{\mathcal{O}}$, the action of the universal R-matrix on ${V^{\psi}}\otimes{W^{\psi}}$ is well-defined and it is given by
    $R_{{V^{\psi}}{W^{\psi}}}=J_{VW}\circ(R_{21})_{VW}\circ(J_{21})_{VW}^{-1},$ (3.10)
[…]
Proof.
(1) It is enough to observe that, for any $h\in\mathsf{Q}^{\vee}_{\sf ext}$, $\psi(K_{h})=K_{\theta(h)}$. Therefore, ${V^{\psi}}\in\mathcal{W}$.
(2) Since $\theta_{q}=\mathcal{T}_{\theta}\circ\omega\circ\tau$, it is enough to observe that the restriction under $\mathcal{T}_{\theta}$ preserves $X$-integrable category ${\mathcal{O}}$ modules, see e.g., [AV22a, Lemma 5.2].
(3)(4) See [AV22a, Prop. 8.7]. In particular, note that
$(\bm{\gamma}\otimes\bm{\gamma})\cdot R_{\theta}=R_{\theta}\cdot(\bm{\gamma}\otimes\bm{\gamma})=R_{\theta}\cdot\Delta(\bm{\gamma})$ (3.11)
since $R_{\theta}$ is a weight preserving operator and $\bm{\gamma}$ is grouplike. ∎
Remark 3.6.2.
The identity (3.10) is somewhat surprising. Namely, by (2), the operator $R_{{V^{\psi}}{W^{\psi}}}$ is well-defined whenever $V\in{\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}{}$, since the locally finite action of $U_{q}{({{\mathfrak{n}}}^{-})}$ guarantees the convergence of $\Theta$, see Section 2.6. Instead, it is not clear a priori why the operator is well-defined even for category ${\mathcal{O}}$ modules which are not $X$-integrable. $\triangledown$
Proposition 3.6.3.
  1. (1)
    There is a well-defined operator ${R}^{\psi}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}})$ given by the collection of maps
    ${R}_{VW}^{\psi}={R}_{{V^{\psi}}W}\in\operatorname{End}(V\otimes W)$
    where $V\in\mathcal{W}$ and $W\in{\mathcal{O}}$.
  2. (2)
    There is a well-defined operator ${R}_{21}^{\psi}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}_{\mathcal{W}})$ given by the collection of maps
[…]
    where $V,W\in{\mathcal{O}}$.
  3. (3)
    There is a well-defined operator ${R}^{\psi\psi}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ given by the collection of maps
    ${R}_{VW}^{\psi\psi}={R}_{{V^{\psi}}{W^{\psi}}}\in\operatorname{End}(V\otimes W)$
    where $V,W\in{\mathcal{O}}$.
Proof.
(1) By 2.6, $R\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}})$. Thus, by Lemma 3.6.1 (1), it follows that ${R}_{{V^{\psi}}W}$ is well-defined and natural in both $V$ and $W$.
(2) As in (1), it is enough to observe that
[…]
and $R_{21}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}_{\mathcal{W}})$.
(3) It follows readily from Lemma 3.6.1 (3). ∎
3.7. The quasi K-matrix
The QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})$ gives rise to a family of basic K-matrices in $U_{q}({\mathfrak{g}})$, see [AV22a, Thms. 8.8-8.9]. The result stems from the construction of the quasi K-matrix due to Bao and Wang [BW18b], subsequently generalized by Balagović and Kolb in [BK19] to the symmetrizable Kac-Moody setting. In [AV22a] it was shown that in this construction the parameters $(\bm{\gamma},\bm{\sigma})\in\bm{\Gamma}\times\bm{\Sigma}$ need not be constrained.
Theorem 3.7.1.
[AV22a, Thms. 8.8-8.9] There exists a unique element $\Upsilon\in\operatorname{End}({\mathsf{f}})$ of the form
$\Upsilon=1+\sum_{\begin{subarray}{c}\mu\in(\mathsf{Q}^{+})^{-\theta}\\ \mu\neq 0\end{subarray}}\Upsilon_{\mu}\in\prod_{\mu\in\mathsf{Q}^{+}}U_{q}{({{\mathfrak{n}}}^{+})}_{\mu}$ (3.14)
which satisfies
  • the intertwining identity
    $K\cdot b=\psi(b)\cdot K\,\qquad\text{for all }b\in U_{q}({\mathfrak{k}^{\prime}}),$ (3.15)
  • the coproduct identity
    $\Delta(K)=J^{-1}\cdot 1\otimes K\cdot{R}^{\psi}\cdot\Upsilon\otimes 1,$ (3.16)
  • Cherednik’s generalized reflection equation
    $R^{\psi\psi}_{21}\cdot(1\otimes K)\cdot R^{\psi}\cdot(K\otimes 1)=(K\otimes 1)\cdot R^{\psi}_{21}\cdot(1\otimes K)\cdot R,$ (3.17)
where
$\psi=\operatorname{Ad}(\bm{\gamma})\circ\theta_{q}^{-1}=\phi_{q}^{-1},\quad J=R_{\theta},\quad\mbox{and}\quad K=\Upsilon.$ (3.18)
Remarks 3.7.2.
  1. (1)
    By the weight decomposition of $\Upsilon$, one has $u\Upsilon=\Upsilon u$ for any $u\in U_{q}({\mathfrak{h}}^{\theta})$. Thus, the identity (3.15) for $U_{q}({\mathfrak{k}^{\prime}})$ implies the same identity for $U_{q}(\mathfrak{k})$ which appears in [AV22a, Thms. 8.8-8.9].
  2. (2)
    By Proposition 3.6.3, the identities (3.16) and (3.17) are in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$. $\triangledown$
In [AV22a], we introduced the notion of a cylindrical structure, encoding the formal properties of the datum $(\psi,J,K)$, see Section B.2. Since the action of the quasi K-matrix is defined at the level of category ${\mathcal{O}}$ modules, the result above can be rephrased as follows.
Corollary 3.7.3.
The datum $(\psi,J,K)=(\phi_{q}^{-1},R_{\theta},\Upsilon)$ is a cylindrical structure on $U_{q}({\mathfrak{g}})$ defined over ${\mathcal{O}}$ with respect to the QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})$.
3.8. Gauged K-matrices
In [AV22a, AV22b], we described a procedure to obtain new universal solutions of the generalized reflection equation (3.17) by gauging the quasi K-matrix, see [AV22b, Cor. 3.6].
Proposition 3.8.1.
Fix a subset $Z\subseteq I$. Let ${\mathcal{G}}_{Z}$ be the group of invertible elements ${\bf g}\in\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}})$ such that $\operatorname{Ad}({\bf g})$ preserves $U_{q}({\mathfrak{g}})\subset\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}})$. For any ${\bf g}\in{\mathcal{G}}_{Z}$, the datum $(\psi,J,K)$ given by
$\psi=\operatorname{Ad}({\bf g}\bm{\gamma})\circ\thinspace\theta_{q}^{-1},\quad J=({\bf g}\otimes{\bf g})\cdot R_{\theta}\cdot\Delta({\bf g})^{-1}\quad\mbox{and}\quad K={\bf g}\cdot\Upsilon$ (3.19)
satisfies the identities (3.15), (3.16), and (3.17). In particular, $(\psi,J,K)$ is a cylindrical structure on $U_{q}({\mathfrak{g}})$ defined over ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}{}$ with respect to the QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})$.
Proof.
It is enough to observe that the results of Lemma 3.6.1 and Proposition 3.6.3 remain valid for this choice of $\psi$ and $J$, up to replacing ${\mathsf{f}}$ with ${\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}$. ∎
4. The classical limit of universal K-matrices
In this section we discuss specialization. In the formal ($\hbar$-adic) setting, it is clear that the specializations below are well-defined. We will use the formalism for specialization as discussed in [Wa21, Sec. 2.4] via subrings of ${\mathbb{F}}$ defined in terms of certain Puiseux series. The main result of [Kol14, Sec. 10], characterizing $U_{q}({\mathfrak{k}^{\prime}})$ as the maximal subspace of $U_{q}({\mathfrak{g}}^{\prime})$ whose specialization is $U({\mathfrak{k}}^{\prime})$, directly carries over to this setting.
In this section we rely on this result to characterize $U_{q}({\mathfrak{k}^{\prime}})$ as the maximal right coideal subalgebra of $U_{q}({\mathfrak{g}}^{\prime})$ contained in a fixed-point subalgebra determined by the quasi K-matrix.
4.1. Maximality for QSP subalgebras
We denote by $\operatorname{cl}(x)$ the specialization of $x$ at $q=1$ whenever well-defined. Henceforth, we assume that the QSP parameters $(\bm{\gamma},\bm{\sigma})\in\bm{\Gamma}\times\bm{\Sigma}$ are specializable. Note that, unlike in [Kol14], we do not assume that $\operatorname{cl}(\gamma_{i})=1$.
Note that in Section 3.2 we defined subalgebras ${\mathfrak{k}}\subseteq{\mathfrak{g}}$ and ${\mathfrak{k}}^{\prime}={\mathfrak{k}}\cap{\mathfrak{g}}^{\prime}$ depending on a tuple $\bm{y}\in({\mathbb{C}}^{\times})^{I}$ satisfying (3.3). We recall the following characterization of $U_{q}(\mathfrak{k})^{\prime}$.
Theorem 4.1.1.
[Kol14, RV20] The QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})\subset U_{q}({\mathfrak{g}}^{\prime})$ is a maximal subspace such that $\operatorname{cl}(U_{q}({\mathfrak{k}^{\prime}}))=U({\mathfrak{k}}^{\prime})$ with $\operatorname{cl}(\bm{\gamma})=\bm{y}$.
Proof.
In the involutive case, we refer to [AV22a, Rmk. 6.13] for the equality between $U_{q}({\mathfrak{k}^{\prime}})$ and Kolb’s subalgebra $B_{{\bf c},{\bf s}}$. In this case, the statement is proved in [Kol14, Thms. 10.8, 10.11]. The proofs of loc. cit. naturally extend to generalized Satake diagrams, as was pointed out in [RV20]. In the same way, the proofs are still valid for general specializable $\bm{\gamma}\in\Gamma$, the difference being that the procedure yields the universal enveloping algebra of a Lie subalgebra ${\mathfrak{k}}^{\prime}\subset{\mathfrak{g}}^{\prime}$ depending on $\bm{y}=\operatorname{cl}(\bm{\gamma})$. ∎
Let $(\psi,J,K)$ be as in Proposition 3.8.1. We consider the automorphism of the completion $\operatorname{End}({\mathsf{f}}_{{\operatorname{int}}})$ given by161616 There is a priori no reason why $U_{q}({\mathfrak{g}})$ should be $\operatorname{Ad}(\Upsilon)$-stable. In fact, it is easy to show that this is false even for ${\mathfrak{g}}=\mathfrak{sl}_{2}$ (see [DK17, Lemma 3.8]).
$\xi=\operatorname{Ad}(K)^{-1}\circ\psi=\operatorname{Ad}(\Upsilon)^{-1}\circ\phi_{q}^{-1},$ (4.1)
where we recall $\phi_{q}^{-1}=\operatorname{Ad}(\bm{\gamma})\circ\theta_{q}^{-1}$. Then, we consider
$B^{\prime}_{\xi}=\{x\in U_{q}({\mathfrak{g}}^{\prime})\,|\,(\xi\otimes{\operatorname{id}})(\Delta(x))=\Delta(x)\}\subseteq U_{q}({\mathfrak{g}}^{\prime})^{\xi}.$ (4.2)
One checks that $B^{\prime}_{\xi}$ is a right coideal subalgebra of $U_{q}({\mathfrak{g}}^{\prime})$, maximal among all those contained in $U_{q}({\mathfrak{g}}^{\prime})^{\xi}$. Hence, the QSP intertwining identity (3.15) implies that the right coideal subalgebra $U_{q}({\mathfrak{k}^{\prime}})$ satisfies $U_{q}({\mathfrak{k}^{\prime}})\subseteq B^{\prime}_{\xi}\subset U_{q}({\mathfrak{g}}^{\prime})^{\xi}$.
Theorem 4.1.2.
In fact, we have $U_{q}({\mathfrak{k}^{\prime}})=B^{\prime}_{\xi}$.
The proof is carried out in Section 4.3. It relies on the classical limit of the inclusion $U_{q}({\mathfrak{k}^{\prime}})\subseteq B^{\prime}_{\xi}$ and the properties of the classical limit of $\Upsilon$.
4.2. Classical limit of the quasi K-matrix
We shall now study the classical limit of $\Upsilon$. Note that $\operatorname{cl}(\theta_{q})=\theta$. Since $\operatorname{Ad}(\bm{\gamma})\in\operatorname{Aut}_{\sf alg}(U_{q}({\mathfrak{g}}))$ only depends on the values $\bm{\gamma}_{i}$, not on the choice of extension of $\bm{\gamma}:\mathsf{Q}\to{\mathbb{F}}^{\times}$ to a group homomorphism with domain $\mathsf{P}$, we obtain $\operatorname{cl}(\phi_{q})=\phi$, defined by (3.2) in terms of $\bm{y}=\operatorname{cl}(\bm{\gamma})$.
Proposition 4.2.1.
  1. (1)
    There is a unique element $\mathbf{Y}=\sum_{\mu\in{\bm{\Phi}}^{+}}\mathbf{Y}_{\mu}\in\prod_{\mu\in{\bm{\Phi}}^{+}}{\mathfrak{n}}^{+}_{\mu}$ such that $\operatorname{cl}(\Upsilon)=\exp(\mathbf{Y})$.
  2. (2)
    For any $b\in{\mathfrak{k}}^{\prime}$, $\exp(\operatorname{ad}(\mathbf{Y}))(b)=\phi^{-1}(b)$.
  3. (3)
    Let $\alpha\in{\bm{\Phi}}^{+}$ such that $\theta(\alpha)\neq-\alpha$. Then $\mathbf{Y}_{\alpha}=0$.
Proof.
(1) The coproduct identity (3.16) reads
$\Delta(\Upsilon)={R}_{\theta}^{-1}\cdot 1\otimes\Upsilon\cdot{R}^{\phi_{q}^{-1}}\cdot\Upsilon\otimes 1.$ (4.3)
Since the classical limit of ${R}$ and ${R}_{\theta}$ is trivial, it follows that $\operatorname{cl}(\Upsilon)\in\prod_{\mu\geqslant 0}U({\mathfrak{n}}^{+})_{\mu}$ is a grouplike element with $\operatorname{cl}(\Upsilon)_{0}=1$. Thus its logarithm is well-defined and primitive.
(2) The intertwining identity (3.15), viz. $\Upsilon\cdot b=\phi_{q}^{-1}(b)\cdot\Upsilon$ for all $b\in U_{q}({\mathfrak{k}^{\prime}})$, yields
$\exp(\mathbf{Y})\cdot b\cdot\exp(\mathbf{Y})^{-1}=\phi^{-1}(b)\qquad\text{for all }b\in{\mathfrak{k}}^{\prime}$ (4.4)
from which the desired equation readily follows.
(3) First consider an arbitrary element $\lambda\in\mathsf{Q}^{+}\backslash(\mathsf{Q}^{+})^{-\theta}$. By Theorem 3.7.1, $\Upsilon_{\lambda}=0$ and therefore $\operatorname{cl}(\Upsilon)_{\lambda}=0$. Since $\operatorname{cl}(\Upsilon)=\exp(\mathbf{Y})$, it follows that
$\sum_{k\geq 1}\frac{1}{k!}\sum_{\begin{subarray}{c}\beta_{1},\ldots,\beta_{k}\in{\bm{\Phi}}^{+}\\ \beta_{1}+\ldots+\beta_{k}=\lambda\end{subarray}}\mathbf{Y}_{\beta_{1}}\cdots\mathbf{Y}_{\beta_{k}}=0.$ (4.5)
In particular, if $\alpha\in{\bm{\Phi}}^{+}$ is a simple root such that $\theta(\alpha)\neq-\alpha$, then for $\lambda=\alpha$ the identity (4.5) reduces to $\mathbf{Y}_{\alpha}=0$, as required. We now proceed by induction on ${\mathsf{ht}}(\alpha)$. Let $\alpha\in{\bm{\Phi}}^{+}$ be a positive root such that $\theta(\alpha)\neq-\alpha$ and ${\mathsf{ht}}(\alpha)>1$. For any decomposition $\alpha=\beta_{1}+\cdots+\beta_{k}$ with $k>1$ and $\beta_{1},\dots,\beta_{k}\in{\bm{\Phi}}^{+}$ we have $\theta(\beta_{j})\neq-\beta_{j}$ for some $j\in\{1,\ldots,k\}$. Since ${\mathsf{ht}}(\beta_{j})<{\mathsf{ht}}(\alpha)$, by the induction hypothesis one has $\mathbf{Y}_{\beta_{j}}=0$. Thus, for $\lambda=\alpha$ the identity (4.5) reduces to
$\mathbf{Y}_{\alpha}=-\sum_{k\geq 2}\frac{1}{k!}\sum_{\begin{subarray}{c}\beta_{1},\ldots,\beta_{k}\in{\bm{\Phi}}^{+}\\ \beta_{1}+\ldots+\beta_{k}=\alpha\end{subarray}}\mathbf{Y}_{\beta_{1}}\cdots\mathbf{Y}_{\beta_{k}}=0$ (4.6)
and the result follows. ∎
The following statement is a generalization, to symmetric pairs of non-diagonal type, of the fact that the classical limit of the quasi R-matrix is $1\in U({\mathfrak{g}}\oplus{\mathfrak{g}})$.
Proposition 4.2.2.
$\mathbf{Y}=0$ if and only if $\phi$ is an involution.
Proof.
If $\mathbf{Y}=0$ then (4.4) implies that ${\mathfrak{k}}^{\prime}$ is contained in the fixed-point subalgebra of $\phi$. Re-running the proof of Lemma 3.3.1, we deduce that $\phi$ is an involution.
Conversely, suppose $\phi$ is an involution. Now Lemma 3.3.1 and (4.4) imply that $[\exp(\mathbf{Y}),b]=0$ for all $b\in{\mathfrak{k}}^{\prime}$. Since ${\mathfrak{k}}^{\prime}$ contains the elements $f_{i}$ for all $i\in X$ and $f_{i}+\phi(f_{i})$ for all $i\in I\backslash X$, this in turn implies
$\sum_{k\geq 1}\frac{1}{k!}[\mathbf{Y}^{k},f_{i}]=\begin{cases}0&\text{if }i\in X,\\ \displaystyle\sum_{k\geq 1}\frac{1}{k!}[\phi(f_{i}),\mathbf{Y}^{k}]&\text{if }i\in I\backslash X,\end{cases}$ (4.7)
which induces a recursive relation among the components of $\mathbf{Y}$ with respect to the principal grading:
$\mathbf{Y}=\sum_{h\geq 1}\mathbf{Y}(h),\qquad\mathbf{Y}(h)\in\bigoplus_{\begin{subarray}{c}\lambda\in{\bm{\Phi}}^{+}\\ {\mathsf{ht}}(\lambda)=h\end{subarray}}{\mathfrak{n}}^{+}_{\lambda}\subseteq{\mathfrak{n}}^{+}.$ (4.8)
To specify this relation, for $\bm{\ell}=(\ell_{1},\ldots,\ell_{k})\in{\mathbb{Z}}_{>0}^{k}$ we denote $\mathbf{Y}(\bm{\ell})=\mathbf{Y}(\ell_{1})\cdots\mathbf{Y}(\ell_{k})$, so that $\mathbf{Y}^{k}=\sum_{\bm{\ell}\in{\mathbb{Z}}_{>0}^{k}}\mathbf{Y}(\bm{\ell})$. For $m\in{\mathbb{Z}}$ consider the set of ordered partitions of $m$ in $k$ parts:
${\mathcal{P}}_{k}(m)=\{\bm{\ell}=(\ell_{1},\ldots,\ell_{k})\in{\mathbb{Z}}_{>0}^{k}\,|\,\ell_{1}+\ldots+\ell_{k}=m\}.$
Finally, for $i\in I\backslash X$, note that $\phi(f_{i})$ is homogeneous of degree $m_{i}={\mathsf{ht}}(w_{X}(\alpha_{\tau(i)}))\in{\mathbb{Z}}_{>0}$. With this notation, the component of (4.7) of degree $m-1$ ($m\in{\mathbb{Z}}_{>0}$) is given by
$\sum_{k\geq 1}\frac{1}{k!}\sum_{\bm{\ell}\in{\mathcal{P}}_{k}(m)}[\mathbf{Y}(\bm{\ell}),f_{i}]=\begin{cases}0&\text{if }i\in X,\\ \displaystyle\sum_{k\geq 1}\sum_{\bm{\ell}\in{\mathcal{P}}_{k}(m-m_{i}-1)}\frac{1}{k!}[\phi(f_{i}),\mathbf{Y}(\bm{\ell})]&\text{if }i\in I\backslash X.\end{cases}$ (4.9)
We now show by induction that $\mathbf{Y}(m)=0$ for all $m\in{\mathbb{Z}}_{>0}$, yielding $\mathbf{Y}=0$ as desired. The case $m=1$ follows directly from (4.9): the left-hand side equals $[\mathbf{Y}(1),f_{i}]$; the right-hand side equals 0 since ${\mathcal{P}}_{k}(-m_{i})=\emptyset$ for all $k$. Hence we obtain $[\mathbf{Y}(1),f_{i}]=0$ for all $i\in I$ so that $\mathbf{Y}(1)=0$ by [Kac90, Lemma 1.5].
Now let $m\in{\mathbb{Z}}_{\geq 2}$ and suppose that $\mathbf{Y}(m^{\prime})=0$ for all $m^{\prime}\in\{1,\ldots,m-1\}$. Since for all $k>1$ the elements of ${\mathcal{P}}_{k}(m)$ are tuples whose entries are less than $m$, the induction hypothesis implies that the left-hand side of (4.9) equals its $k=1$ term, namely $[\mathbf{Y}(m),f_{i}]$. Also, for all $k\geq 1$ the elements of ${\mathcal{P}}_{k}(m-m_{i}-1)$ are tuples whose entries are less than $m$. Hence the right-hand side of (4.9) vanishes by the induction hypothesis. We obtain that $[\mathbf{Y}(m),f_{i}]=0$ for all $i\in I$ and by [Kac90, Lemma 1.5] we deduce $\mathbf{Y}(m)=0$, which completes the proof. ∎
4.3. Proof of Theorem 4.1.2
Since $B^{\prime}_{\xi}$ is a right coideal subalgebra which contains $U_{q}({\mathfrak{k}^{\prime}})$, it is enough to prove that $\operatorname{cl}(B^{\prime}_{\xi})=U({\mathfrak{k}}^{\prime})$. Indeed, in this case Kolb’s maximality theorem, see Theorem 4.1.1, implies $B^{\prime}_{\xi}=U_{q}({\mathfrak{k}^{\prime}})$. We start with the following basic result.
Lemma 4.3.1.
Let $H$ be a Hopf algebra over a field $k$ generated by primitive elements. Then every coideal subalgebra $H^{\prime}$ of $H$ is a Hopf subalgebra.
Proof.
Since $H$ is cocommutative, $H^{\prime}$ is a subbialgebra. It remains to prove that every subbialgebra $H^{\prime}\subseteq H$ is preserved by the antipode $S$, which is a consequence from the following argument (see e.g. [EE05, Proof of Prop. 4.2] and [FGB05, Prop. 5.2]). Let $\eta$, $\varepsilon$, $m^{(n)}$ and $\Delta^{(n)}$ denote the unit, counit, $n$-fold product and $n$-fold coproduct, respectively. With respect to the convolution product on $\operatorname{End}_{k}(H)$, the antipode is the inverse of ${\operatorname{id}}_{H}$, the neutral element being $\eta\circ\varepsilon$. Since any product of primitive elements is annihilated by $(\eta\circ\epsilon-{\operatorname{id}}_{H})^{\otimes n}\circ\Delta^{(n)}\in\operatorname{Hom}_{k}(H,H^{\otimes n})$ for sufficiently large $n$, the following identity makes sense:
$S=\sum_{n\geq 0}m^{(n)}\circ(\eta\circ\varepsilon-{\operatorname{id}}_{H})^{\otimes n}\circ\Delta^{(n)}.$
Using this, we can express $S(h)$ for all $h\in H^{\prime}$ in terms of the bialgebra structure; the desired result follows. ∎
By [MM65, Thm. 5.18, Prop. 6.13] we conclude that the coideal subalgebra $\operatorname{cl}(B^{\prime}_{\xi})$ of $U({\mathfrak{g}}^{\prime})$ is equal to $U({\mathfrak{f}})$ for some Lie subalgebra ${\mathfrak{f}}\subseteq{\mathfrak{g}}^{\prime}$. Moreover, since $\operatorname{cl}(U_{q}({\mathfrak{k}^{\prime}}))=U({\mathfrak{k}}^{\prime})$ and $\operatorname{cl}((U_{q}({\mathfrak{g}}^{\prime}))^{\xi})=U({\mathfrak{g}}^{\prime})^{\operatorname{cl}(\xi)}$, we have
${\mathfrak{k}}^{\prime}\subseteq{\mathfrak{f}}\subseteq({\mathfrak{g}}^{\prime})^{\operatorname{cl}(\xi)}.$ (4.10)
Note that $\operatorname{cl}(\xi)=\operatorname{Ad}(\operatorname{cl}(\Upsilon))\circ\phi$. By Proposition 4.2.1, $\operatorname{cl}(\Upsilon)=\exp(\mathbf{Y})$ for some $\mathbf{Y}=\sum_{\mu\in{\bm{\Phi}}^{+}}\mathbf{Y}_{\mu}\in\prod_{\mu\in{\bm{\Phi}}^{+}}{\mathfrak{n}}^{+}_{\mu}$. Thus,
$({\mathfrak{g}}^{\prime})^{\operatorname{cl}(\xi)}=\{x\in{\mathfrak{g}}^{\prime}\,|\,\operatorname{Ad}(\operatorname{cl}(\Upsilon))(x)=\phi^{-1}(x)\}=\{x\in{\mathfrak{g}}^{\prime}\,|\,\exp(\operatorname{ad}(\mathbf{Y}))(x)=\phi^{-1}(x)\}.$ (4.11)
Lemma 4.3.2.
${\mathfrak{k}}={\mathfrak{g}}^{\operatorname{cl}(\xi)}$.
Proof.
Clearly, ${\mathfrak{k}}\subseteq{\mathfrak{g}}^{\operatorname{cl}(\xi)}$. Denote ${\mathfrak{n}}^{+}_{\theta}={\mathfrak{n}}^{+}\cap\theta({\mathfrak{n}}^{-})={\mathfrak{n}}^{+}\cap\phi({\mathfrak{n}}^{-})$. By [RV22, Cor. 3.10], one has the Iwasawa decomposition:
${\mathfrak{g}}={\mathfrak{k}}\oplus{\mathfrak{h}}^{-\theta}\oplus{\mathfrak{n}}^{+}_{\theta}$ (4.12)
(as linear spaces). Now suppose $x\in{\mathfrak{h}}^{-\theta}\oplus{\mathfrak{n}}^{+}_{\theta}$ is fixed by $\operatorname{cl}(\xi)$, i.e.,
$\exp(\operatorname{ad}(\mathbf{Y}))(x)=\phi^{-1}(x).$ (4.13)
Denote by ${\mathfrak{b}}^{\pm}={\mathfrak{h}}\oplus{\mathfrak{n}}^{\pm}$ the standard Borel subalgebras. Note that $\exp(\operatorname{ad}(\mathbf{Y}))(x)\in\prod_{\mu\in\mathsf{Q}^{+}}{\mathfrak{b}}^{+}_{\mu}$ whereas $\phi^{-1}(x)\in{\mathfrak{b}}^{-}$. Hence (4.13) implies $x\in{\mathfrak{h}}$, so that $x\in{\mathfrak{h}}^{-\theta}={\mathfrak{h}}^{-\phi}$. Rewriting (4.13) as $\exp(\mathbf{Y})x=-x\exp(\mathbf{Y})$ and comparing weight-0 components, we obtain $x=0$. Now $({\mathfrak{h}}^{-\theta}\oplus{\mathfrak{n}}^{+}_{\theta})^{\operatorname{cl}(\xi)}=\{0\}$ implies the desired result. ∎
Intersecting with ${\mathfrak{g}}^{\prime}$, we obtain ${\mathfrak{k}}^{\prime}=({\mathfrak{g}}^{\prime})^{\operatorname{cl}(\xi)}$ so that ${\mathfrak{f}}={\mathfrak{k}}^{\prime}$. Theorem 4.1.1 now implies Theorem 4.1.2.
Remark 4.3.3.
If $\phi$ is an involution, Lemma 4.3.2 follows immediately from Proposition 4.2.2. $\triangledown$
5. Tensor K-matrices
5.1. The QSP weight lattice $\mathsf{P}_{\theta}$
The involution $\theta:{\mathfrak{h}}\to{\mathfrak{h}}$ determines several lattices in ${\mathfrak{h}}$ and ${\mathfrak{h}}^{*}$ equipped with a natural pairing, see e.g.,  [BW18a, Wa24]. Adopting a slightly different approach, we consider the quotient lattice
$\mathsf{P}_{\theta}=\mathsf{P}/\mathsf{P}^{-\theta},$ (5.1)
with canonical projection $[\cdot]_{\theta}:\mathsf{P}\twoheadrightarrow\mathsf{P}_{\theta}$. The fixed-point sublattice of the ${\mathbb{Z}}$-linear automorphism $-\theta$ satisfies
$\mathsf{P}^{-\theta}=\mathsf{P}\cap\operatorname{span}_{\tfrac{1}{2}{\mathbb{Z}}}\{\lambda-\theta(\lambda)\,|\,\lambda\in\mathsf{P}\}.$ (5.2)
Given the bilinear pairing $\mathsf{P}\times\mathsf{Q}^{\vee}_{\sf ext}\to{\mathbb{Z}}$, define a pairing $\mathsf{P}_{\theta}\times\mathsf{Q}^{\vee}_{\sf ext}\to\tfrac{1}{2}{\mathbb{Z}}$ as follows: $\zeta(h)=\tfrac{1}{2}\lambda(h+\theta(h))$ for $\zeta=[\lambda]_{\theta}\in\mathsf{P}_{\theta}$ with $\lambda\in\mathsf{P}$ and $h\in\mathsf{Q}^{\vee}_{\sf ext}$ (note that it is independent of the choice of $\lambda$). This restricts to a nondegenerate pairing $\mathsf{P}_{\theta}\times(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}\to{\mathbb{Z}}$.
Remark 5.1.1.
Set $I_{\sf ns}=\{i\in I\,|\,\theta(\alpha_{i})=-\alpha_{i}\}$ so that $[\alpha_{i}]_{\theta}=0$ if and only if $i\in I_{\sf ns}$. Note that $\mathsf{P}_{\theta}$ identifies with a subgroup of the lattice $X^{\imath}$ considered in [Wa24, Sec. 3.2] of index $2^{|I_{\sf ns}|}$. $\triangledown$
5.2. QSP weight modules
Recall that we have
$U_{q}({\mathfrak{h}}^{\theta})=\langle K_{h}\;|\;h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}\rangle=U_{q}(\mathfrak{k})\cap U_{q}({\mathfrak{h}}).$
Definition 5.2.1.
A $U_{q}(\mathfrak{k})$-module $M$ is a (type $\bf 1$) QSP weight module if, as a $U_{q}({\mathfrak{h}}^{\theta})$-module,
$M=\bigoplus_{\zeta\in\mathsf{P}_{\theta}}M_{\zeta}$ (5.3)
where the QSP weight spaces are defined by
$M_{\zeta}=\{m\in M\,|\,\forall h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta},\,K_{h}\cdot m=q^{\zeta(h)}m\}.$ (5.4)
We denote the full subcategory of QSP weight modules by $\mathcal{W}_{\theta}\subseteq{\operatorname{Mod}}(U_{q}(\mathfrak{k}))$.
Remark 5.2.2.
Consider the case of a split quantum symmetric pair, i.e., $\theta=\omega$. Then $\mathsf{P}_{\theta}=\{0\}$ and $\mathcal{W}_{\theta}={\operatorname{Mod}}(U_{q}(\mathfrak{k}))$. In particular, every $U_{q}(\mathfrak{k})$-module $M$ is a QSP weight module which consists of a single QSP weight space: $M=M_{0}$. $\triangledown$
In the following proposition we collect some elementary properties of $\mathcal{W}_{\theta}$.
Proposition 5.2.3.
  1. (1)
    Let $M\in\mathcal{W}_{\theta}$. For any $\zeta\in\mathsf{P}_{\theta}$, $i\in I$ and $j\in X$, one has
    $B_{i}\cdot M_{\zeta}\subseteq M_{\zeta-[\alpha_{i}]_{\theta}}\qquad\mbox{and}\qquad E_{j}\cdot M_{\zeta}\subseteq M_{\zeta+[\alpha_{j}]_{\theta}}.$
  2. (2)
    All finite-dimensional irreducible $U_{q}(\mathfrak{k})$-modules are objects in $\mathcal{W}_{\theta}$ up to twisting by an automorphism of $U_{q}(\mathfrak{k})$.
  3. (3)
    $\mathcal{W}_{\theta}$ is a right module category over the monoidal category $\mathcal{W}$. Further, every weight $U_{q}({\mathfrak{g}})$-module is a QSP weight module under restriction to $U_{q}(\mathfrak{k})$.
  4. (4)
    The action of $U_{q}(\mathfrak{k})$ on $\mathcal{W}_{\theta}$ is faithful, i.e., $U_{q}(\mathfrak{k})$ embeds as a subalgebra into $\operatorname{End}({\mathsf{f}}_{\theta})$, where ${\mathsf{f}}_{\theta}:\mathcal{W}_{\theta}\to\operatorname{Vect}$ is the forgetful functor.
Proof.
(1) Note that $\bm{\sigma}_{i}=0$ if $\theta(\alpha_{i})\neq-\alpha_{i}$. Therefore, the result follows from the relations
$K_{h}E_{j}=q_{j}^{\alpha_{j}(h)}E_{j}K_{h}\quad K_{h}B_{i}=q_{i}^{-\alpha_{i}(h)}B_{i}K_{h}$ (5.5)
for any $i\in I$, $j\in X$, $h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}$.
(2) This is a consequence of Theorem A.4.1 which we prove in the self-contained appendix A.
(3) Since $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ is a (right) coideal subalgebra, ${\operatorname{Mod}}(U_{q}(\mathfrak{k}))$ is a right module category over ${\operatorname{Mod}}(U_{q}({\mathfrak{g}}))$ with monoidal action given by the tensor product, cf. C.1. In particular, for any $M\in\mathcal{W}_{\theta}$ and $V\in\mathcal{W}$, one has $M\otimes V\in{\operatorname{Mod}}(U_{q}(\mathfrak{k}))$. Then,
$M\otimes V=\bigoplus_{\zeta\in\mathsf{P}_{\theta}}\underbrace{\left(\bigoplus_{\mu\in\mathsf{P}}M_{\zeta-[\mu]_{\theta}}\otimes V_{\mu}\right)}_{(M\otimes V)_{\zeta}}$
Let $h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}$ and $\zeta\in\mathsf{P}_{\theta}$. Then, $K_{h}$ acts on $(M\otimes V)_{\zeta}$ as multiplication by $q^{\zeta(h)}$. Namely, write $\zeta=[\lambda]_{\theta}$ for some $\lambda\in\mathsf{P}$. Then, it is enough to observe that $K_{h}$ is grouplike and it acts on $M_{\zeta-[\mu]_{\theta}}$ and $V_{\mu}$ via multiplication by $q^{(\lambda-\mu)(h)}$ and $q^{\mu(h)}$, respectively.
Finally, note that the restriction of $V$ to $U_{q}(\mathfrak{k})$ is given by the action of $V$ on the trivial representation $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})\stackrel{{\scriptstyle\varepsilon}}{{\longrightarrow}}{\mathbb{F}}$, thus yielding a restriction functor $\mathcal{W}\to\mathcal{W}_{\theta}$.
(4) By (3), every module in ${\mathcal{O}}$ restricts to $\mathcal{W}_{\theta}$. This yields a morphism $\operatorname{End}({\mathsf{f}}_{\theta})\to\operatorname{End}({\mathsf{f}})$ and a commutative diagram
${U_{q}(\mathfrak{k})}$${U_{q}({\mathfrak{g}})}$${\operatorname{End}({\mathsf{f}}_{\theta})}$${\operatorname{End}({\mathsf{f}})}$ (5.6)
where the right vertical arrow is injective by [ATL24]. It follows that the map $U_{q}(\mathfrak{k})\to\operatorname{End}({\mathsf{f}}_{\theta})$ is also injective. ∎
5.3. Tensor weight operators
Let $f\colon\mathsf{P}\to{\mathbb{F}}$ be a ${\mathbb{Z}}$-linear functional which is self-adjoint with respect to the pairing on $\mathsf{P}$, i.e., $(f(\lambda),\mu)=(\lambda,f(\mu))$ for any $\lambda,\mu\in\mathsf{P}$. We consider the symmetric functional $\operatorname{C}^{f}:\mathsf{P}\times\mathsf{P}\to{\mathbb{F}}$ given by
$\operatorname{C}^{f}(\lambda,\mu)=q^{(f(\lambda),\mu)}.$ (5.7)
For any $V,W\in\mathcal{W}$, we denote by the same symbol the diagonal operator on $V\otimes W$, acting on the subspace $V_{\lambda}\otimes W_{\mu}$ as multiplication by $\operatorname{C}^{f}(\lambda,\mu)$. Let ${\mathsf{f}}_{\mathcal{W}}^{2}\colon\mathcal{W}\boxtimes\mathcal{W}\to\operatorname{Vect}$ be the 2-fold forgetful functor. Then, one readily checks that $\operatorname{C}^{f}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{2})$ and in $\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{3})$ the following identities hold:
$(\Delta\otimes{\operatorname{id}})(\operatorname{C}^{f})=\operatorname{C}^{f}_{13}\operatorname{C}^{f}_{23},\qquad({\operatorname{id}}\otimes\Delta))(\operatorname{C}^{f})=\operatorname{C}^{f}_{13}\operatorname{C}^{f}_{12}.$ (5.8)
5.4. Tensor QSP weight operators
Let $f\colon\mathsf{P}\to{\mathbb{F}}$ be a ${\mathbb{Z}}$-linear, self-adjoint functional such that $f\circ\theta=f$. Then, $f(\mathsf{P}^{-\theta})=0$ and $\operatorname{C}^{f}$ descends to a function $\mathsf{P}_{\theta}\times\mathsf{P}\to{\mathbb{F}}$. Proceeding as in Section 5.3, $\operatorname{C}^{f}$ determines a diagonal operator on any tensor product $M\otimes V$ with $M\in\mathcal{W}_{\theta}$ and $V\in\mathcal{W}$, which is natural in both $M$ and $V$. Let
${\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{\mathcal{W}}\colon\mathcal{W}_{\theta}\boxtimes\mathcal{W}\to\operatorname{Vect}$
be the 2-fold forgetful functor ${\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{\mathcal{W}}(M,V)=M\otimes V$. Then, $\operatorname{C}^{f}\in\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{\mathcal{W}})$.
5.5. Tensor Drinfeld algebras
For any $\beta\in\mathsf{Q}^{+}$, let ${\mathcal{B}}_{\beta}$ be a basis of $U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}$ and set ${\mathcal{B}}=\bigsqcup_{\beta\in\mathsf{Q}^{+}}{\mathcal{B}}_{\beta}$. Consider the Drinfeld algebra
${\mathcal{D}}=\left\{\left.\sum_{x\in{\mathcal{B}}}c_{x}x\,\right|\,c_{x}\in U_{q}{({\mathfrak{b}}^{-})}\right\}=\prod_{\beta\in\mathsf{Q}^{+}}U_{q}{({\mathfrak{b}}^{-})}U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\supseteq U_{q}({\mathfrak{g}}).$ (5.9)
One readily checks that ${\mathcal{D}}$ has a unique algebra structure which extends that of $U_{q}({\mathfrak{g}})$. Moreover, every element in ${\mathcal{D}}$ naturally acts on category ${\mathcal{O}}$ modules, since it reduces to a finite sum due to the condition (C3). By [ATL24, Sec. 3.1], ${\mathcal{D}}$ embeds as a subalgebra in both $\operatorname{End}({\mathsf{f}})$ and $\operatorname{End}({\mathsf{f}}_{{\operatorname{int}}})$, where ${\mathsf{f}}:{\mathcal{O}}\to\operatorname{Vect}$ and ${\mathsf{f}}_{{\operatorname{int}}}\colon{\mathcal{O}}_{{\operatorname{int}}}\to\operatorname{Vect}$ are the forgetful functors.
Remark 5.5.1.
By construction, the quasi K-matrix $\Upsilon$ is an element of ${\mathcal{D}}$, see Theorem 3.7.1. $\triangledown$
In a similar way, we equip
${\mathcal{E}}_{\theta}$ =$=\left\{\left.\sum_{x\in{\mathcal{B}}}b_{x}\otimes c_{x}x\,\right|\,b_{x}\in U_{q}(\mathfrak{k})\,,\,c_{x}\in U_{q}{({\mathfrak{b}}^{-})}\right\}$ (5.10)
$=\prod_{\beta\in\mathsf{Q}^{+}}U_{q}(\mathfrak{k})\otimes U_{q}{({\mathfrak{b}}^{-})}U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\supseteq U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{g}})$ (5.11)
with the unique algebra structure extending that of $U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{g}})$. Every element in ${\mathcal{E}}_{\theta}$ acts in particular on tensor products of the form $M\otimes V$ with $M\in\mathcal{W}_{\theta}$ and $V\in{\mathcal{O}}$. The action is natural in $M$ and $V$, thus yielding an morphism of algebras ${\mathcal{E}}_{\theta}\to\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}})$. Proceeding as in [ATL24, Thm. 3.1], one further shows that the action of ${\mathcal{E}}_{\theta}$ is faithful on category ${\mathcal{O}}$ integrable $U_{q}({\mathfrak{g}})$-modules. Thus, we obtain a commutative diagram
${{\mathcal{E}}_{\theta}}$${\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})}$${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}})}$${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\operatorname{int}}})}$ (5.12)
where the first and second horizontal arrows are induced by the restriction functor ${\mathcal{O}}\to\mathcal{W}_{\theta}$ and the inclusion ${\mathcal{O}}_{{\operatorname{int}}}\hookrightarrow{\mathcal{O}}$, respectively.
Finally, set ${\mathcal{B}}_{X}=\bigsqcup_{\beta\in\mathsf{Q}_{X}^{+}}{\mathcal{B}}_{\beta}$. We consider the opposite algebra
${\mathcal{E}}^{\operatorname{op}}_{X}$ =$=\left\{\left.\sum_{x\in{\mathcal{B}}_{X}}a_{x}\otimes c_{x}\omega(x)\,\right|\,a_{x}\in U_{q}({\mathfrak{g}}_{X})\,,\,c_{x}\in U_{q}({\mathfrak{b}}_{X}^{+})\right\}$ (5.13)
$=\prod_{\beta\in\mathsf{Q}_{X}^{+}}U_{q}({\mathfrak{g}}_{X})\otimes U_{q}({\mathfrak{b}}_{X}^{+})U_{q}({\mathfrak{n}}_{X}^{-})_{-\beta}\supseteq U_{q}({\mathfrak{g}}_{X})\otimes U_{q}({\mathfrak{g}}_{X})$ (5.14)
We observe that ${\mathcal{E}}^{\operatorname{op}}_{X}$ acts on any tensor product $V\otimes W$ with $V\in{\operatorname{Mod}}(U_{q}({\mathfrak{g}}_{X}))$ and $W\in{\mathcal{O}}_{{\operatorname{int}}}(U_{q}({\mathfrak{g}}_{X}))$. Namely, by [Lus94, Thm. 6.2.2], a category ${\mathcal{O}}$ integrable module is completely reducible. Since $X\subset I$ is a subdiagram of finite type, $W$ decomposes as a (possibly infinite) sum of finite-dimensional irreducible $U_{q}({\mathfrak{g}}_{X})$-modules. Thus, the action of ${\mathcal{E}}^{\operatorname{op}}_{X}$ on $V\otimes W$ is well-defined. Through the restriction functor ${\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}\to{\mathcal{O}}_{{\operatorname{int}}}(U_{q}({\mathfrak{g}}_{X}))$ and the inclusion $U_{q}({\mathfrak{g}}_{X})\subset U_{q}(\mathfrak{k})$, we obtain an embedding
${{\mathcal{E}}^{\operatorname{op}}_{X}}$${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}}).}$ (5.15)
5.6. The tensor K-matrix
Let
$\psi=\operatorname{Ad}(\bm{\gamma})\circ\theta_{q}^{-1}=\phi_{q}^{-1},\quad J=R_{\theta},\quad\mbox{and}\quad K=\Upsilon,$ (5.16)
be the cylindrical structure on $U_{q}({\mathfrak{g}})$ with respect to the QSP subalgebra $U_{q}({\mathfrak{k}^{\prime}})$ constructed in Theorem 3.7.1. Let $R_{21}^{\psi}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ be the operator defined in Proposition 3.6.3 (2).
Theorem 5.6.1.
The operator
${{\mathbb{K}}}=R_{21}^{\psi}\cdot 1\otimes K\cdot R\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ (5.17)
has the following properties.
  1. (1)
    It satisfies the following QSP intertwining identity in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$:
    ${{\mathbb{K}}}\cdot\Delta(b)=({\operatorname{id}}\otimes\psi)(\Delta(b))\cdot{{\mathbb{K}}}\qquad\text{for all }b\in U_{q}(\mathfrak{k}).$ (5.18)
  2. (2)
    It satisfies the following coproduct identities in $\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}\boxtimes{\mathsf{f}})$:
    $(\Delta\otimes{\operatorname{id}})({{\mathbb{K}}})$ $=R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23},$ (5.19)
    $({\operatorname{id}}\otimes\Delta)({{\mathbb{K}}})$ $=J^{-1}_{23}\cdot{{\mathbb{K}}}_{13}\cdot R^{\psi}_{23}\cdot{{\mathbb{K}}}_{12}$ (5.20)
  3. (3)
    It satisfies the generalized reflection equation:
    $R^{\psi\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R^{\psi}_{23}\cdot{{\mathbb{K}}}_{12}={{\mathbb{K}}}_{12}\cdot R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}\boxtimes{\mathsf{f}}).$ (5.21)
  4. (4)
    There are elements $\Xi_{X}\in{\mathcal{E}}^{\operatorname{op}}_{X}$ and $\Xi\in{\mathcal{E}}_{\theta}$ such that
    ${{\mathbb{K}}}=\operatorname{C}^{{\operatorname{id}}+\theta}\cdot\Xi_{X}\cdot\Xi.$
    In particular, ${{\mathbb{K}}}$ admits a canonical lift171717 By construction, ${{\mathbb{K}}}$ acts on any tensor product $V\otimes W$ with $V,W\in{\mathcal{O}}$. This means that the action of ${{\mathbb{K}}}$ is extended to any tensor product $M\otimes W$ with $M\in\mathcal{W}_{\theta}$ but $W\in{\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}$. in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$, the identity (5.18) remains valid in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$, and181818 Since $\mathcal{W}_{\theta}$ is a module category over ${\mathcal{O}}$, the coproduct $\Delta$ naturally lifts to a morphism $\operatorname{End}({\mathsf{f}}_{\theta})\to\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}})$, cf. Section 2.6. the identities (5.19-5.21) remain valid in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}}\boxtimes{\mathsf{f}}_{{{X}\negthinspace\operatorname{-int}}})$.
The proof is carried out in Sections 5.7 and 5.8.
Remark 5.6.2.
One recovers $K$ from ${{\mathbb{K}}}$ via $K=(\varepsilon\otimes{\operatorname{id}})({{\mathbb{K}}})$. $\triangledown$
5.7. Proof of Theorem 5.6.1 (1), (2), and (3)
(1) For any $b\in U_{q}(\mathfrak{k})$ we have
$R^{\psi}_{21}\cdot 1\otimes\Upsilon\cdot R\cdot\Delta(b)$ $=R^{\psi}_{21}\cdot({\operatorname{id}}\otimes\psi)(\Delta^{\operatorname{op}}(b))\cdot 1\otimes\Upsilon\cdot R$ (5.22)
$=({\operatorname{id}}\otimes\psi)(R_{21}\cdot\Delta^{\operatorname{op}}(b))\cdot 1\otimes\Upsilon\cdot R$ (5.23)
$=({\operatorname{id}}\otimes\psi)(\Delta(b))\cdot R^{\psi}_{21}\cdot 1\otimes\Upsilon\cdot R,$ (5.24)
where the equality follows from the intertwining identity of $\Upsilon$ (3.15), the coideal property of $U_{q}(\mathfrak{k})$, and the intertwining property of $R$ (2.10).
(2) The first coproduct identity (5.19) follows from that of the R-matrix (2.11):
$(\Delta\otimes{\operatorname{id}})(R^{\psi}_{21}\cdot 1\otimes\Upsilon\cdot R)$ $=({\operatorname{id}}\otimes{\operatorname{id}}\otimes\psi)(R_{32}\cdot R_{31})\cdot(1\otimes 1\otimes\Upsilon)\cdot R_{13}\cdot R_{23}$ (5.25)
$=R^{\psi}_{32}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes\Upsilon)\cdot R_{13}\cdot R_{23}.$ (5.26)
On the other hand, the second coproduct identity (5.20) follows from that of $\Upsilon$ (4.3) and Lemma 3.6.1 (3). Namely, since $\Delta_{J}=\Delta^{\operatorname{op},\psi}$, one has
$({\operatorname{id}}\otimes\Delta)(R^{\psi}_{21})\cdot J^{-1}_{23}$ $=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\Delta_{J})(R^{\psi}_{21})$ (5.27)
$=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\Delta^{\operatorname{op},\psi})(R^{\psi}_{21})$ (5.28)
$=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\psi\otimes\psi)({\operatorname{id}}\otimes\Delta^{\operatorname{op}})(R_{21})$ (5.29)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot R^{\psi}_{21}.$ (5.30)
Therefore,
$({\operatorname{id}}\otimes\Delta)$ $(R^{\psi}_{21}\cdot 1\otimes\Upsilon\cdot R)=$ (5.31)
$=({\operatorname{id}}\otimes\Delta)(R^{\psi}_{21})\cdot J^{-1}_{23}\cdot(1\otimes 1\otimes\Upsilon)\cdot R^{\psi}_{23}\cdot(1\otimes\Upsilon\otimes 1)\cdot R_{13}\cdot R_{12}$ (5.32)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot R^{\psi}_{21}\cdot(1\otimes 1\otimes\Upsilon)\cdot R^{\psi}_{23}\cdot(1\otimes\Upsilon\otimes 1)\cdot R_{13}\cdot R_{12}$ (5.33)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes\Upsilon)\cdot R^{\psi}_{21}\cdot R^{\psi}_{23}\cdot R_{13}\cdot(1\otimes\Upsilon\otimes 1)\cdot R_{12}$ (5.34)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes\Upsilon)\cdot R_{13}\cdot R^{\psi}_{23}\cdot R^{\psi}_{21}\cdot(1\otimes\Upsilon\otimes 1)\cdot R_{12}$ (5.35)
where in the last equality follows from the Yang-Baxter equation. Finally, the generalized reflection equation (5.21) follows from (5.18), and (5.19).
5.8. Proof of Theorem 5.6.1 (4)
(3) We observe that, since $\theta|_{\mathfrak{h}}$ is involutive, the weight operator $\operatorname{C}^{{\operatorname{id}}+\theta}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}\boxtimes{\mathsf{f}}_{\mathcal{W}})$ has a canonical lift in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{\mathcal{W}})$, see Section 5.4. We have to prove that
$\operatorname{C}^{-{\operatorname{id}}-\theta}\cdot\,{{\mathbb{K}}}=\Xi_{X}\cdot\Xi$
for some $\Xi_{X}\in{\mathcal{E}}^{\operatorname{op}}_{X}$ and $\Xi\in{\mathcal{E}}_{\theta}$.
The universal R-matrix is equipped with the factorization $R=\operatorname{C}^{{\operatorname{id}}}\cdot\,\Theta$, where
$\Theta=\sum_{\beta\in\mathsf{Q}^{+}}\Theta_{\beta}\in\prod_{\beta\in\mathsf{Q}^{+}}U_{q}{({{\mathfrak{n}}}^{-})}_{-\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\,$
see Section 2.6. By Lemma 3.6.1 (1), one readily checks that
$\operatorname{C}^{{\operatorname{id}}}_{V{W^{\psi}}}=\operatorname{C}^{\theta}_{VW}$
for any $V,W\in\mathcal{W}$. Hence, $R_{21}^{\psi}=\operatorname{C}^{\theta}\cdot\,\Theta_{21}^{\psi}$. Thus, we have
$\operatorname{C}^{-{\operatorname{id}}-\theta}\cdot\,{{\mathbb{K}}}=\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(\Theta^{\psi}_{21})\cdot\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(1\otimes\Upsilon)\cdot\Theta.$ (5.36)
Let ${R}_{X}$ be the universal R-matrix of $U_{q}({\mathfrak{g}}_{X})$ with the corresponding factorization ${R}_{X}=\operatorname{C}^{{\operatorname{id}}}\cdot\Theta_{X}$. By [AV22a, Prop. 4.3],
$\widetilde{\Theta}=\Theta_{X}^{-1}\cdot\Theta\in\prod_{\beta\in\mathsf{Q}^{+}\setminus\mathsf{Q}_{X}^{+}}U_{q}{({{\mathfrak{n}}}^{-})}_{-\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}.$ (5.37)
Since $\psi$ is the identity on $U_{q}({\mathfrak{g}}_{X})$, we have $\Theta_{21}^{\psi}=\Theta_{X,21}\cdot\widetilde{\Theta}_{21}^{\psi}$. Set
$\Xi_{X}=\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(\Theta_{X,21})\quad\mbox{and}\quad\Xi=\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(\widetilde{\Theta}^{\psi}_{21})\cdot\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(1\otimes\Upsilon)\cdot\Theta.$ (5.38)
Clearly, $\Xi_{X}\in{\mathcal{E}}^{\operatorname{op}}_{X}$. It remains to prove that $\Xi\in{\mathcal{E}}_{\theta}$. Set
${\mathcal{D}}^{(2)}=\left\{\left.\sum_{x\in{\mathcal{B}}}a_{x}\otimes c_{x}x\,\right|\,a_{x}\in U_{q}({\mathfrak{g}})\,,\,c_{x}\in U_{q}{({\mathfrak{b}}^{-})}\right\}\supset{\mathcal{E}}_{\theta}.$
We claim that $\Xi\in{\mathcal{D}}^{(2)}$. Clearly, $\operatorname{Ad}(\operatorname{C}^{-{\operatorname{id}}})(1\otimes\Upsilon)\cdot\Theta\in{\mathcal{D}}^{(2)}$. Then, it is enough to observe that $-\theta$ preserves $\mathsf{Q}^{+}\setminus\mathsf{Q}_{X}^{+}$. Therefore,
$\widetilde{\Theta}_{21}^{\psi}\in\prod_{\beta\in\mathsf{Q}^{+}\setminus\mathsf{Q}_{X}^{+}}U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\otimes U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}\subset{\mathcal{D}}^{(2)}.$
Finally, set $\xi=\operatorname{Ad}(\Upsilon)^{-1}\circ\psi$ as in (4.1). By Theorem 4.1.2,
$U_{q}({\mathfrak{k}^{\prime}})=\{u\in U_{q}({\mathfrak{g}}^{\prime})\,|\,(\xi\otimes{\operatorname{id}})(\Delta(u))=\Delta(u)\}.$
Therefore, $\Xi\in{\mathcal{E}}_{\theta}$ if and only if
$(\xi\otimes{\operatorname{id}}\otimes{\operatorname{id}})\circ(\Delta\otimes{\operatorname{id}})(\Xi)=(\Delta\otimes{\operatorname{id}})(\Xi).$ (5.39)
Note that (5.39) is automatically satisfied by $\Xi_{X}$, since $\xi$ is the identity on $U_{q}(\mathfrak{k})$. Moreover, it is also satisfied by $\operatorname{C}^{{\operatorname{id}}+\theta}$, since
  • the coproduct of $\operatorname{C}^{{\operatorname{id}}+\theta}$ is provided by (5.8);
  • for any $V,W\in\mathcal{W}$, one has $\operatorname{C}^{{\operatorname{id}}+\theta}_{{V^{\psi}}W}=\operatorname{C}^{{\operatorname{id}}+\theta}_{VW}$;
  • $\operatorname{C}^{{\operatorname{id}}+\theta}$ commutes with $\Upsilon$ because $\Upsilon_{\mu}=0$ whenever $\mu\not\in(\mathsf{Q}^{+})^{-\theta}$.
Thus, $\Xi\in{\mathcal{E}}_{\theta}$ if and only if ${{\mathbb{K}}}$ satisfies (5.39). Finally, it is enough to observe that, by (5.19), the identity (5.39) for ${{\mathbb{K}}}$ follows from the generalized reflection equation (3.17) for $\Upsilon$. Namely,
$(\xi\otimes{\operatorname{id}}\otimes{\operatorname{id}})$ $\circ(\Delta\otimes{\operatorname{id}})({{\mathbb{K}}})=$ (5.40)
=(5.19)${{=}}(\xi\otimes{\operatorname{id}}\otimes{\operatorname{id}})(R^{\psi}_{32}\cdot R^{\psi}_{31}\cdot 1\otimes 1\otimes K\cdot R_{13}\cdot R_{23})$ ) (5.41)
=(4.1)${{=}}R^{\psi}_{32}\cdot K^{-1}\otimes 1\otimes 1\cdot R^{\psi\psi}_{31}\cdot 1\otimes 1\otimes K\cdot R^{\psi}_{13}\cdot K\otimes 1\otimes 1\cdot R_{23}$ (5.42)
=(3.17)${{=}}R^{\psi}_{32}\cdot R^{\psi}_{31}\cdot 1\otimes 1\otimes K\cdot R_{13}\cdot R_{23}$ (5.43)
=(5.19)${{=}}(\Delta\otimes{\operatorname{id}})({{\mathbb{K}}}).$ ) . (5.44)
5.9. Boundary bimodule structure
In the Appendix, we describe a categorical framework which encodes the defining properties of the tensor K-matrix ${{\mathbb{K}}}$. In Section C.4, we introduce the notion of a boundary bimodule category, which is a generalization of that of braided module category, see e.g., [Enr07, Bro13, Kol20].
Corollary 5.9.1.
Set
$\psi=\operatorname{Ad}(\bm{\gamma})\circ\theta_{q}^{-1}=\phi_{q}^{-1},\quad J=R_{\theta},\quad\mbox{and}\quad K=\Upsilon,$ (5.45)
and let $R_{21}^{\psi}\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ be as in Proposition 3.6.3 (2).
  1. (1)
    The category $\mathcal{W}_{\theta}$ has the following natural bimodule structure over ${\mathcal{O}}$, cf. C.3.
    • The right action is given by $M\triangleleft V=M\otimes V$ for any $M\in\mathcal{W}_{\theta}$ and $V\in{\mathcal{O}}$.
    • The left action is given by $M\blacktriangleleft V=M\otimes V^{\psi}$ with the associativity constraint
      $M\blacktriangleleft(V\otimes W)\to(M\blacktriangleleft W)\blacktriangleleft V$
      induced by $J_{VW}$, where $M\in\mathcal{W}_{\theta}$ and $V,W\in{\mathcal{O}}$.
    • The commutativity constraint
      $(M\triangleleft V)\blacktriangleleft W\to(M\blacktriangleleft W)\triangleleft V$
      is induced by $R^{\psi}_{VW}$, where $M\in\mathcal{W}_{\theta}$ and $V,W\in{\mathcal{O}}$.
  2. (2)
    The tensor K-matrix ${{\mathbb{K}}}$ defined in Theorem 5.6.1 is a boundary structure on the bimodule category $\mathcal{W}_{\theta}$ over ${\mathcal{O}}_{{{X}\negthinspace\operatorname{-int}}}$.
Proof.
The notion of a boundary structure on a bimodule category is tailored to encode the datum of the auxiliary R-matrix $R^{\psi}$, the cylindrical structure $(\psi,J,K)$, and the corresponding tensor K-matrix ${{\mathbb{K}}}$. ∎
5.10. Gauged tensor K-matrices
The construction of the tensor K-matrix is compatible with gauging. We have the following analogue of Proposition 3.8.1.
Proposition 5.10.1.
Fix a subset $Z\subseteq I$. Let ${\mathcal{G}}_{Z}$ be the group of invertible elements ${\bf g}\in\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}})$ such that $\operatorname{Ad}({\bf g})$ preserves $U_{q}({\mathfrak{g}})\subset\operatorname{End}({\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}})$. For any ${\bf g}\in{\mathcal{G}}_{Z}$, let $(\psi,J,K)$ be the datum given by
$\psi=\operatorname{Ad}({\bf g}\bm{\gamma})\circ\thinspace\theta_{q}^{-1},\qquad J=({\bf g}\otimes{\bf g})\cdot R_{\theta}\cdot\Delta({\bf g})^{-1},\qquad K={\bf g}\cdot\Upsilon,$ (5.46)
and set
${{\mathbb{K}}}=R_{21}^{\psi}\cdot 1\otimes K\cdot R\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}_{{{Z}\negthinspace\operatorname{-int}}}).$ (5.47)
  1. (1)
    ${{\mathbb{K}}}$ admits a canonical lift in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X\cup Z}\negthinspace\operatorname{-int}}})$ which satisfies the identities (5.18) in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X\cup Z}\negthinspace\operatorname{-int}}})$ and (5.19), (5.20) and (5.21) in $\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{{X\cup Z}\negthinspace\operatorname{-int}}}\boxtimes{\mathsf{f}}_{{{X\cup Z}\negthinspace\operatorname{-int}}})$.
  2. (2)
    Consider the bimodule structure on $\mathcal{W}_{\theta}$ over ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}$ induced by $(\psi,J)$ as in Corollary 5.9.1. Then, ${{\mathbb{K}}}$ is a boundary structure on $\mathcal{W}_{\theta}$ over ${\mathcal{O}}_{{{X\cup Z}\negthinspace\operatorname{-int}}}$.
Proof.
It is enough to observe that ${{\mathbb{K}}}=1\otimes{\bf g}\cdot{{\mathbb{K}}}_{{\bf g}=1}$, where ${{\mathbb{K}}}_{{\bf g}=1}=R_{21}^{\phi_{q}^{-1}}\cdot 1\otimes\Upsilon\cdot R$ is the tensor K-matrix from Theorem 5.6.1. ∎
Remark 5.10.2.
The gauge group ${\mathcal{G}}_{Z}$ acts simultaneously on both the bimodule and the boundary structures of $\mathcal{W}_{\theta}$ over ${\mathcal{O}}_{{{Z}\negthinspace\operatorname{-int}}}$. $\triangledown$
6. Trigonometric tensor K-matrices
In [AV22b] we showed that, for ${\mathfrak{g}}$ of untwisted affine type and for suitable twist automorphisms $\psi$, the universal K-matrix $K$ gives rise to a formal Laurent series with a well-defined action on finite-dimensional modules of $U_{q}({\mathfrak{g}})$. Here we consider similar questions for the tensor K-matrix ${{\mathbb{K}}}$.
6.1. Quantum loop algebras and finite-dimensional modules
Let ${\mathfrak{g}}$ be an untwisted affine Kac-Moody algebra. By [Kac90], ${\mathfrak{g}}^{\prime}$ is a central extension of the loop algebra ${\mathfrak{L}}=\mathring{{\mathfrak{g}}}[t,t^{-1}]$, where $\mathring{{\mathfrak{g}}}={\mathfrak{g}}_{I\backslash\{0\}}\subset{\mathfrak{g}}$ is a distinguished finite-dimensional Lie subalgebra appearing where the node $0\in I$ has been chosen as in [Kac90]. Let ${\mathcal{C}}\subset\mathcal{W}$ be the category of finite-dimensional (type $\bf 1$) $U_{q}({\mathfrak{g}}^{\prime})$-modules and ${\mathsf{f}}_{{\mathcal{C}}}^{n}\colon{\mathcal{C}}^{\boxtimes n}\to\operatorname{Vect}$ the corresponding $n$-fold forgetful functor.
Let $\delta=\sum_{i\in I}a_{i}\alpha_{i}$ be the minimal imaginary root, i.e., the unique element of ${\operatorname{Ker}}(A)\cap\mathsf{Q}^{+}$ with setwise coprime coefficients, and set $K_{\delta}=\prod_{i\in I}K_{i}^{a_{i}}\in U_{q}({\mathfrak{g}}^{\prime})$. The element $K_{\delta}$ is central and acts trivially on any module in ${\mathcal{C}}$. Therefore, ${\mathcal{C}}$ is equivalent to the category of finite-dimensional (type $\bf 1$) modules over the quotient Hopf algebra $U_{q}(\mathfrak{L})=U_{q}({\mathfrak{g}}^{\prime})/(K_{\delta}-1)$, which is commonly referred to as the quantum loop algebra.
6.2. Trigonometric R-matrices [Dri86, FR92, KS95]
Let $z$ be a formal variable and consider the $\tau$-minimal grading shift $\Sigma_{z}:U_{q}(\mathfrak{L})\to U_{q}(\mathfrak{L})[z,z^{-1}]$ defined by
$\Sigma_{z}(E_{i})=z^{s_{i}}E_{i},\quad\Sigma_{z}(F_{i})=z^{-s_{i}}F_{i}\,\quad\mbox{and}\quad\Sigma_{z}(K_{i})=K_{i}$
where $s_{0}=s_{\tau(0)}=1$ and $s_{i}=0$ if $i\notin\{0,\tau(0)\}$. For any $U_{q}(\mathfrak{L})$-module $V$, we consider the modules $V(z)=V\otimes{\mathbb{F}}(z)$ and ${V}(\negthinspace(z)\negthinspace)=V\otimes{{\mathbb{F}}}(\negthinspace(z)\negthinspace)$ with action given by the pullback through $\Sigma_{z}$ and extension of scalars.
Let $R\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}})$ be the universal R-matrix of $U_{q}({\mathfrak{g}})$, see Section 2.6. Then,
$R(z)=({\operatorname{id}}\otimes\Sigma_{z})(R)=(\Sigma_{z^{-1}}\otimes{\operatorname{id}})(R)\in{\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}[\negthinspace[z]\negthinspace]$ (6.1)
that is $R(z)$ is a formal power series in $z$ whose coefficients have a well-defined and natural action on any tensor product in ${\mathcal{C}}$. Moreover, $R(z)$ satisfies the spectral Yang-Baxter equation
$R(w)_{12}\cdot R(wz)_{13}\cdot R(z)_{23}=R(z)_{23}\cdot R(wz)_{13}\cdot R(w)_{12}$ (6.2)
in ${\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}^{3})}[\negthinspace[w,z]\negthinspace]$. Thus, for any $U,V,W\in{\mathcal{C}}$, one obtains a matrix-valued formal series satisfying (6.2) in ${\operatorname{End}(U\otimes V\otimes W)}[\negthinspace[w,z]\negthinspace]$.
By [KS95, Sec. 4.2] and [Cha02, Thm. 3], when $V,W\in{\mathcal{C}}$ are irreducible, the tensor product $V\otimes W$ is generically irreducible, i.e., $V\otimes{W}(\negthinspace(z)\negthinspace)$ remains irreducible as a module over ${U_{q}(\mathfrak{L})}(\negthinspace(z)\negthinspace)$. This yields a factorization, uniquely defined up to a factor in ${\mathbb{F}}^{\times}$, of the form
$R_{VW}(z)=f_{VW}(z)\cdot{\mathbf{R}}_{VW}(z),$ (6.3)
where $f_{VW}(z)\in{{\mathbb{F}}}(\negthinspace(z)\negthinspace)$ and ${\mathbf{R}}_{VW}(z)\in\operatorname{End}(V\otimes W)[z]$ is a non-vanishing operator, commonly referred to as a trigonometric R-matrix.
Since irreducible modules are highest weight and the R-matrix is weight zero, it is natural to normalize ${\mathbf{R}}_{VW}(z)$ such that the tensor product of the highest weight vectors has eigenvalue $1$. This yields a non-vanishing operator $\mathbf{R}^{\operatorname{norm}}_{VW}(z)$ satisfying the unitarity relation
[…]
Moreover, if $V\otimes W(\zeta)$ is irreducible for some $\zeta\in{\mathbb{F}}^{\times}$, then $\mathbf{R}^{\operatorname{norm}}_{VW}(\zeta)$ is well-defined and invertible.
6.3. Quantum affine symmetric pairs and categories of representations
Let $\theta$ be an automorphism of the second kind of ${\mathfrak{g}}$ of the form (3.1), $\theta_{q}$ the corresponding automorphism of $U_{q}({\mathfrak{g}})$ defined in (3.7), and $U_{q}(\mathfrak{k})\subseteq U_{q}({\mathfrak{g}})$ the corresponding QSP subalgebra.
Since $\delta$ is fixed by both the Weyl group and the diagram automorphisms, one has $\theta(\delta)=-\delta$. In particular, $[\delta]_{\theta}=[0]_{\theta}$, where $[\cdot]_{\theta}$ is the projection to the quotient lattice $\mathsf{P}_{\theta}$, see (5.1). Therefore, $\theta_{q}$ descends to an automorphism of $U_{q}(\mathfrak{L})$, and $U_{q}(\mathfrak{k})$ embeds in $U_{q}(\mathfrak{L})$ as a coideal subalgebra. We obtain the following analogue of Proposition 5.2.3 (3).
Lemma 6.3.1.
The category $\mathcal{W}_{\theta}$ of QSP weight modules is a right module category over the category ${\mathcal{C}}$ of finite-dimensional $U_{q}(\mathfrak{L})$-modules in $\mathcal{W}$.
By Proposition 5.2.3 (2), up to twisting, irreducible finite-dimensional $U_{q}(\mathfrak{k})$-modules are objects in $\mathcal{W}_{\theta}$. Now denote by ${\mathcal{C}}_{\theta}$ the full subcategory of finite-dimensional $U_{q}(\mathfrak{k})$-modules in $\mathcal{W}_{\theta}$. Clearly, it is a right module category over ${\mathcal{C}}$.
6.4. Spectral tensor K-matrices
Let ${\mathcal{G}}_{\theta,\bm{\gamma}}$ be the set of gauge transformations of the form ${\bf g}=S_{Y}^{-1}S_{X}{\bm{\beta}}^{-1}$ where
  • $Y\subseteq I\backslash\{0,\tau(0)\}$;
  • ${\bm{\beta}}\colon\mathsf{P}\to{\mathbb{F}}^{\times}$ satisfies ${\bm{\beta}}(\delta)=\bm{\gamma}(\delta)$.
Note that the identity belongs to ${\mathcal{G}}_{\theta,\bm{\gamma}}$ if and only if $0\not\in X$. Let ${\bf g}\in{\mathcal{G}}_{\theta,\bm{\gamma}}$ and set
$\psi=\operatorname{Ad}({\bf g}\bm{\gamma})\circ\thinspace\theta_{q}^{-1},\quad J=({\bf g}\otimes{\bf g})\cdot R_{\theta}\cdot\Delta({\bf g})^{-1}\quad\mbox{and}\quad K={\bf g}\cdot\Upsilon,$ (6.5)
cf. Proposition 5.10.1.
Lemma 6.4.1.
  1. (1)
    $\Sigma_{z}\circ\psi=\psi\circ\Sigma_{z^{-1}}$.
  2. (2)
    $(\Sigma_{z}\otimes{\operatorname{id}})(R^{\psi})=R(z)^{\psi}$
  3. (3)
    $(\Sigma_{z}\otimes{\operatorname{id}})(J)=J$.
  4. (4)
    $K(z)=\Sigma_{z}(K)\in{\operatorname{End}(f)}(\negthinspace(z)\negthinspace)$. If $0\not\in X$, $K(z)\in{\operatorname{End}(f)}[\negthinspace[z]\negthinspace]$.
Proof.
  1. (1)
    It is enough to observe that $\Sigma_{z}$ is the $\tau$-minimal grading shift.
  2. (2)
    This follows from (1) and (6.1).
  3. (3)
    Since ${\bf g}=S_{Y}^{-1}S_{X}{\bm{\beta}}^{-1}$ and $R_{\theta}=(\mathcal{S}_{\theta}\otimes\mathcal{S}_{\theta})^{-1}\cdot\Delta(\mathcal{S}_{\theta})$ by (3.6), the Drinfeld twist $J=({\bf g}\otimes{\bf g})\cdot R_{\theta}\cdot\Delta({\bf g})^{-1}$ is given by a Cartan correction of the quasi R-matrix $\Theta_{Y}$ of $U_{q}({\mathfrak{g}}_{Y})$. The result follows, since $0,\tau(0)\not\in Y$.
  4. (4)
    This is readily verified by direct inspection, see e.g., [AV22b, Sec. 4]. ∎
Let $\Delta_{z}=({\operatorname{id}}\otimes\Sigma_{z})\circ\Delta$ be the shifted coproduct. We have the following.
Theorem 6.4.2.
Let ${{\mathbb{K}}}=R^{\psi}_{21}\cdot 1\otimes K\cdot R\in\operatorname{End}({\mathsf{f}}\boxtimes{\mathsf{f}}_{X\cup Y})$ be the tensor K-matrix corresponding to ${\bf g}\in{\mathcal{G}}_{\theta,\bm{\gamma}}$ constructed in Theorem 5.6.1. Set
${{\mathbb{K}}}(z)=({\operatorname{id}}\otimes\Sigma_{z})({{\mathbb{K}}})=R(z)_{21}^{\psi}\cdot 1\otimes K(z)\cdot R(z)$ (6.6)
in ${\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z)\negthinspace)$ Then, ${{\mathbb{K}}}(z)$ has a canonical lift in ${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z)\negthinspace)$, which satisfies the following identities:
  • the intertwining identity
    ${{\mathbb{K}}}(z)\cdot\Delta_{z}(b)$ $=({\operatorname{id}}\otimes\psi)(\Delta_{z^{-1}}(b))\cdot{{\mathbb{K}}}(z)\qquad(b\in U_{q}(\mathfrak{k}))$ (6.7)
    in ${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z)\negthinspace)$;
  • the coproduct identities
    $(\Delta_{z_{1}}\otimes{\operatorname{id}})({{\mathbb{K}}}(z_{2}))$ $=R(z_{1}z_{2})^{\psi}_{32}\cdot{{\mathbb{K}}}(z_{2})_{13}\cdot R(\tfrac{z_{2}}{z_{1}})_{23},$ (6.8)
    $({\operatorname{id}}\otimes\Delta_{z_{2}/z_{1}})({{\mathbb{K}}}(z_{1}))$ $=J_{23}^{-1}\cdot{{\mathbb{K}}}(z_{2})_{13}\cdot R(z_{1}z_{2})^{\psi}_{23}\cdot{{\mathbb{K}}}(z_{1})_{12},$ (6.9)
    in ${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z_{1},z_{2}/z_{1})\negthinspace)$;
  • the generalized reflection equation
    $R(\tfrac{z_{2}}{z_{1}})^{\psi\psi}_{32}\cdot{{\mathbb{K}}}(z_{2})_{13}\cdot R(z_{1}z_{2})^{\psi}_{23}\cdot{{\mathbb{K}}}(z_{1})_{12}$ $={{\mathbb{K}}}(z_{1})_{12}\cdot R(z_{1}z_{2})^{\psi}_{32}\cdot{{\mathbb{K}}}(z_{2})_{13}\cdot R(\tfrac{z_{2}}{z_{1}})_{23}$ (6.10)
    in ${\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z_{1},z_{2}/z_{1})\negthinspace)$.
Moreover, if $0\not\in X$, then ${{\mathbb{K}}}(z)\in{\operatorname{End}({\mathsf{f}}_{\theta}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}[\negthinspace[z]\negthinspace]$.
Proof.
It is enough to observe that $S_{X},S_{Y}$ are well-defined operators on finite-dimensional $U_{q}(\mathfrak{L})$-modules, ${\bf g}$ acts on $V(z)$ through Laurent polynomials, and therefore ${{\mathbb{K}}}(z)\in{\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})}(\negthinspace(z)\negthinspace)$. Then, the result readily follows from Theorem 5.6.1. ∎
6.5. Trigonometric tensor K-matrices
The construction of the spectral tensor K-matrix from Theorem 6.4.2 immediately implies the following.
Lemma 6.5.1.
Let $M\in{\mathcal{C}}_{\theta}$ and $V\in{\mathcal{C}}$. Then, there exists a QSP intertwiner $\mathbb{K}^{\operatorname{trig}}_{MV}(z)\colon M\otimes V(z)\to M\otimes V^{\psi}(z^{-1})$.
Proof.
The proof is standard, see, e.g., [AV22b, Lemma 5.1.1]. ∎
At present, in general it is unclear if $\mathbb{K}^{\operatorname{trig}}_{MV}(z)$ is obtained by normalizing the action of ${{\mathbb{K}}}(z)$ on $M\otimes{V}(\negthinspace(z)\negthinspace)$. When $M$ and $V$ are irreducible, it would follow automatically if the tensor product $M\otimes V$ is (generically) irreducible over $U_{q}(\mathfrak{k})$, which is currently out of reach.
Nonetheless, by relying on the construction of the trigonometric K-matrices carried out in [AV22b], we can prove it when $M\in{\mathcal{C}}_{\theta}$ is the restriction of an irreducible $U_{q}(\mathfrak{L})$-module. Namely, in [AV22b, Thm. 5.1.2 (2)], we proved that, for any irreducible $V\in{\mathcal{C}}$, there exist a formal Laurent series $g_{V}(z)\in{{\mathbb{F}}}(\negthinspace(z)\negthinspace)$ and a non-vanishing operator-valued polynomial $\mathbf{K}_{V}(z)\in\operatorname{End}(V)[z]$, uniquely defined up to a scalar in ${\mathbb{F}}^{\times}$, such that
$K_{V}(z)=g_{V}(z)\cdot\mathbf{K}_{V}(z).$ (6.11)
Then, we have the following.
Theorem 6.5.2.
Let $M\in{\mathcal{C}}_{\theta}$ be the restriction of an irreducible $U_{q}(\mathfrak{L})$-module. For any irreducible $V\in{\mathcal{C}}$, set
$\mathbb{K}^{\operatorname{trig}}_{MV}(z)=(\mathbf{R}^{\operatorname{norm}}_{V^{\psi}M}(z))_{21}\cdot({\operatorname{id}}_{M}\otimes\mathbf{K}_{V}(z))\cdot\mathbf{R}^{\operatorname{norm}}_{MV}(z)$ (6.12)
in $\operatorname{End}(M\otimes V)(z)$, where $\mathbf{R}^{\operatorname{norm}}(z)$ denotes the unitary R-matrix from Section 6.2.
  1. (1)
    The operator-valued rational function $\mathbb{K}^{\operatorname{trig}}_{MV}(z)$ is non-vanishing and there exist a formal Laurent series $h_{MV}(z)\in{{\mathbb{F}}}(\negthinspace(z)\negthinspace)$ such that
    ${{\mathbb{K}}}_{MV}(z)=h_{MV}(z)\cdot\mathbb{K}^{\operatorname{trig}}_{MV}(z).$ (6.13)
  2. (2)
    For any irreducible $V,W\in{\mathcal{C}}$, the generalized spectral reflection equation hold
    ${\mathbf{R}^{\operatorname{norm}}_{{W^{\psi}}{V^{\psi}}}(\tfrac{z_{2}}{z_{1}})}_{32}\cdot\mathbb{K}^{\operatorname{trig}}_{MW}(z_{2})_{13}\cdot\mathbf{R}^{\operatorname{norm}}_{{V^{\psi}}W}(z_{1}z_{2})_{23}\cdot\mathbb{K}^{\operatorname{trig}}_{MV}(z_{1})_{12}=$ (6.14)
    $\qquad=\mathbb{K}^{\operatorname{trig}}_{MV}(z_{1})_{12}\cdot\mathbf{R}^{\operatorname{norm}}_{{W^{\psi}}V}(z_{1}z_{2})_{32}\cdot\mathbb{K}^{\operatorname{trig}}_{MW}(z_{2})_{13}\cdot\mathbf{R}^{\operatorname{norm}}_{VW}(\tfrac{z_{2}}{z_{1}})_{23}.$
Proof.
The result follows immediately from (6.6) and [AV22b, Thm. 5.1.2]. ∎
Remarks 6.5.3.
  1. (1)
    Consider the q-Onsager algebra, which is the QSP subalgebra $U_{q}(\mathfrak{k})\subset U_{q}(\widehat{\mathfrak{sl}}_{2})$ with $\theta=\omega$, see [BK05]. In [IT09], Ito and Terwilliger proved that, up to isomorphism, every finite-dimensional irreducible module over $U_{q}(\mathfrak{k})$ is the restriction of a $U_{q}(\widehat{\mathfrak{sl}}_{2})$-module. This implies that, in this case, Theorem 6.5.2 is valid for every irreducible module in ${\mathcal{C}}_{\theta}$.
  2. (2)
    For more general modules in $M\in{\mathcal{C}}_{\theta}$, Theorem 6.5.2 follows at once whenever the space of QSP intertwiners
    $M\otimes{V}(\negthinspace(z)\negthinspace)\to M\otimes{V}(\negthinspace(z)\negthinspace)^{\psi}$
    is proved to be one-dimensional. $\triangledown$
6.6. Unitary tensor K-matrices
The construction of unitary K-matrices requires to consider only $V\in{\mathcal{C}}$ such that $V^{\psi^{2}}=({V^{\psi}})^{\psi}$ is isomorphic to $V$ as a $U_{q}(\mathfrak{k})$-module. This constraint yields no loss of generality. Choosing the gauge transformation ${\bf g}=\mathcal{S}_{\theta}\cdot\bm{\gamma}^{-1}\in{\mathcal{G}}_{\theta,\bm{\gamma}}$, we have $\psi=\omega\circ\tau$. In particular, $\psi$ is an involution so that $V^{\psi^{2}}=V$ for every $V\in{\mathcal{C}}$.
In [AV22b, Prop. 5.4.1], we proved that, in this setting, the operators $\mathbf{K}_{V}(z)$ and $\mathbf{K}_{V^{\psi}}(z)$ can be normalized to a pair of unitary K-matrices $\mathbf{K}^{\operatorname{norm}}_{V}(z)$ and $\mathbf{K}^{\operatorname{norm}}_{V^{\psi}}(z)$. This yields the following.
Proposition 6.6.1.
Let $M\in{\mathcal{C}}_{\theta}$ be the restriction of an irreducible $U_{q}(\mathfrak{L})$-module. For any irreducible $V\in{\mathcal{C}}$ such that $V^{\psi^{2}}\cong V$, set
$\mathbb{K}^{\operatorname{norm}}_{MV}(z)=(\mathbf{R}^{\operatorname{norm}}_{V^{\psi}M}(z))_{21}\cdot({\operatorname{id}}_{M}\otimes\mathbf{K}^{\operatorname{norm}}_{V}(z))\cdot\mathbf{R}^{\operatorname{norm}}_{MV}(z)$ (6.15)
in $\operatorname{End}(M\otimes V)(z)$. Then,
$\mathbb{K}^{\operatorname{norm}}_{MV}(z)^{-1}=\mathbb{K}^{\operatorname{norm}}_{MV^{\psi}}(z^{-1})\,.$ (6.16)
Moreover, if $V(\zeta)$ is $U_{q}(\mathfrak{k})$-irreducible and $M\otimes V(\zeta)$, $V^{\psi}(\zeta)\otimes M$ are $U_{q}(\mathfrak{L})$-irreducible for some $\zeta\in{\mathbb{F}}^{\times}$, then $\mathbb{K}^{\operatorname{norm}}_{MV}(z)$ is well-defined and invertible.
Remark 6.6.2.
As before, for more general modules in $M\in{\mathcal{C}}_{\theta}$, Proposition 6.6.1 follows at once whenever the space of QSP intertwiners
$M\otimes{V}(\negthinspace(z)\negthinspace)\to M\otimes{V}(\negthinspace(z)\negthinspace)^{\psi}$
is proved to be one-dimensional. $\triangledown$
Appendix A Finite-dimensional $U_{q}(\mathfrak{k})$-modules
In this section we will sometimes write $U_{q}({\mathfrak{k}})_{\bm{\gamma}}$ to emphasize the dependence of the QSP subalgebra on the tuple $\bm{\gamma}\in\Gamma$. Here we will show that, up to twisting by straightforward algebra automorphisms of $U_{q}(\mathfrak{k})$, finite-dimensional irreducible $U_{q}(\mathfrak{k})$-modules are QSP weight modules. Note that restriction of the action induces a functor ${\operatorname{Mod}_{\sf fd}}(U_{q}({\mathfrak{g}}))\to{\operatorname{Mod}_{\sf fd}}(U_{q}(\mathfrak{k}))$ and in affine type, we in fact have a functor ${\operatorname{Mod}_{\sf fd}}(U_{q}({\mathfrak{g}}^{\prime}))\to{\operatorname{Mod}_{\sf fd}}(U_{q}(\mathfrak{k}))$, by Remark 3.5.1 (iii). The full description of ${\operatorname{Mod}_{\sf fd}}(U_{q}(\mathfrak{k}))$ for general ${\mathfrak{k}}$ is an open problem which we do not discuss here (for the case $\dim({\mathfrak{g}})<\infty$, see [Wa21] for recent developments and a survey).
Recall that on any finite-dimensional module $V$ of $U_{q}({\mathfrak{g}})$ with $\dim({\mathfrak{g}})<\infty$ the generators $K_{i}$ act via scalar multiplication by signed integer powers of $q$, see e.g., [CP95, Sec. 10.1]. Given $\bm{\eta}_{i}\in\{\pm 1\}$, consider the unique algebra automorphism $\Psi_{\bm{\eta}}$ of $U_{q}({\mathfrak{g}})$ such that
$\Psi_{\bm{\eta}}(E_{i})=\eta_{i}E_{i},\qquad\Psi_{\bm{\eta}}(F_{i})=F_{i},\qquad\Psi_{\bm{\eta}}(K_{i}^{\pm 1})=\eta_{i}K_{i}^{\pm 1}.$ (A.1)
As a consequence, for all finite-dimensional modules $V$ there exists $\bm{\eta}\in\{\pm 1\}^{I}$ such that $\Psi_{\bm{\eta}}^{*}(V)$ is a type $\bf 1$ weight module. We now set out to give a QSP analogue of this statement for irreducible finite-dimensional $U_{q}(\mathfrak{k})$-modules (we do not restrict to the case $\dim({\mathfrak{g}})<\infty$), with $\mathcal{W}_{\theta}$ substituting for the category of type $\bf 1$ weight modules of $U_{q}({\mathfrak{g}})$.
A.1. Weak QSP weight vectors
First, we consider a weaker notion of QSP weight module. Namely, for a $U_{q}(\mathfrak{k})$-module $M$ and a multiplicative character $\beta\in\operatorname{Hom}((\mathsf{Q}^{\vee}_{\sf ext})^{\theta},{\mathbb{F}}^{\times})$, denote the corresponding weak QSP weight space by
$M(\beta)=\{m\in M\,|\,K_{h}\cdot m=\beta(h)m\text{ for all }h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}\}.$ (A.2)
We call $m\in M$ a weak QSP weight vector if it is nonzero and lies in $M(\beta)$ for some $\beta\in\operatorname{Hom}((\mathsf{Q}^{\vee}_{\sf ext})^{\theta},{\mathbb{F}}^{\times})$ and we call $M$ a weak QSP weight module if as a $U_{q}({\mathfrak{h}}^{\theta})$-module it decomposes as a direct sum of weak QSP weight spaces. As in Proposition 5.2.3 (1) one obtains
$B_{i}\cdot M(\beta)\subseteq M(q^{-[\alpha_{i}]_{\theta}}\beta),\qquad E_{j}\cdot M(\beta)\subseteq M(q^{[\alpha_{j}]_{\theta}}\beta)$ (A.3)
for all $i\in I$, $j\in X$, where we note $[\alpha_{i}]_{\theta}\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}$ for all $i\in I$. Quite straightforwardly one obtains, relying on the algebraic closedness of ${\mathbb{F}}$, any irreducible finite-dimensional $U_{q}(\mathfrak{k})$-module has a nonzero weak QSP weight space and hence, by (A.3), is a weak QSP weight module.
We will now deduce a more powerful statement, still in full generality. For this it is convenient to describe $(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}$, and hence $U_{q}({\mathfrak{h}}^{\theta})$, more explicitly as follows (see e.g. [RV22, Lem. 4.8]). Fix a set $I_{\theta}$ of representatives of the nontrivial $\tau$-orbits of $I\backslash X$. Then
$(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}=\operatorname{Sp}_{\mathbb{Z}}\big{\{}[\alpha_{i}]_{\theta}\,\big{|}\,i\in I_{\theta}\cup X\big{\}}$
and the commutative subalgebra $U_{q}({\mathfrak{h}}^{\theta})\subset U_{q}(\mathfrak{k})$ is generated by $K_{j}^{\pm 1}$ ($j\in X$) and $(K_{i}K_{\tau(i)}^{-1})^{\pm 1}$ ($i\in I_{\theta}$). We call the elements of
$\{E_{j},F_{j},K_{j}^{\pm 1}\,|\,j\in X\}\cup\{B_{\tau(i)},B_{i},(K_{i}K_{\tau(i)}^{-1})^{\pm 1}\,|\,i\in I_{\theta}\}\cup\{B_{i}\,|\,i\in I\backslash X,\tau(i)=i\}$
the canonical generators of $U_{q}(\mathfrak{k})$.
Consider $(\mathsf{Q}^{\vee}_{\sf ext})^{\theta,+}=\operatorname{Sp}_{{\mathbb{Z}}_{\geq 0}}\big{\{}[\alpha_{i}]_{\theta}\,\big{|}\,i\in I_{\theta}\cup X\big{\}}$. Define a partial order $\geq_{\theta}$ on $\operatorname{Hom}((\mathsf{Q}^{\vee}_{\sf ext})^{\theta},{\mathbb{F}}^{\times})$ as follows
$\beta\geq_{\theta}\beta^{\prime}\qquad\Longleftrightarrow\qquad\beta/\beta^{\prime}\in q^{(\mathsf{Q}^{\vee}_{\sf ext})^{\theta,+}}.$
Lemma A.1.1.
Let $M\in{\operatorname{Mod}_{\sf fd}}(U_{q}(\mathfrak{k}))$. Then the generators $B_{\tau(i)}$, $B_{i}$ ($i\in I_{\theta}$) and $E_{j}$, $F_{j}$ ($j\in X$) act locally finitely on $M$.
Now assume $M$ is irreducible. Then it is a weak QSP weight module. Furthermore, $M$ is generated by a weak QSP weight vector $m_{0}$ which is annihilated by the action of $B_{\tau(i)}$ and $E_{j}$ for all $i\in I_{\theta}$ and $j\in X$ and the eigenvalues of the action of the following elements of $U_{q}({\mathfrak{h}}^{\theta})$ on $m_{0}$ lie in ±$\pm q^{\mathbb{Z}}$: $K_{j}$ for all $j$ in $X$ and $K_{i}K_{\tau(i)}^{-1}$ for all $i$ in
$I_{\theta}^{0}:=\big{\{}i\in I_{\theta}\,\big{|}\,(\theta(\alpha_{i}),\alpha_{i})=0\big{\}}.$
Proof.
This is a variation on a standard argument. Note that $M$ decomposes as a direct sum of generalized weak QSP weight spaces
$M^{\sf gen}(\beta)=\{m\in M\,|\,\text{for all }h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}\text{ and all }r>>0,(K_{h}-\beta(h))^{r}\cdot m=0\}$ (A.4)
where $\beta\in\operatorname{Hom}((\mathsf{Q}^{\vee}_{\sf ext})^{\theta},{\mathbb{F}}^{\times})$. The relations (5.5) imply that
$B_{i}\cdot M^{\sf gen}(\beta)\subseteq M^{\sf gen}\big{(}q_{i}^{-[\alpha_{i}]_{\theta}}\beta\big{)},\qquad E_{j}\cdot M^{\sf gen}(\beta)\subseteq M^{\sf gen}\big{(}q_{j}^{[\alpha_{j}]_{\theta}}\beta\big{)}$
for all $i\in I$, $j\in X$ and $\beta\in\operatorname{Hom}((\mathsf{Q}^{\vee}_{\sf ext})^{\theta},{\mathbb{F}}^{\times})$. Note that $-[\alpha_{\tau(i)}]_{\theta}=[w_{X}(\alpha_{i})]_{\theta}$ for $i\in I_{\theta}$. Owing to relations (5.5) the action of $B_{\tau(i)}$ and $E_{j}$ on generalized weak QSP weight spaces will strictly increase the weight with respect to $\geq_{\theta}$ and the action of $B_{i}$ and $F_{j}$ will strictly decrease the weight, where $i$ is an arbitrary element of $I_{\theta}$ and $j$ is an arbitrary element of $X$. Since $\dim(M)<\infty$, we obtain the local nilpotency and the existence of a nonzero element of $M$ annihilated by, say, $B_{\tau(i)}$ for all $i\in I_{\theta}$ and $E_{j}$ for all $j\in X$.
Consider the nonzero subspace
$M_{0}=\{m\in M\,|\,B_{\tau(i)}\cdot m=E_{j}\cdot m=0\text{ for all }i\in I_{\theta},\,j\in\}.$
By (5.5), $M_{0}$ is a $U_{q}({\mathfrak{h}}^{\theta})$-submodule. Since $\dim(M_{0})<\infty$ and ${\mathbb{F}}$ is algebraically closed, $M_{0}$ contains a joint eigenvector $m_{0}$ of all $K_{h}$ ($h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta}$), i.e. $m_{0}$ is a weak QSP weight vector. Since the canonical generators of $U_{q}(\mathfrak{k})$ send weak QSP weight spaces to weak QSP weight spaces, we obtain that $U_{q}(\mathfrak{k})\cdot m_{0}$ is a weak QSP weight module which must equal $M$ by irreducibility.
Finally, for $i\in I_{\theta}^{0}$, [BK15, Thm. 3.6] implies the relation
$[B_{\tau(i)},B_{i}]=c\frac{K_{i}K_{\tau(i)}^{-1}-K_{i}^{-1}K_{\tau(i)}}{q_{i}-q_{i}^{-1}}$ (A.5)
where $c\in{\mathbb{F}}^{\times}$ is proportional to ${\bm{\gamma}}_{i}={\bm{\gamma}}_{\tau(i)}$. Hence, for all $i\in I_{\theta}^{0}$ the subalgebra ${\mathbb{F}}\langle B_{\tau(i)},\,B_{i},\,(K_{i}K_{\tau(i)}^{-1})^{\pm 1}\rangle\subseteq U_{q}(\mathfrak{k})$ is isomorphic to $U_{q}(\mathfrak{sl}_{2})$. Hence, a standard $U_{q}(\mathfrak{sl}_{2})$-argument, see e.g. [Jan96, Prop. 2.3], implies the characterization of the eigenvalues of the action on $m_{0}$ of $K_{j}$ ($j\in X$) and $K_{i}K_{\tau(i)}^{-1}$ ($i\in I_{\theta}^{0}$). ∎
A.2. Defining relations of QSP algebras
Recall from the definition of $U_{q}({\mathfrak{g}})$ the noncommutative polynomial $\operatorname{Serre}_{ij}(F_{i},F_{j})$ of degree $(1-a_{ij})\alpha_{i}+\alpha_{j}$. By [Kol14, Thms. 7.1 & 7.3], the QSP algebra $U_{q}({\mathfrak{k}})_{\bm{\gamma}}$ has a presentation as follows. It is generated over $U_{q}({\mathfrak{n}}_{X}^{+})U_{q}({\mathfrak{h}}^{\theta})$ by elements $B_{i}$ ($i\in I$) subject to the relations
$K_{h}B_{i}$ $=q_{i}^{-\alpha_{i}(h)}B_{i}K_{h}$ for $\text{for all }h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\theta},\,i\in I,$ (A.6)
$[E_{j},B_{i}]$ $=\delta_{ij}\frac{K_{i}-K_{i}^{-1}}{q_{i}-q_{i}^{-1}}$ for $\text{for all }i\in I,\,j\in X,$ (A.7)
$\operatorname{Serre}_{ij}(B_{i},B_{j})$ $=C_{ij}$ for $\text{for all }i,j\in I,\,i\neq j,$ (A.8)
where $C_{ij}$ is a noncommutative polynomial over $U_{q}({\mathfrak{n}}_{X}^{+})U_{q}({\mathfrak{h}}^{\theta})$ in $B_{i}$ and $B_{j}$, of degree strictly less than $(1-a_{ij})\alpha_{i}+\alpha_{j}$. If $i\in X$ or if $i\not\in\{\tau(i),\tau(j)\}$ then by [Kol14, Lem. 5.11, Thm. 7.3] we have $\operatorname{Serre}_{ij}(B_{i},B_{j})=0$. For other $(i,j)$, various more explicit forms have been provided for the relation (A.8), see [Let02, Kol14, BK15, DC19, CLW21a, CLW21b, KY21].
Drawing on these works, here we give properties of the $C_{ij}$ which are sufficient to our purposes. For all $i\in I\backslash X$ there exists $Z_{i}\in U_{q}({\mathfrak{n}}^{+}_{X})_{w_{X}(\alpha_{i})-\alpha_{i}}$ such that the following statements are satisfied.
  1. (1)
    By [DC19, Cor. 3.5], for all $i,j\in I\backslash X$ such that $\tau(i)=i\neq j$, $\operatorname{Serre}_{ij}(B_{i},B_{j})$ is a ${\mathbb{F}}$-linear combination of products of one factor $B_{j}$, $M$ factors $B_{i}$ and $(1-a_{ij}-M)/2$ factors $Z_{i}$, where $M$ runs through $\{0,1,\ldots,-1-a_{ij}\}$.
  2. (2)
    By [DC19, Cor. 3.9], for all $i\in I\backslash X$ and $j\in X$ such that $\tau(i)=i$, $\operatorname{Serre}_{ij}(B_{i},B_{j})$ is an ${\mathbb{F}}$-linear combination of products of one factor $B_{j}$, $M$ factors $B_{i}$ and either $(1-a_{ij}-M)/2$ factors $Z_{i}$ or $(-1-a_{ij}-M)/2$ factors $Z_{i}$ and one factor $W_{ij}K_{j}$, where $M$ runs through $\{0,1,\ldots,-1-a_{ij}\}$ and $W_{ij}\in U_{q}({\mathfrak{n}}^{+}_{X})_{w_{X}(\alpha_{i})-\alpha_{i}-\alpha_{j}}$.
  3. (3)
    By [BK15, Thm. 3.6], for all $i\in I_{\theta}\cup\tau(I_{\theta})$, $\operatorname{Serre}_{i\tau(i)}(B_{i},B_{\tau(i)})$ is an ${\mathbb{F}}$-linear combination, with coefficients independent of $\bm{\gamma}$, of $\gamma_{i}B_{i}^{-a_{i\tau(i)}}Z_{i}K_{i}^{-1}K_{\tau(i)}$ and $\gamma_{\tau(i)}B_{i}^{-a_{i\tau(i)}}Z_{\tau(i)}K_{i}K_{\tau(i)}^{-1}$.
This completes the enumeration of all $(i,j)\in I\times I$ such that $i\neq j$, $i\in\{\tau(i),\tau(j)\}$ and $i\notin X$.
Before we use these relations to describe certain automorphisms of $U_{q}(\mathfrak{k})$, we look at a special case which shows that in Lemma A.1.1 the irreducibility assumption is necessary: for some $U_{q}(\mathfrak{k})$ there exist reducible finite-dimensional $U_{q}(\mathfrak{k})$-modules which are not weak QSP weight modules.
Example A.2.1.
Suppose $I=\{0,1\}$, $a_{01}=a_{10}=-2$, so that ${\mathfrak{g}}^{\prime}\cong\widehat{\mathfrak{sl}}_{2}$. Let $\tau$ be the nontrivial diagram automorphism, and consider the Satake diagram $(\emptyset,\tau)$. Specializing the definitions in Sections 3.4-3.5 to this case, we obtain the QSP subalgebra $U_{q}(\mathfrak{k})\subset U_{q}(\widehat{\mathfrak{sl}}_{2})$ generated by
$B_{0}=F_{0}-q\gamma_{0}E_{1}K_{0}^{-1},\qquad B_{1}=F_{1}-q\gamma_{1}E_{0}K_{1}^{-1},\qquad L^{\pm 1}$ (A.9)
where $\gamma_{0},\gamma_{1}\in{\mathbb{F}}^{\times}$ and $L=K_{0}K_{1}^{-1}$. It is also known as augmented q-Onsager algebra. By (A.6-A.8), relying on [BK15, Thm. 3.6] for the explicit expression of $C_{ij}$ in this case (also see [BB17]), this is isomorphic to the algebra generated over ${\mathbb{C}}(q)$ by $B_{0}$, $B_{1}$ and invertible $L$ subject to the relations
$LB_{0}$ $=q^{-4}B_{0}L,\hskip 30.0ptLB_{1}=q^{4}B_{1}L,$ (A.10)
$\operatorname{Serre}_{01}(B_{0},B_{1})$ $=(q+q^{-1})(q^{3}-q^{-3})B_{0}\big{(}\gamma_{0}L^{-1}-\gamma_{1}L\big{)}B_{0},$
$\operatorname{Serre}_{10}(B_{1},B_{0})$ $=(q+q^{-1})(q^{3}-q^{-3})B_{1}\big{(}\gamma_{1}L-\gamma_{0}L^{-1}\big{)}B_{1}.$
Any finite-dimensional ${\mathbb{C}}(q)$-linear space $M$ becomes a $U_{q}(\mathfrak{k})$-module if we let $B_{0}$, $B_{1}$ act by the zero map and $L$ by any invertible operator $L_{M}$. Clearly, this is a weak QSP weight module if and only if $L_{M}$ is diagonalizable. $\triangledown$
A.3. Some automorphisms of QSP algebras
We will use these defining relations to describe certain automorphisms of the QSP algebra which act on the canonical generators by scalar multiplication. The only defining relations which constrain any such algebra automorphism $f$ of $U_{q}(\mathfrak{k})$ are the non-homogeneous relations: (A.7) in the case $i=j\in X$, constraining only the scalar factor for $E_{j}$, $F_{j}$ and $K_{j}^{\pm 1}$ ($j\in X$), the q-Serre relations (A.8) in the case $i\neq\tau(i)=j\in I\backslash X$, which constrains the products of the scalar factors appearing in $f(B_{i})$ and $f(B_{\tau(i)})$ in terms of the induced scalar factors of $Z_{i}$ and $Z_{\tau(i)}$, and the q-Serre relations (A.8) in the case $\tau(i)=i\in I\backslash X$, which only constrains the scalar factor of $B_{i}$ in terms of the scalar factors appearing in $f(Z_{i})$ and $f(W_{ij}K_{j})$.
The following is a special case of [Wa21, Lem. 2.5.1], which we reproduce here in the present conventions.
Lemma A.3.1.
Let $i\in I_{\theta}\backslash I_{\theta}^{0}$ and $\kappa\in{\mathbb{F}}^{\times}$. Given ${\bm{\gamma}}\in\Gamma$, define ${\bm{\gamma}}^{\prime}\in\Gamma$ by
${\bm{\gamma}}^{\prime}_{j}=\begin{cases}\kappa\bm{\gamma}_{i}&\text{if }j=i,\\ \kappa^{-1}\bm{\gamma}_{\tau(i)}&\text{if }j=\tau(i),\\ \bm{\gamma}_{j}&\text{otherwise}.\end{cases}$ (A.11)
There is a unique isomorphism
$U_{q}({\mathfrak{k}})_{\bm{\gamma}}\stackrel{{\scriptstyle f_{i,\kappa}}}{{\longrightarrow}}U_{q}({\mathfrak{k}})_{\bm{\gamma}^{\prime}}$
such that
$f_{i,\kappa}\big{(}(K_{i}K_{\tau(i)}^{-1})^{\pm 1}\big{)}=\kappa^{\pm 1}(K_{i}K_{\tau(i)}^{-1})^{\pm 1}$ (A.12)
and $f_{i,\kappa}$ fixes all other canonical generators.
Proof.
This follows from a direct inspection of the q-Serre relations given in [BK15, Thm. 3.6], see point (3) above. ∎
We can also define certain algebra automorphisms of $U_{q}({\mathfrak{k}})_{\bm{\gamma}}$ analogous to the automorphisms $\Psi_{\bm{\eta}}:U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})$ defined above. First note that for any choice of ${\bm{\eta}}_{j}\in\{\pm 1\}$ ($j\in X$), the assignments
$E_{j}\mapsto{\bm{\eta}}_{j}E_{j},\qquad F_{j}\mapsto F_{j},\qquad K_{j}^{\pm 1}\mapsto{\bm{\eta}}_{j}K_{j}^{\pm 1}$
extend to an algebra automorphism of $U_{q}({\mathfrak{g}}_{X})$. Now define ${\bm{\eta}}_{X}\in\operatorname{Hom}(\mathsf{Q}_{X},{\mathbb{F}}^{\times})$ by ${\bm{\eta}}_{X}(\alpha_{j})={\bm{\eta}}_{j}$. For each $\tau$-orbit outside $X$, i.e. each element of
$I^{*}_{\theta}=I_{\theta}\cup\{i\in I\backslash X\,|\,\tau(i)=i\}$
we need a correcting factor depending on $X$, in addition to a free choice of sign. If $i\in I_{\theta}$ set $C(X,i):={\bm{\eta}}_{X}(w_{X}(\alpha_{i})-\alpha_{i})\in\{\pm 1\}$. On the other hand, if $i\in I\backslash X$ satisfies $\tau(i)=i$ then denote by $C(X,i)$ a fixed square root of ${\bm{\eta}}_{X}(w_{X}(\alpha_{i})-\alpha_{i})\in\{\pm 1\}$.
Lemma A.3.2.
Given ${\bm{\eta}}\in\{\pm 1\}^{X\cup I^{*}_{\theta}}$, the above algebra automorphism of $U_{q}({\mathfrak{g}}_{X})$, depending only on $({\bm{\eta}}_{j})_{j\in X}$, extends to an algebra automorphism
$\Psi_{\theta,{\bm{\eta}}}:U_{q}(\mathfrak{k})\to U_{q}(\mathfrak{k})$
by means of the assigments
$B_{i}$ $\mapsto{\bm{\eta}}_{i}C(X,i)B_{i},$ $i\in I\backslash X,\,\tau(i)=i,$
$(K_{i}K_{\tau(i)}^{-1})^{\pm 1}$ $\mapsto{\bm{\eta}}_{i}(K_{i}K_{\tau(i)}^{-1})^{\pm 1},$ $i\in I_{\theta}$
$B_{\tau(i)}$ $\mapsto{\bm{\eta}}_{i}C(X,i)B_{\tau(i)},$ $i\in I_{\theta}$
$B_{i}$ $\mapsto B_{i},$ $i\in I_{\theta}.$
Proof.
Let $j\in X$. As was the case for the automorphism $\Psi_{\bf\eta}$ of $U_{q}({\mathfrak{g}})$, given that $F_{j}$ is fixed we readily obtain from (A.7) the form of the assignments for $E_{j}$ and $K_{j}^{\pm 1}$ and we obtain an algebra automorphism $\Psi_{({\bm{\eta}}_{j})_{j\in X}}$ of $U_{q}({\mathfrak{g}}_{X})$. Let $i\in I\backslash X$ and $j\in X$ be arbitrary. Since $Z_{i}\in U_{q}({\mathfrak{n}}^{+}_{X})_{w_{X}(\alpha_{i})-\alpha_{i}}$, $W_{ij}\in U_{q}({\mathfrak{n}}^{+}_{X})_{w_{X}(\alpha_{i})-\alpha_{i}-\alpha_{j}}$ and $w_{X}(\alpha_{i})-\alpha_{i}$ is $\tau$-invariant, we see that $\Psi_{({\bm{\eta}}_{j})_{j\in X}}$ acts on $Z_{i}$, $Z_{\tau(i)}$ and $W_{ij}K_{j}$ by scalar multiplication by the same sign ${\bm{\eta}}(w_{X}(\alpha_{i})-\alpha_{i})$.
Let $i\in I_{\theta}$. It follows that $C(X,i)=C(X,\tau(i))$. Given that $B_{i}$ is fixed, the q-Serre relations (A.8) in the case $i\neq\tau(i)=j\in I\backslash X$ require the form of the assignments for $B_{\tau(i)}$ and $(K_{i}K_{\tau(i)}^{-1})^{\pm 1}$.
Finally, let $i\in I\backslash X$ such that $\tau(i)=i$. The q-Serre relations (A.8) in the case $\tau(i)=i\in I\backslash X$ require the form of the assignments for $B_{i}$. ∎
A.4. From weak QSP weight vectors to QSP weight vectors
Putting it all together we arrive at the following result, which implies
Theorem A.4.1.
Let $M\in{\operatorname{Mod}_{\sf fd}}(U_{q}(\mathfrak{k}))$ be irreducible and let $m_{0}$ be as in Lemma A.1.1. Then there exists an algebra automorphism $f$ of $U_{q}(\mathfrak{k})$ such that $m_{0}\in f^{*}M$ is an eigenvector of the action of $K_{j}$ ($j\in X$) and $K_{i}K_{\tau(i)}^{-1}$ ($i\in I_{\tau}$) with eigenvalues in $q^{\mathbb{Z}}$.
Proof.
Let $M$ be a finite-dimensional irreducible $U_{q}(\mathfrak{k})$-module with $m_{0}\in M$ as in Lemma A.1.1. By Lemma A.3.1, there exists a tuple $\bm{\kappa}=({\bm{\kappa}}_{i})\in({\mathbb{F}}^{\times})^{I_{\theta}\backslash I_{\theta}^{0}}$ such that $(f_{\bm{\kappa}})^{*}M$ is a weak QSP weight module with the joint action of $U_{q}({\mathfrak{h}}^{\theta})$ on $m$ given by scalar multiplication by signed powers of $q$, where
$f_{\bm{\kappa}}:=\prod_{i\in I_{\theta}\backslash I_{\theta}^{0}}f_{i,{\bm{\kappa}}_{i}}.$ (A.13)
Now we can use Lemma A.3.2 to remove signs from the eigenvalues. ∎
Appendix B Cylindrical and reflection bialgebras
We outline a generalization of the algebraic approach of [Kol20, Sec. 2] to a braided structure on the category of modules of certain coideal subalgebras of quasitriangular bialgebras. Part of this was developed independently by Lemarthe, Baseilhac and Gainutdinov in [LBG23, Le23] (for comodule algebras over quasitriangular algebras). We also connect with a formalism for comodule algebras over a weakly quasitriangular bialgebra extending its braided structure, as introduced by Kolb and Yakimov in [KY20], which we call here weakly cylindrical.
B.1. Quasitriangular bialgebras [Dri86]
Fix a base field ${\mathbb{F}}$. A quasitriangular bialgebra $(A,R)$ with coproduct $\Delta$ is a pair $(A,R)$ where $A$ is a bialgebra with coproduct $\Delta$ and $R\in A\otimes A$ is a universal R-matrix, i.e., an invertible element satisfying the intertwining identity
$R\cdot\Delta(x)=\Delta^{\operatorname{op}}(x)\cdot R\qquad(x\in A)$ (B.1)
and the coproduct identities
$(\Delta\otimes{\operatorname{id}})(R)=R_{13}\cdot R_{23}\qquad\mbox{and}\qquad({\operatorname{id}}\otimes\Delta)(R)=R_{13}\cdot R_{12}$ (B.2)
where $\Delta^{\sf op}:=(12)\circ\Delta$. Then $R$ is a solution of the Yang-Baxter equation
$R_{12}\cdot R_{13}\cdot R_{23}=R_{23}\cdot R_{13}\cdot R_{12}.$ (B.3)
Moreover, $(\varepsilon\otimes{\operatorname{id}})(R)=1=({\operatorname{id}}\otimes\varepsilon)(R)$, where $\varepsilon$ is the counit of $A$. Note also that the co-opposite $(A^{\operatorname{cop}},R_{21})$ is a quasitriangular bialgebra.
A quasitriangular bialgebra $(A,R)$ can be twisted as follows, also see [Dav07].
  • Let $J\in A\otimes A$ be a Drinfeld twist, i.e., an invertible element satisfying $(J\otimes 1)\cdot(\Delta\otimes{\operatorname{id}})(J)=(1\otimes J)\cdot({\operatorname{id}}\otimes\Delta)(J)$ and $(\varepsilon\otimes{\operatorname{id}})(J)=1=({\operatorname{id}}\otimes\varepsilon)(J)$. There is a quasitriangular bialgebra $(A_{J},R_{J})$ such that $A_{J}=A$ as an algebra, its coproduct is $\Delta_{J}=\operatorname{Ad}(J)\circ\Delta$, the counit is $\varepsilon$, and $R_{J}=J_{21}\cdot R\cdot J^{-1}$.
  • Let $\psi:A\to A$ be an algebra automorphism. There is a quasitriangular bialgebra $(A^{\psi},R^{\psi\psi})$ such that $A^{\psi}=A$ as an algebra, its coproduct is $\Delta^{\psi}=(\psi\otimes\psi)\circ\Delta\circ\psi^{-1}$, the counit is $\varepsilon^{\psi}=\varepsilon\circ\psi^{-1}$, and $R^{\psi\psi}=(\psi\otimes\psi)(R)$. By construction, $\psi$ is an isomorphism of quasitriangular bialgebras $(A,R)\to(A^{\psi},R^{\psi\psi})$.
If $A^{\operatorname{cop},\psi}=A_{J}$ and $R_{21}^{\psi\psi}=R_{J}$, the pair $(\psi,J)$ is referred to as a twist pair, see [AV22a].
Remark B.1.1.
Note that, up to right-multiplication by elements of the centralizer of $\Delta(A)$ in $A\otimes A$, $J$ is uniquely determined by $\psi$ via the constraint $(\Delta^{\operatorname{op}})^{\psi}=\Delta_{J}$, in the same way that $R$ is almost uniquely determined by (B.1). $\triangledown$
B.2. Cylindrical bialgebras [AV22a]
A cylindrical bialgebra $(A,R,B,\psi,J,K)$ is the datum of
  • a quasitriangular bialgebra $(A,R)$;
  • a (right) coideal subalgebra $B\subseteq A$, i.e., $\Delta(B)\subseteq B\otimes A$;
  • a twist pair $(\psi,J)$;
  • a (basic) universal K-matrix $K\in A$ with respect to $(B,\psi,J)$, i.e., an invertible element satisfying the QSP intertwining identity
    $K\cdot b=\psi(b)\cdot K\qquad(b\in B)$ (B.4)
    and the coproduct identity
    $\Delta(K)=J^{-1}\cdot(1\otimes K)\cdot R^{\psi}\cdot(K\otimes 1).$ (B.5)
Equivalently, we shall also say that $(\psi,J,K)$ is a cylindrical structure on $(A,R)$ with respect to $B$. It follows, see [AV22a, Prop. 2.4], from combining (B.1), (B.5) and the twist pair property that $K$ satisfies the generalized reflection equation:
$R^{\psi\psi}_{21}\cdot(1\otimes K)\cdot R^{\psi}\cdot(K\otimes 1)=(K\otimes 1)\cdot R^{\psi}_{21}\cdot(1\otimes K)\cdot R.$ (B.6)
Moreover, $\varepsilon(K)=1$. This forms a generalization of the notion of cylinder-braided coideal subalgebra, due to Balagović and Kolb in [BK19, Def. 4.10], which in turn generalizes the notion of cylinder twist, due to tom Dieck and Häring-Oldenburg in [tD98, tDHO98].
Example B.2.1.
Universal R-matrices are examples of basic universal K-matrices, cf. [BW18a, Rmk. 4.10] for a similar remark about quantum symmetric pairs. More generally, if $(A,R)$ is a quasitriangular bialgebra, then so is $(A^{\operatorname{cop}},R_{21})$. Hence $(A^{\operatorname{cop}}\otimes A,R_{31}\cdot R_{24})$ is quasitriangular, and naturally equipped with a (diagonal) cylindrical structure with respect to the coideal subalgebra $\Delta(A)$, given by the twist pair $((12),1\otimes 1)$, and K-matrix $R$. $\triangledown$
It is beneficial to require an additional condition of a cylindrical structure.
Definition B.2.2.
Let $(\psi,J,K)$ be a cylindrical structure on $(A,R)$ with respect to $B$. If
$R^{\psi}_{21}\cdot 1\otimes K\cdot R\quad\in\quad B\otimes A$ (B.7)
then we call $(\psi,J,K)$ supported on $B$. $\triangledown$
Note that (B.7) ties the objects $B$ and $K$ more closely together: without it, any cylindrical structure $(\psi,J,K)$ on $(A,R)$ with respect to $B$ would also be with respect to any coideal subalgebra $B^{\prime}$ of $A$ contained in $B$, see also [Kol20, Rmk. 2.11]. Furthermore, we obtain a second derivation of (B.6) by combining (B.7) with (B.4).
Finally, since $B$ is a coideal subalgebra, (B.7) permits the module category ${\operatorname{Mod}}(B)$ over the monoidal category ${\operatorname{Mod}}(A)$ to be endowed with a braided structure compatible with the braided structure on ${\operatorname{Mod}}(A)$. We develop the categorical framework in Appendix C. The key algebraic ingredient is the universal tensor K-matrix.
B.3. Tensor K-matrices and reflection bialgebras
In a similar spirit, we introduce the notion of a reflection bialgebra. This is the natural generalization (to arbitrary twist pairs) of the notions of reflection algebra, given by Enriquez in [Enr07, Def. 4.1], and quasitriangular comodule algebra, given by Kolb in [Kol20, Def. 2.7].
Definition B.3.1.
A reflection bialgebra consists of a tuple $(A,R,B,\psi,J)$ as in Section B.2 and a universal tensor K-matrix ${{\mathbb{K}}}$ with respect to $(B,\psi,J)$, i.e., an invertible element of $B\otimes A$ satisfying the QSP intertwining identity
${{\mathbb{K}}}\cdot\Delta(b)=({\operatorname{id}}\otimes\psi)(\Delta(b))\cdot{{\mathbb{K}}}\qquad(b\in B)$ (B.8)
and the coproduct identities
$(\Delta\otimes{\operatorname{id}})({{\mathbb{K}}})$ $=R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23},$ (B.9)
$({\operatorname{id}}\otimes\Delta)({{\mathbb{K}}})$ $=J_{23}^{-1}\cdot{{\mathbb{K}}}_{13}\cdot R^{\psi}_{23}\cdot{{\mathbb{K}}}_{12}.$ (B.10)
Equivalently, we shall also say that $(\psi,J,{{\mathbb{K}}})$ is a reflection structure on $(A,R)$ with respect to $B$. $\triangledown$
Note that by applying ${\operatorname{id}}\otimes{\operatorname{id}}\otimes\epsilon$ to (B.10) we obtain the normalization
$({\operatorname{id}}\otimes\varepsilon)({{\mathbb{K}}})=1.$
Similar to the situation for a cylindrical structure $(\psi,J,K)$ supported on $B$, one readily deduces the following reflection equation in $A^{\otimes 3}$:
$R^{\psi\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R^{\psi}_{23}\cdot{{\mathbb{K}}}_{12}={{\mathbb{K}}}_{12}\cdot R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23}.$ (B.11)
Indeed, one can combine ${{\mathbb{K}}}\in B\otimes A$ with (B.8) and (B.9) or use (B.1), (B.10) and the twist pair property.
We will see that a reflection structure is equivalent to a cylindrical structure supported on $B$. First of all, in analogy with (part of) [Kol20, Lem. 2.9] (cf. [Enr07, Rmk. 4.2]), reflection structures straightforwardly induce cylindrical structures.
Lemma B.3.2.
Let $(A,R,B,\psi,J,{{\mathbb{K}}})$ be a reflection bialgebra and set
$K=(\varepsilon\otimes{\operatorname{id}})({{\mathbb{K}}})\in A.$
Then $(A,R,B,\psi,J,K)$ is a cylindrical bialgebra, i.e., $K$ is a universal K-matrix with respect to $(B,\psi,J)$. Furthermore,
${{\mathbb{K}}}=R^{\psi}_{21}\cdot 1\otimes K\cdot R.$ (B.12)
Proof.
This follows from a straightforward computation, almost identical to [Kol20, Lem. 2.9]:
${{\mathbb{K}}}=(((\varepsilon\otimes{\operatorname{id}})\circ\Delta)\otimes{\operatorname{id}})({{\mathbb{K}}})=(\varepsilon\otimes{\operatorname{id}}\otimes{\operatorname{id}})(R^{\psi}_{32}\cdot{{\mathbb{K}}}_{13}\cdot R_{23})=R^{\psi}_{21}\cdot 1\otimes K\cdot R.$ (B.13)
Then (B.4) follows from (B.8), while (B.5) follows from (B.10). ∎
Remark B.3.3.
If $\psi$ is a bialgebra automorphism and $J=R_{21}^{-1}$, the setup of Kolb in [Kol20, Sec. 2] is recovered. Note that in Kolb’s approach the twist automorphism $\psi$ is removed from the axiomatics by replacing ${\operatorname{Mod}}(A)$ by an equivariantized category. However, we are mainly interested in the case where $A$ is a q-deformed Kac-Moody algebra $U_{q}({\mathfrak{g}})$, which is a reflection bialgebra up to completion with respect to a full subcategory ${\mathcal{C}}$ of ${\operatorname{Mod}}(A)$, see Theorem 5.6.1. Since ${\mathcal{C}}$ is not preserved by the pullback of $\psi$ (except when ${\mathfrak{g}}$ is of finite type, which is assumed in [Kol20]), equivariantization is not applicable in our setting. $\triangledown$
The concept of a universal tensor K-matrix generalizes to the case where $B$ is a right $A$-comodule algebra with coaction map $\Delta_{B}$ (by definition, a linear map$:B\to B\otimes A$ satisfying $(\Delta_{B}\otimes{\operatorname{id}}_{A})\circ\Delta_{B}=({\operatorname{id}}_{B}\otimes\Delta)\circ\Delta_{B}$), which is the setting in [LBG23, Le23]. In order to formulate the generalized notion of universal tensor K-matrix, simply replace the coproduct $\Delta$ by $\Delta_{B}$ in axioms (B.8-B.9). If $B$ is a subalgebra of $A$ and $\Delta_{B}=\Delta|_{B}$, then the situation described above is recovered191919 More generally, if there exists an algebra map $\epsilon_{B}:B\to{\mathbb{F}}$ (i.e., $B$ is an augmented algebra) then $\widetilde{B}=((\epsilon_{B}\otimes{\operatorname{id}}_{A})\circ\Delta_{B})(B)$ is a right coideal subalgebra of $A$ and a universal tensor K-matrix ${{\mathbb{K}}}\in B\otimes A$ with respect to $(A,R,B,\psi,J)$ induces a cylindrical structure $(\psi,J,K)$ on $(A,R)$ with respect to $\widetilde{B}$ via $K=(\epsilon_{B}\otimes{\operatorname{id}})({{\mathbb{K}}})\in A$. It has the property $R^{\psi}_{21}\cdot 1\otimes K\cdot R\in B\otimes A$. A natural sufficient condition for this cylindrical structure to be supported on $\widetilde{B}$ is isomorphicity of $B$ and $\widetilde{B}$ as right $A$-comodule algebras, i.e., injectivity of $(\epsilon_{B}\otimes{\operatorname{id}}_{A})\circ\Delta_{B}$. .
B.4. Gauge transformations
Given a quasitriangular bialgebra $(A,R)$ and a coideal subalgebra $B$. Then it is straightforward to check that the following assignments define an action of $A^{\times}$ on cylindrical structures called gauge transformation (also see [AV22a, Rmk. 8.11]):
$(\psi,J,K)\mapsto(\operatorname{Ad}(g)\circ\psi,(g\otimes g)\cdot J\cdot\Delta(g)^{-1},g\cdot K),\qquad g\in A^{\times}.$
This immediately lifts to an action on reflection structures, with the action on the tensor K-matrix given by
${{\mathbb{K}}}\mapsto(1\otimes g)\cdot{{\mathbb{K}}},$ (B.14)
which is compatible with Lemma B.3.2. Note that the support condition (B.7) is invariant under this action so that, in order to prove (B.7), it suffices to do so for a convenient choice of gauge transformation.
B.5. Weakly cylindrical bialgebras
Several authors [Ta92, Re95, Ga97] have considered a generalization of the notion of a quasitriangular bialgebra where the role of the universal R-matrix is played by an algebra automorphism. A boundary analogue of this, generalizing the notion of reflection algebra with a distinguished comodule algebra, was considered by [KY20] in a very general context of pairs of suitable Nichols algebras and comodule algebras. To describe both types of structures we will mostly use the conventions from that paper, in particular the terminology weak.
In particular, a bialgebra $A$ with coproduct $\Delta$ is called weakly quasitriangular if there exists an algebra automorphism $\mathcal{S}$ of $A\otimes A$ fixing $\Delta(A)\subset A\otimes A$ pointwise such that
$({\operatorname{id}}\otimes\Delta)\circ\mathcal{S}$ $=(\mathcal{S}\otimes{\operatorname{id}})\circ({\operatorname{id}}\otimes\mathcal{S})\circ(\Delta\otimes{\operatorname{id}}),$
$(\Delta\otimes{\operatorname{id}})\circ\mathcal{S}$ $=({\operatorname{id}}\otimes\mathcal{S})\circ(\mathcal{S}\otimes{\operatorname{id}})\circ({\operatorname{id}}\otimes\Delta).$
Quasitriangular bialgebras are weakly quasitriangular: just set $\mathcal{S}$ equal to $(12)\circ\operatorname{Ad}(R)$.
We call a weakly quasitriangular bialgebra $(A,\mathcal{S})$ weakly cylindrical if there exists a right coideal subalgebra $B\subseteq A$ and an algebra automorphism $\kappa:A\to A$ fixing $B$ pointwise such that
$\Delta\circ\kappa=(\kappa\otimes{\operatorname{id}})\circ\mathcal{S}\circ(\kappa\otimes{\operatorname{id}})\circ\Delta.$ (B.15)
Further, we say that the weakly cylindrical structure $\kappa$ on $(A,\mathcal{S})$ is supported on $B$ if the automorphism $\mathcal{S}\circ(\kappa\otimes{\operatorname{id}})\circ\mathcal{S}$ preserves $B\otimes A$.
We readily obtain that cylindrical bialgebras with twist pair $(\psi,J)$ are weakly cylindrical by setting $\kappa=\psi^{-1}\circ\operatorname{Ad}(K)$, i.e. $\kappa=\xi^{-1}$ in the notation of Section 4, and this weakly cylindrical structure is supported on $B$ if the cylindrical structure $(\psi,J,K)$ is supported on $B$. Note that the verification of (B.15) relies on the fact that $A^{\operatorname{cop},\psi}=A_{J}$.
A weak reflection bialgebra is a weakly quasitriangular bialgebra $(A,\mathcal{S})$ together with a right coideal subalgebra $B\subseteq A$ and an automorphism ${\mathcal{K}}$ of $B\otimes A$ fixing $\Delta(B)$ pointwise such that
$(\Delta\otimes{\operatorname{id}})\circ{\mathcal{K}}$ $=({\operatorname{id}}\otimes\mathcal{S})\circ({\mathcal{K}}\otimes{\operatorname{id}})\circ({\operatorname{id}}\otimes\mathcal{S})\circ(\Delta\otimes{\operatorname{id}}),$ (B.16)
$({\operatorname{id}}\otimes\Delta)\circ{\mathcal{K}}$ $=({\mathcal{K}}\otimes{\operatorname{id}})\circ({\operatorname{id}}\otimes\mathcal{S})\circ({\mathcal{K}}\otimes{\operatorname{id}})\circ({\operatorname{id}}\otimes\Delta).$ (B.17)
Generalizing the coideal subalgebra of $A$ to a comodule algebra over $A$ as at the end of Section B.3, this coincides with the notion of a weakly quasitriangular comodule algebra, see [KY20, Def. 6.11].
Remarks B.5.1.
  1. (1)
    The definition of weak cylindricity is clearly simpler than the definition of cylindricity involving the twist pair $(\psi,J)$. On the other hand, for the representation theory of quantum loop algebras it is natural to separate the role of the universal K-matrix and the twist automorphism.
  2. (2)
    In the traditional (non-weak) formalism for quasitriangularity and cylindricity, typically universal solutions are constructed in a completion of the bialgebra with respect to a certain category of representations. If the representation theory of the bialgebra is intractable then the weak formalisms provide workarounds, although the use of completions, see [KY20, Sec. 6.3], is still necessary [^20] Indeed, if $(A,B)$ is a quantum symmetric pair, then $(12)\circ\operatorname{Ad}(R)$ is not an automorphism of $A\otimes A$, although it fixes $\Delta(A)$ pointwise, and similarly $\psi^{-1}\circ\operatorname{Ad}(K)$ is not an automorphism of $A$, although it fixes $B$ pointwise..
  3. (3)
    If $A$ is a Hopf algebra then $B\otimes A^{\otimes 2}$ is generated212121 To see this, note that $a\otimes 1=\Delta(a^{(1)})\cdot(1\otimes S(a^{(2)}))$ for all $a\in A$. as a unital algebra by the subalgebras $\Delta(B)\otimes A$ and $B\otimes\Delta(A)$. In this case, the above axioms for the maps $\mathcal{S}$, $\kappa$ and ${\mathcal{K}}$ imply the weak analogues of Yang-Baxter and (basic and tensor) reflection equations:
    $({\operatorname{id}}\otimes\mathcal{S})\cdot(\mathcal{S}\otimes{\operatorname{id}})\cdot({\operatorname{id}}\otimes\mathcal{S})$ $=(\mathcal{S}\otimes{\operatorname{id}})\cdot({\operatorname{id}}\otimes\mathcal{S})\cdot(\mathcal{S}\otimes{\operatorname{id}}),$ (B.18)
    $\mathcal{S}\cdot(\kappa\otimes{\operatorname{id}})\cdot\mathcal{S}\cdot(\kappa\otimes{\operatorname{id}})$ $=(\kappa\otimes{\operatorname{id}})\cdot\mathcal{S}\cdot(\kappa\otimes{\operatorname{id}})\cdot\mathcal{S},$ (B.19)
    $({\operatorname{id}}\otimes\mathcal{S})\cdot({\mathcal{K}}\otimes{\operatorname{id}})\cdot({\operatorname{id}}\otimes\mathcal{S})\cdot({\mathcal{K}}\otimes{\operatorname{id}})$ $=({\mathcal{K}}\otimes{\operatorname{id}})\cdot({\operatorname{id}}\otimes\mathcal{S})\cdot({\mathcal{K}}\otimes{\operatorname{id}})\cdot({\operatorname{id}}\otimes\mathcal{S}),$ (B.20)
    equations for algebra automorphisms of $A^{\otimes 3}$, $A^{\otimes 2}$ and $B\otimes A^{\otimes 2}$, respectively. Hence, we get a representation of the Artin-Tits braid group of type Bn on $B\otimes A^{\otimes n}$ for any $n\in{\mathbb{Z}}_{>0}$, cf. [AV22a, Prop. 2.4]. In fact, for suitable Nichols algebras, by [KY20, Thms. 6.9, 6.15] one has weak analogues of a system of parameter-independent Yang-Baxter equations and left reflection equation in the style of Cherednik, see [Che92, Sec. 4], cf. [AV22a, Sec. 2.5]. $\triangledown$
B.6. From K-matrices to tensor K-matrices
Returning to the approach of [Kol20], we ask the natural question when a cylindrical bialgebra $(A,R,B,\psi,J,K)$ can be promoted to a reflection bialgebra $(A,R,B,\psi,J,{{\mathbb{K}}})$ with ${{\mathbb{K}}}=R^{\psi}_{21}\cdot 1\otimes K\cdot R$. From Lemma B.3.2 it follows that the support condition (B.7) must be satisfied. It turns out that this is sufficient.
Proposition B.6.1.
Let $(A,R,B,\psi,J,K)$ be a cylindrical bialgebra and define ${{\mathbb{K}}}\in A\otimes A$ by (B.12).
  1. (1)
    The identities (B.8-B.10) are satisfied and $K=(\varepsilon\otimes{\operatorname{id}})({{\mathbb{K}}})$.
  2. (2)
    ${{\mathbb{K}}}$ is a universal tensor K-matrix with respect to $(B,\psi,J)$ (in other words, $(A,R,B,\psi,J,{{\mathbb{K}}})$ is a reflection bialgebra) if and only if (B.7) is satisfied.
Proof.
(1) For any $b\in B$ we have
$R^{\psi}_{21}\cdot 1\otimes K\cdot R\cdot\Delta(b)$ $=R^{\psi}_{21}\cdot({\operatorname{id}}\otimes\psi)(\Delta^{\operatorname{op}}(b))\cdot 1\otimes K\cdot R$ (B.21)
$=({\operatorname{id}}\otimes\psi)(R_{21}\cdot\Delta^{\operatorname{op}}(b))\cdot 1\otimes K\cdot R$ (B.22)
$=({\operatorname{id}}\otimes\psi)(\Delta(b))\cdot R^{\psi}_{21}\cdot 1\otimes K\cdot R,$ (B.23)
where we have used (B.1), (B.4) and the coideal property $\Delta(B)\subseteq B\otimes A$, thus establishing the QSP intertwining identity (B.8). The first coproduct identity (B.9) follows from the identities (B.2) for the R-matrix:
$(\Delta\otimes{\operatorname{id}})(R^{\psi}_{21}\cdot 1\otimes K\cdot R)$ $=({\operatorname{id}}\otimes{\operatorname{id}}\otimes\psi)(R_{32}\cdot R_{31})\cdot(1\otimes 1\otimes K)\cdot R_{13}\cdot R_{23}$ (B.24)
$=R^{\psi}_{32}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes K)\cdot R_{13}\cdot R_{23}.$ (B.25)
On the other hand, the second coproduct identity (B.10) follows from (B.5) and the fact that $(\psi,J)$ is a twist pair. Indeed, since $\Delta_{J}=\Delta^{\operatorname{op},\psi}$, one has
$({\operatorname{id}}\otimes\Delta)(R^{\psi}_{21})\cdot J^{-1}_{23}$ $=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\Delta_{J})(R^{\psi}_{21})$ (B.26)
$=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\Delta^{\operatorname{op},\psi})(R^{\psi}_{21})$ (B.27)
$=J^{-1}_{23}\cdot({\operatorname{id}}\otimes\psi\otimes\psi)({\operatorname{id}}\otimes\Delta^{\operatorname{op}})(R_{21})$ (B.28)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot R^{\psi}_{21},$ (B.29)
therefore
$({\operatorname{id}}\otimes\Delta)$ $(R^{\psi}_{21}\cdot 1\otimes K\cdot R)=$ (B.30)
$=({\operatorname{id}}\otimes\Delta)(R^{\psi}_{21})\cdot J^{-1}_{23}\cdot(1\otimes 1\otimes K)\cdot R^{\psi}_{23}\cdot(1\otimes K\otimes 1)\cdot R_{13}\cdot R_{12}$ (B.31)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot R^{\psi}_{21}\cdot(1\otimes 1\otimes K)\cdot R^{\psi}_{23}\cdot(1\otimes K\otimes 1)\cdot R_{13}\cdot R_{12}$ (B.32)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes K)\cdot R^{\psi}_{21}\cdot R^{\psi}_{23}\cdot R_{13}\cdot(1\otimes K\otimes 1)\cdot R_{12}$ (B.33)
$=J^{-1}_{23}\cdot R^{\psi}_{31}\cdot(1\otimes 1\otimes K)\cdot R_{13}\cdot R^{\psi}_{23}\cdot R^{\psi}_{21}\cdot(1\otimes K\otimes 1)\cdot R_{12}$ (B.34)
where in the last equality we used (B.3). Finally, note that
$(\varepsilon\otimes{\operatorname{id}})({{\mathbb{K}}})=(\varepsilon\otimes{\operatorname{id}})(R^{\psi}_{21}\cdot 1\otimes K\cdot R)=K.$ (B.35)
(2) If ${{\mathbb{K}}}=R^{\psi}_{21}\cdot 1\otimes K\cdot R$ is a universal tensor K-matrix with respect to $(B,\psi,J)$ then (B.7) is just the defining condition ${{\mathbb{K}}}\in B\otimes A$. ∎
Appendix C Boundary bimodule categories
In this section, we introduce the notion of a boundary bimodule category tailored around the category of modules over a reflection bialgebra, see Section B.3. This framework encodes in particular the action of the tensor K-matrix of a quantum symmetric pair of arbitrary type. Thus, it is very similar, in spirit, with the approach used in [Enr07, Bro13, Kol20], and more recently in [LYW23].
As braided monoidal categories are equipped with a natural action of the braid group on the tensor powers of their objects, boundary bimodule categories give rise to representations of a cylindrical braid groupoid whose definition relies on Cherednik’s generalized reflection equation (see Section C.6).
Conventions. A boundary bimodule category is a category $\mathcal{M}$ acted upon by a braided monoidal category ${\mathcal{C}}$ and equipped with an extra structure. In the cases of our interest, the monoidal structure on ${\mathcal{C}}$ is always given by the ordinary tensor product of vector spaces. Hence the associativity and the unit constraints will be omitted for simplicity. In contrast, the monoidal action of ${\mathcal{C}}$ on $\mathcal{M}$, while being canonically unital, has a non–canonical associativity constraint, which will therefore not be omitted.
C.1. Module categories
We briefly review the notion of a module category $\mathcal{M}$ over a (braided) monoidal category ${\mathcal{C}}$ as presented in [Har01, Kol20]. Let $\mathcal{C}$ be a monoidal category with tensor product $\otimes$ and unit $\mathbf{1}$.
A (right) monoidal action of $\mathcal{C}$ on a category $\mathcal{M}$ is a functor :×$\vartriangleleft\colon\mathcal{M}\times\mathcal{C}\to\mathcal{M}$ together with an associativity constraint
$\Phi\colon\vartriangleleft\circ({\operatorname{id}}\times\otimes)\to\vartriangleleft\circ(\vartriangleleft\times{\operatorname{id}})$ (C.1)
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, $M\vartriangleleft{\bf 1}=M$;
  2. $\bullet$
    for any $M\in\mathcal{M}$, $V\in{\mathcal{C}}$, $\Phi_{M{\bf 1}V}={\operatorname{id}}_{M\vartriangleleft V}=\Phi_{MV{\bf 1}}$;
  3. $\bullet$
    for any $M\in\mathcal{M}$ and $U,V,W\in{\mathcal{C}}$,
[…]
${\bf J}_{MV}\colon F(M)\vartriangleleft V\to F(M\vartriangleleft V)$ (C.3)
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, ${\bf J}_{M{\bf 1}}={\operatorname{id}}_{F(M)}$;
  2. $\bullet$
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$,
[…]
Henceforth, we refer to the datum $(F,{\bf J})$ as a module functor $\mathcal{M}\to{\mathcal{N}}$.
Example C.1.1.
Set $\mathcal{C}={\operatorname{Mod}}{(U_{q}({\mathfrak{g}}))}$ and $\mathcal{M}={\operatorname{Mod}}{(U_{q}(\mathfrak{k}))}$. Since $U_{q}(\mathfrak{k})$ is a coideal subalgebra in $U_{q}({\mathfrak{g}})$, the usual tensor product induces a functor
$\vartriangleleft\colon{\operatorname{Mod}}{(U_{q}(\mathfrak{k}))}\times{\operatorname{Mod}}{(U_{q}({\mathfrak{g}}))}\to{\operatorname{Mod}}{(U_{q}(\mathfrak{k}))}$
which is readily verified to be a monoidal action of ${\operatorname{Mod}}{(U_{q}({\mathfrak{g}}))}$ on ${\operatorname{Mod}}{(U_{q}(\mathfrak{k}))}$. $\triangledown$
C.2. Braided module categories
Let ${\mathcal{C}}$ be a braided monoidal category with braiding ${\bf c}\colon\otimes\to\otimes^{\scriptscriptstyle\operatorname{op}}$ and $\mathcal{M}$ a right module category over ${\mathcal{C}}$. For any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$, set
$({\operatorname{id}}_{M}\vartriangleleft{\bf c}_{VW})_{\Phi}=\Phi_{MWV}\circ({\operatorname{id}}_{M}\vartriangleleft{\bf c}_{VW})\circ\Phi_{MVW}^{-1}$
Following [Bro13, BZBJ18, Kol20], a module braiding on $\mathcal{M}$ is a natural automorphism
${\bf d}_{MV}\colon M\vartriangleleft V\to M\vartriangleleft V$
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, ${\bf d}_{M{\bf 1}}={\operatorname{id}}_{M}$;
  2. $\bullet$
    for any $M\in\mathcal{M}$, $V,W\in{\mathcal{C}}$,
[…]
  3. $\bullet$
    for any $M\in\mathcal{M}$, $V,W\in{\mathcal{C}}$,
[…]
A module functor $F:\mathcal{M}\to{\mathcal{N}}$ is braided if the module structure on $F$ intertwines the braidings, i.e., if the diagram
[…]
commutes for any $M\in\mathcal{M}$ and $V\in{\mathcal{C}}$.
Remarks C.2.1.
  1. (1)
    The identity ${\operatorname{id}}_{M\vartriangleleft V}$ does not define a module braiding on $\mathcal{M}$ unless the monoidal category ${\mathcal{C}}$ is symmetric, i.e., ${\bf c}_{WV}={\bf c}_{VW}^{-1}$.
  2. (2)
    It is pointed out in [BZBJ18, Rmk. 3.6] that there is an infinite family of possible axioms for a braided module category. The relations (C.5) and (C.6) constitute just one example of such axioms.
  3. (3)
    In [Kol20], Kolb proved that the action of the tensor K-matrix of a quantum symmetric pair $U_{q}(\mathfrak{k})\subset U_{q}({\mathfrak{g}})$ with $\dim{\mathfrak{g}}<\infty$ gives rise to a braided module structure on ${\operatorname{Mod}_{\sf fd}}(U_{q}(\mathfrak{k}))$ over ${\operatorname{Mod}_{\sf fd}}(U_{q}({\mathfrak{g}}))$. This is no longer true when $\dim{\mathfrak{g}}=\infty$, which motivates the notion of a boundary bimodule category, which we introduce in Sections C.3-C.4. $\triangledown$
C.3. Bimodule categories
Let ${\mathcal{C}}$ be a monoidal category. A left monoidal action of ${\mathcal{C}}$ on a category $\mathcal{M}$ is a functor :×$\blacktriangleleft\colon\mathcal{M}\times{\mathcal{C}}\to\mathcal{M}$ together with a natural isomorphism
$\Phi_{MVW}:M\blacktriangleleft(V\otimes W)\to(M\blacktriangleleft W)\blacktriangleleft V$
for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$.
A bimodule category over ${\mathcal{C}}$ is a category $\mathcal{M}$ equipped with
  1. $\bullet$
    a right monoidal action :×$\vartriangleleft\colon\mathcal{M}\times{\mathcal{C}}\to\mathcal{M}$ with associativity constraints $\Phi^{\vartriangleleft}$;
  2. $\bullet$
    a left monoidal action :×$\blacktriangleleft\colon\mathcal{M}\times{\mathcal{C}}\to\mathcal{M}$ with associativity constraints $\Phi^{\blacktriangleleft}$;
  3. $\bullet$
    a natural isomorphism
    ${\bf e}_{MVW}\colon(M\blacktriangleleft V)\vartriangleleft W\to(M\vartriangleleft W)\blacktriangleleft V$ (C.8)
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, $V\in{\mathcal{C}}$, ${\bf e}_{MV{\bf 1}}={\operatorname{id}}_{M\blacktriangleleft V}$ and ${\bf e}_{M{\bf 1}W}={\operatorname{id}}_{M\vartriangleleft W}$;
  2. $\bullet$
    for any $M\in\mathcal{M}$, $U,V,W\in{\mathcal{C}}$,
[…]
  3. $\bullet$
    for any $M\in\mathcal{M}$, $U,V,W\in{\mathcal{C}}$,
[…]
Remark C.3.1.
[…]
Let $\mathcal{M},{\mathcal{N}}$ be bimodule categories [^23] By abuse of notation, we use the same symbol to denote the monoidal actions on $\mathcal{M}$ and ${\mathcal{N}}$. and $F:\mathcal{M}\to{\mathcal{N}}$ a functor. A bimodule structure on $F$ is the datum of a pair of natural isomorphisms
${\bf J}^{\vartriangleleft}_{MV}\colon F(M)\vartriangleleft V\to F(M\vartriangleleft V)\quad\mbox{and}\quad{\bf J}^{\blacktriangleleft}_{MV}\colon F(M)\blacktriangleleft V\to F(M\blacktriangleleft V)$ (C.11)
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$, $V,W\in{\mathcal{C}}$,
[…]
  2. $\bullet$
    for any $M\in\mathcal{M}$, $V,W\in{\mathcal{C}}$,
[…]
  3. $\bullet$
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$,
[…]
C.4. Boundary bimodule categories
Let ${\mathcal{C}}$ be a braided monoidal category with braiding ${\bf c}\colon\otimes\to\otimes^{\scriptscriptstyle\operatorname{op}}$ and $\mathcal{M}$ a bimodule category. A boundary structure on $\mathcal{M}$ is a natural isomorphism
${\bf f}_{MV}\colon M\vartriangleleft V\to M\blacktriangleleft V$ (C.15)
such that
  1. $\bullet$
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$,
[…]
  2. $\bullet$
    for any $M\in\mathcal{M}$ and $V,W\in{\mathcal{C}}$,
[…]
Remark C.4.1.
Consider ${\mathcal{C}}$ as a module category over itself. Then, every braided structure on ${\mathcal{C}}$ (see Section C.2) is also a boundary with respect the bimodule structure described in Remark C.3.1. $\triangledown$
Let $\mathcal{M},{\mathcal{N}}$ be boundary bimodule categories and $F\colon\mathcal{M}\to{\mathcal{N}}$ a bimodule functor. Then, $F$ is boundary if the bimodule structures intertwines the boundary structures, i.e.,
[…]
for any $M\in\mathcal{M}$ and $V\in{\mathcal{C}}$.
C.5. From reflection bialgebras to boundary structures
The notion of a boundary structure on a bimodule category is tailored to encode the defining datum of a reflection bialgebra (see Section B.3). From
Proposition C.5.1.
Let $(A,R,\psi,J,B,{\mathbb{K}})$ be a reflection bialgebra. The following datum defines a boundary bimodule structure on ${\operatorname{Mod}}(B)$ over ${\operatorname{Mod}}(A)$.
  1. (1)
    Module structure. There is a right monoidal action
    $\vartriangleleft\colon{\operatorname{Mod}}(B)\times{\operatorname{Mod}}(A)\to{\operatorname{Mod}}(B)$
    with trivial associativity constraint given by the standard tensor product, i.e., $M\vartriangleleft V=M\otimes V$
  2. (2)
    Bimodule structure. The left monoidal action
    $\blacktriangleleft:{\operatorname{Mod}}(B)\times{\operatorname{Mod}}(A)\to{\operatorname{Mod}}(B)$
[…]
    $\Phi^{\blacktriangleleft}_{MVW}\colon M\otimes(V\otimes W)^{\psi}\to M\otimes{W^{\psi}}\otimes{V^{\psi}}$
    is provided by the action of $J$. The commutativity constraint
    ${\bf e}_{MVW}\colon M\otimes{V^{\psi}}\otimes W\to M\otimes W\otimes{V^{\psi}}$
    is provided by the action of the R-matrix $R$ on the second and third factor.
  3. (3)
    Boundary structure. The boundary structure
    ${\bf f}_{MV}\colon M\otimes V\to M\otimes{V^{\psi}}$
    is provided by the action of the tensor K-matrix ${\mathbb{K}}$.
C.6. The model of a boundary bimodule category
In a braided monoidal category, every tensor power $V^{\otimes n}$ is naturally acted upon by the ordinary braid group (Artin-Tits group of type ${\mathsf{A}}_{n-1}$). Similarly, in a braided module category, every object of the form $W\vartriangleleft V^{\otimes n}$ is naturally acted upon by the cylindrical braid group (Artin-Tits group of type ${\mathsf{B}}_{n}$). In the case of boundary bimodule categories, the symmetry is encoded instead by a cylindrical braid groupoid ${\mathsf{M}}$, whose generators satisfy constant versions of Cherednik’s generalized reflection equation [Che92, Sec. 4]. Its definition can be thought of as a boundary analogue of the free PROP, see, e.g., [Law63, Mac65, KS20].
Let ${\mathsf{C}}$ be the strict monoidal category defined as follows.
  1. $\bullet$
    The objects of ${\mathsf{C}}$ are the non–negative integers. For any $n\in{\mathbb{Z}}_{\geqslant 0}$, the corresponding object is denoted $[n]$.
  2. $\bullet$
    The tensor product in ${\mathsf{C}}$ is the sum and the unit is zero.
  3. $\bullet$
    The morphisms in ${\mathsf{C}}$ are generated by an invertible element $\mathbf{t}\in\operatorname{End}_{{\mathsf{C}}}([2])$ satisfying the following braid relation in $\operatorname{End}_{{\mathsf{C}}}([3])$:
    $\mathbf{t}_{12}\mathbf{t}_{23}\mathbf{t}_{12}=\mathbf{t}_{23}\mathbf{t}_{12}\mathbf{t}_{23}.$ (C.19)
Remark C.6.1.
Let $V$ be an object in a braided monoidal category ${\mathcal{C}}$ with braiding $\mathbf{c}$. Then, there is an essentially unique242424 The functor ${\mathcal{G}}$ depends upon the choice of a bracketing on $n$ elements for every $n$, see, e.g., [ATL19, Sec. 7] for the precise definition. monoidal functor ${\mathcal{G}}:{\mathsf{C}}\to{\mathcal{C}}$ such that ${\mathcal{G}}([n])=V^{\otimes n}$ and ${\mathcal{G}}(\mathbf{t})=\mathbf{c}_{VV}$. $\triangledown$
Let ${\mathsf{M}}$ be the strict bimodule category over ${\mathsf{C}}$ defined as follows.
  1. $\bullet$
    The objects of ${\mathsf{M}}$ are given by {}×$\{\star\}\times\coprod_{n\geqslant 0}{\mathsf{A}}^{n}$, where ${\mathsf{A}}=\{\pm\}$.
  2. $\bullet$
    The morphisms in ${\mathsf{M}}$ are generated by the following elements.
    • For any ${{\mathsf{s}}}\in{\mathsf{M}}$, three invertible elements
      $\mathbf{t}^{++}\in\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}++\,,\,{\mathsf{s}}++)$ (C.20)
      $\mathbf{t}^{--}\in\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}--\,,\,{\mathsf{s}}--)$ (C.21)
      $\mathbf{t}^{-+}\in\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}-+\,,\,{\mathsf{s}}+-)$ (C.22)
      subject to the mixed braid relations
      $\mathbf{t}^{\tau\zeta}_{12}\mathbf{t}^{\sigma\zeta}_{23}\mathbf{t}^{\sigma\tau}_{12}=\mathbf{t}^{\sigma\tau}_{23}\mathbf{t}^{\sigma\zeta}_{12}\mathbf{t}^{\tau\zeta}_{23}$ (C.23)
      in $\operatorname{Hom}_{\mathsf{C}}({\mathsf{s}}\,\sigma\,\tau\,\zeta,{\mathsf{s}}\,\zeta\,\tau\,\sigma)$, for all $\sigma,\tau,\zeta\in{\mathsf{A}}$ such that the corresponding operators are all defined.
    • For any ${{\mathsf{s}}}\in{\mathsf{M}}$, an invertible element
      $\mathbf{b}\in\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}\,+\,,\,{\mathsf{s}}\,-)$
      subject to Cherednik’s reflection equation in $\operatorname{Hom}_{\mathsf{M}}({\mathsf{s}}\,++\,,\,{\mathsf{s}}\,--)$:
      $\mathbf{t}^{--}\mathbf{b}\mathbf{t}^{-+}\mathbf{b}=\mathbf{b}\mathbf{t}^{-+}\mathbf{b}\mathbf{t}^{++}.$ (C.24)
  3. $\bullet$
    The right and left actions of $[n]\in{\mathsf{C}}$ are given by concatenation with length-$n$ sequences of $+$ and --, respectively. The associativity constraints are trivial. The commutativity constraint (C.8) is given by $\mathbf{t}^{-+}$. Finally, the boundary structure (C.15) is given by $\mathbf{b}$.
Proposition C.6.2.
Let ${\mathcal{C}}$ be a braided monoidal category and $\mathcal{M}$ a boundary bimodule category.
  1. (1)
    For any $V\in{\mathcal{C}}$, the functor ${\mathcal{G}}:{\mathsf{C}}\to{\mathcal{C}}$ defined in Remark C.6.1 induces on $\mathcal{M}$ a boundary bimodule structure over ${\mathsf{C}}$.
  2. (2)
    For any $M\in\mathcal{M}$, there is an essentially unique boundary bimodule functor ${\mathcal{F}}:{\mathsf{M}}\to\mathcal{M}$ such that ${\mathcal{F}}(\star)=M$ and ${\mathcal{F}}(\mathbf{b})={\bf f}$.